### v1

# SelfDefend: LLMs Can Defend Themselves against Jailbreaking in a Practical Manner

Abstract
Jailbreaking is an emerging adversarial attack that bypasses the safety alignment deployed in off-the-shelf large language models (LLMs) and has evolved into four major categories: optimization-based attacks such as the Greedy Coordinate Gradient (GCG) attack, jailbreak template-based attacks such as “Do-Anything-Now” (DAN), advanced indirect attacks like DrAttack, and multilingual jailbreaks. However, delivering a practical jailbreak defense is challenging because it needs to not only handle all the above jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs.
Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM defense instance to concurrently protect the target LLM instance in the normal stack and collaborate with it for checkpoint-based access control. The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful prompts or intentions in user queries, which we empirically validate using the commonly used GPT-3.5/4 models across all major jailbreak attacks. Our measurements show that SelfDefend enables GPT-3.5 to suppress the attack success rate (ASR) by 8.97-95.74% (average: 60%) and GPT-4 by even 36.36-100% (average: 83%), while incurring negligible effects on normal queries. To further improve the defense’s robustness and minimize costs, we employ a data distillation approach to tune dedicated open-source defense models. These models outperform four SOTA defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays. We also empirically show that the tuned models are robust to targeted GCG and prompt injection attacks.
1 Introduction
Recent years have witnessed the significant potential of large language models (LLMs) in various domains [2], such as natural language processing (NLP) [3, 4, 5], information retrieval [6], image generation [7], science [8, 9, 10, 11], code tasks [12, 13, 14, 15], security tasks [16, 17, 18, 19, 20, 21], and more. To avoid causing social anxiety, ethical, and legal issues due to LLM responses to harmful questions, LLM vendors typically conduct safety alignment to prevent the misuse of LLMs through techniques like RLHF (Reinforcement Learning from Human Feedback) [22]. In response to a harmful prompt that violates safety policies, an aligned LLM often replies with a standard response such as “I’m sorry, I can’t assist with that request.” To bypass LLMs’ safety alignment checks, an adversarial attack known as jailbreaking [23] was proposed.
In the past two years, research on LLM jailbreak attacks and defenses has attracted considerable interest, with most of them focused on the offensive side. Specifically, jailbreak strategies have evolved from manual prompt engineering [23, 24, 25] to automatic LLM-based red teaming [26, 27, 28]. Besides these template-based jailbreaks aimed at identifying a valid jailbreak prompt template, a more generic, optimization-based adversarial jailbreak approach, notably Greedy Coordinate Gradient (GCG) [29], was recently proposed. It uses public available models to train adversarial suffixes that maximize the probability for LLMs to produce an affirmative response rather than refusing to answer. These studies [29, 30, 31] found that the identified suffixes are transferable to closed-source off-the-shelf LLMs. More recently, advanced indirect jailbreaks like DrAttack [32] and Puzzler [33], as well as multilingual jailbreaks [34, 35], have also been invented. In addition to proposing these four major types of jailbreaks, various benchmark studies on LLM jailbreak attacks [23, 36, 37, 38, 39, 40] have also been conducted.
On the contrary, the defensive side is somewhat overshadowed, despite around a dozen defense mechanisms being proposed in the past year. They can be roughly categorized into tuning-based and non-tuning-based mechanisms. Specifically, tuning-based defenses [41, 42, 43, 44, 45, 46] aim to fundamentally improve a model’s safety alignment against jailbreaking, while non-tuning-based defenses [47, 48, 49, 50, 51, 52, 53, 54, 55] can be typically applied to any off-the-shelf LLMs. We conduct an analysis of these defense techniques in §3 and find that it is still challenging for them to be widely used in practice. In short, we advocate that a practical jailbreak defense needs to not only handle all the aforementioned jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs in various scenarios (see details in §3).
In this paper, we propose a new perspective on practically defending against jailbreak attacks, inspired by how the traditional security concept of shadow stacks [56] defends against memory overflow attacks. Similar to the shadow stack creating a shadow memory space, we establish a shadow LLM defense instance, $LLM_{defense}$, alongside the target LLM instance, $LLM_{target}$, in the normal stack. Under this framework, $LLM_{target}$ can process any user prompt query $P_{query}$ as usual to produce a token-by-token output. Meanwhile, $LLM_{defense}$ employs a dedicated detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$). Such a unique setup brings several benefits. ① It simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate due to this dual-layer protection. ② Since $LLM_{defense}$’s output is typically short, such as only “No” (indicating no issue) for normal queries, a checkpoint in the normal stack tends to be quickly triggered from the shadow stack without delaying $LLM_{target}$’s output. ③ Since $LLM_{defense}$ does not need to modify or monitor $LLM_{target}$’s internals, it can protect both open-source and closed-source LLMs. We concretize the above ideas into a generic jailbreak defense framework called SelfDefend and will introduce its details in §4.
The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful portions (prompts/intentions) in user queries. To validate this hypothesis, we conduct an empirical measurement in §5 using the commonly used GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [29, 31, 36, 28, 32, 33, 34]. The results are quite promising in that SelfDefend enables both GPT-3.5 and GPT-4 to significantly suppress the attack success rate (ASR). Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97-95.74% (average: 60%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.28, and GPT-4-based SelfDefend even reduces the ASR by 36.36-100% (average: 83%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.06. Besides the jailbreak scenarios, we also test SelfDefend against 805 normal prompts from the AlpacaEval dataset [57] and find that the pass rate is not affected for GPT-3.5 and slightly decreases by 3% for GPT-4, indicating that SelfDefend incurs negligible effects on normal queries. Moreover, SelfDefend incurs zero delay for over 95% of normal prompts across three out of four configurations, averaging 0.08 seconds for GPT-3.5 and 0.04 seconds for GPT-4. In various attack scenarios, SelfDefend causes an overall average extra delay of 0.04 seconds for GPT-3.5 and 0.39 seconds for GPT-4, respectively.
While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.06, GPT-4 itself is commonly known to be expensive [58]. Moreover, the closed-source nature of GPT-3.5/4 raises privacy concerns for non-OpenAI model vendors. Therefore, we attempt to tune an open-source model that can be used under the SelfDefend framework for robust, low-cost, and self-contained jailbreak defense. By conducting GPT-4-based data distillation (with SelfDefend’s prompts) on a red-team dataset from Anthropic [59] comprising 38,961 harmful and harmless prompts, we generate a large set of high-quality tuning data. The distilled datasets are then used to tune our defense models through the LoRA fine-tuning technique [60] on the publicly available Llama-2-7b model [61]. The details will be introduced in §6.
To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with four representative jailbreak defenses: Perplexity Filter [62], SmoothLLM [51], Llama Guard [42], and Llama Guard 2 [63]. The results show that SelfDefend consistently outperforms all other defenses in 19 out of 21 tested attack scenarios (seven jailbreak methods and three target LLMs: GPT-3.5, GPT-4, and Llama 2) and reaches the level of GPT-4-based SelfDefend’s defense. The only two exceptions occurred in comparison with Perplexity Filter [62], which is specifically designed to defend against the GCG attack and offers no defense against other types of jailbreaks, particularly for the GCG attack targeting GPT-3.5 and Llama 2 LLMs. Nevertheless, we still outperform it when the target LLM is GPT-4.
Besides defense effectiveness, we also measure the extra delay $\Delta d$ caused by our defense models and find that the average $\Delta d$ is negligible, at 0-0.01 seconds for normal prompts. For attack scenarios, the maximum $\Delta d$ has decreased from 1.1 seconds in GPT-4-based SelfDefend to 0.25 seconds, with $\Delta d$ in all attack scenarios except for JSAA now below 0.1 seconds. These findings indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for real-world deployment. Furthermore, we specifically assess whether the detected harmful portion actually aligns with the original prompt through the ensemble CLIP-score [64], and empirically show that the tuned models are robust to targeted GCG and prompt injection [65, 66] attacks that have white-box access to our defense models. More details are available in §7.
Availability and Ethics. To facilitate future research, we are on the way of releasing our datasets, raw evaluation results, and code at this anonymous GitHub link: https://github.com/SelfDefend. While our study required testing various jailbreak attacks, these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server. We also tried to minimize the overhead to OpenAI by spreading all the experiments over a period of two months instead of just several days and paid around 1,500 USD to OpenAI for our API usage.
2 Background
2.1 Threat Model
Attack Scenario. This research focuses on the attack scenario where an adversary seeks to perform jailbreaking on a text-based large language model (LLM). Multimodal jailbreaks [67, 68, 69, 70] are outside the scope of this paper. The objective of jailbreaking is to bypass the LLM’s safety mechanisms and induce it to generate harmful, toxic, or objectionable content. In this context, we assume the adversary has access to the LLM and can input arbitrary prompts to it. Given the vocabulary $\mathcal{T}$, the LLM, denoted as $\text{LLM}:\mathcal{T}^{\star}\to\Delta(\mathcal{T})$, takes a sequence of tokens $\mathcal{T}^{\star}$ as input and outputs a distribution $\Delta(\mathcal{T})$ over the next token in the sequence.
The adversary aims to find a prompt $P\in\mathcal{T}^{\star}$ that, when processed by the LLM, generates a response $R$ fulfilling a harmful goal $G$. We define a classifier $\text{JUDGE}:\mathcal{T}^{\star}\times\mathcal{T}^{\star}\to\{\text{True},\text{False}\}$, which returns True if and only if the response $R$ meets the criteria of the harmful goal $G$ given input $R,G$.
Adversary’s Objective. The adversary’s main objective is to generate responses from the LLM that are classified as successful jailbreaks. Specifically, the adversary seeks to maximize the probability that a response $R$ generated by the LLM for a given prompt $P$ is classified as harmful according to the goal $G$. Formally, the adversary’s objective can be expressed as:
$\sup_{P\in\mathcal{T}^{\star}}\Pr_{R\sim\text{LLM}(P)}[\text{JUDGE}(R,G)=\text{True}]$
where the randomness is due to the stochastic nature of the LLM’s responses to the input prompt $P$. In essence, the adversary iterates over potential prompts to find one that maximizes the likelihood of producing a harmful output as judged by the classifier.
Constraints and Assumptions. We assume the following constraints and assumptions in the general jailbreak threat model:
  • The adversary has black-box access to the LLM, meaning they can input prompts and observe outputs but do not have direct access to the model’s internal parameters or training data.
  • The goal string $G$ is predefined and represents a specific type of harmful content that the adversary aims to induce from the LLM.
  • The classifier JUDGE is accurate in determining whether a response constitutes a successful jailbreak according to the harmful goal $G$.
Under these assumptions, the adversary’s challenge is to discover prompts that effectively exploit the LLM’s vulnerabilities, leading to the generation of undesirable content while navigating the constraints of black-box access.
2.2 Jailbreak Attacks
Existing jailbreak attacks can be roughly grouped into four categories: optimization-based, template-based, indirect, and multilingual jailbreaks.
Optimization-based Jailbreak typically appends a suffix to a malicious prompt and iteratively updates this suffix using gradient-based or search-based methodologies for jailbreaking. The pioneering work by GCG [29] introduced a method known as the greedy coordinate gradient to optimize adversarial suffixes, facilitating transferable jailbreaks across various prompts and models. Sitawarin et al. [30] further extended this technique to GCG++ by employing a proxy model to guide the optimization process. Beyond gradient-based optimization, Andriushchenko et al. [31] utilized a simple random search on a suffix to maximize the target probability.
Template-based Jailbreak involves the use of prompt engineering [23, 24, 25] or LLM-based red teaming [26, 27, 36, 71, 28] to create effective prompt templates that can mislead LLMs into producing restricted content. Early template-based attacks combine manually crafted templates [23, 24, 25] with malicious instructions to achieve jailbreaking. For example, Liu et al. [24] proposed role-playing prompts to confuse LLMs. A more sophisticated approach, AutoDAN [28], employs genetic algorithms to optimize these hand-crafted templates. Recent advancements include the use of LLMs to assist in generating jailbreak prompts. PAIR [36] is the first method to utilize an auxiliary LLM to generate jailbreaks. Mehrotra et al. [71] then refined PAIR’s approach through a tree-based search method.
Indirect Jailbreak conceals malicious intents within the query text to circumvent the safety mechanisms of LLMs and elicit the desired malicious response. A straightforward method for executing indirect jailbreaks is by performing word substitution on the original malicious instruction [72]. Recently, DrAttack [32] introduced a technique that decomposes a malicious prompt into separate sub-prompts, effectively masking its underlying malicious intent. Puzzler [33] provides clues related to the malicious prompt, thereby guiding the target LLM towards a jailbreak.
Multilingual Jailbreak translates the harmful prompt into a language in which LLMs are less aligned for safety [34, 35]. Deng et al. [34] found that it is easier to jailbreak LLMs in low-resource languages than in high-resource languages and released a multilingual jailbreak prompt dataset, MultiJail. Besides the multilingual strategy, Wei et al. [23] encode the original harmful prompt into Base64 format to reduce the sensitivity of LLMs.
TABLE I: A comparison of existing jailbreak defenses under the four objectives we envisioned in §3.1. Note that the first six rows denote tuning-based defenses, while the last seven rows represent non-tuning-based defenses.
Venue1 Core Idea Objectives
O1 O2 O3 O4
Perplexity [41, 62] arXiv:2308.14132 Calculate the perplexity of the prompt to detect adversarial suffixes
Llama Guard [42] arXiv:2312.06674 Align the classification of both prompt and response
RPO [73], DPP [74] arXiv:2401.17263 Optimize universal and transferable suffixes that enforce safe outputs
DRO [43], PAT [44] SeT @ ICLR’24 Add a safety prompt and optimize it as a prefix using prompt tuning
SafeDecoding [45] arXiv:2402.08983 Compare and amplify the token probabilities of safety disclaimers
LED [46] arXiv:2405.18166 Identify safety-critical layers and realign them through model editing
Self Defense [47] arXiv:2308.07308 Add one more step to check the safety of original LLM responses
RAIN [48] ICLR’24 Self-evaluate each token output, rewind, and determine the final output
erase-and-check [49] arXiv:2309.02705 Erase some tokens from the prompt; check the rest using a safety filter
[50, 51, 52] arXiv:2309.14348 Perturb copies of each prompt and aggregate their output responses
IAPrompt [53] arXiv:2401.06561 Conduct a chain-of-thought intention analysis for each input prompt
GradSafe [54] arXiv:2402.13494 Compare the prompt’s gradient similarity with safety-critical gradients
Gradient Cuff [55] arXiv:2403.00867 Compare the gradient norm of refusal loss with that in benign queries
  • = applies; = partially applies; = does not apply.
  • 1
    We only present the venue information for the earliest article in each line of defense techniques.
3 Objectives and Related Work
3.1 Design Objectives
With various jailbreak attacks presented in §2, we now envision the design objectives of an ideal defense as follows:
O1
Handling all kinds of jailbreak attacks. The proposed defense should be able to handle all categories of jailbreak attacks listed in §2, including not only optimization-based adversarial jailbreaks but also template-based, indirect, and multilingual jailbreaks.
O2
Incurring negligible delay to user prompts. The proposed defense should not impact the user experience, causing either no delay or only a negligible one to normal user prompts.
O3
Providing explanations for potential jailbreak queries. Similar to O2, when the proposed defense detects any query potentially related to jailbreaking, it should provide helpful explanations on why the query is considered harmful.
O4
Compatible with both open-source and closed-source LLMs. The proposed jailbreak defense approach should protect both white-box, open-source LLMs as well as black-box, closed-source LLMs.
We clarify that compared with the other three objectives, O3 is not a mandatory objective. Nevertheless, we believe O3 is particularly important for users to forensically understand potential jailbreak queries and better protect against jailbreak attacks in the future.
3.2 Analysis of Existing Defenses
Table I summarizes our analysis of major LLM jailbreak defenses under the four objectives we envisioned above. They can be roughly categorized into tuning-based (the first six rows) and non-tuning-based (the last seven rows) mechanisms. Specifically, tuning-based defenses aim to fundamentally improve a model’s safety alignment against jailbreaking, while non-tuning-based defenses can be typically applied to any off-the-shelf LLMs to enhance their safety against jailbreak attacks. In the rest of this subsection, we present our analysis results from four perspectives:
First, most jailbreak defenses target multiple kinds of jailbreak attacks but typically do not cover advanced indirect attacks (O1: ), while a few existing defense mechanisms are specifically designed to defend against optimization-based adversarial attacks only (O1: ). The latter includes two perplexity-based filtering approaches [41, 62] and several input perturbation-based approaches [49, 50, 51, 52]. Specifically, Alon and Kamfonas [41] proposed the first tuning-based defense in the sense that they not only leveraged perplexity values as an indicator to detect prompts with adversarial suffixes but also tuned a classifier to consider both the sequence length of prompts and their perplexity for improved filtering. Another line of GCG-specific jailbreak defenses perturbed copies of the input prompt and aggregated the output responses, with SmoothLLM [51] as a representative example.
Second, the majority of prior defenses inherently incur additional delays to user prompts (O2: ), with only two exceptions. Since the design principles of most prior defenses are to check the target LLM’s responses [47, 48, 42] and internal statuses [41, 62, 45, 54, 55, 46], or to conduct extra-round analyses of the input prompt [49, 50, 51, 52, 53], it is thus inherently difficult for these approaches to avoid additional delay. The only exceptions are [73, 74, 43, 44], which conducted optimization-based suffix or prefix training for defensive purposes, with the training objective being the exact opposite of the GCG attack. As a result, the tuned models behave like normal LLMs, incurring no extra delay to user prompts.
Third, most non-tuning-based defenses have the potential to provide explanations for jailbreak queries (O1: or ), whereas half of the tuning-based defenses cannot because they rely solely on LLMs’ internal indicators for tuning (O3: ). To provide explanations for potential jailbreak queries, a defense mechanism needs to understand the semantics of incoming prompt queries. Hence, it is hard for approaches that rely solely on LLMs’ internal indicators, whether they are tuning-based [45, 46] or non-tuning-based [54, 55], to provide straightforward explanations to users.
Fourth, most non-tuning-based defenses are compatible with both open-source and closed-source LLMs (O4: ), except for those that need to leverage LLMs’ internal indicators, while the opposite is true for tuning-based defenses (O4: ). Unless they need to monitor LLMs’ internal indicators [48, 54, 55], non-tuning-based defenses, such as Self Defense [47] and IAPrompt [53], are naturally compatible with both open-source and closed-source LLMs. On the other hand, tuning-based defenses typically require white-box access to the LLMs to enable their in-depth defense. The only exception is Meta’s Llama Guard [42], which aligns the classifications of both prompt and response, and thus, theoretically, it can serve as a plug-in model to work with both open-source and closed-source LLMs.
Figure 1: A high-level overview of the SelfDefend framework and our study.
4 The SelfDefend Framework
After analyzing the pitfalls of existing defenses, we propose a new perspective on defending LLMs against jailbreaks. Our key idea is to deploy a dedicated LLM alongside the target LLM to concurrently detect potential jailbreak queries. This idea is made possible because we found that LLMs do have the capability to protect themselves by identifying harmful portions in user queries.
Based on this idea, we propose a generic LLM jailbreak defense framework called SelfDefend. As shown in Figure 1, SelfDefend creatively establishes a shadow stack alongside the normal stack in the LLM space to conduct checkpoint-based access control, which mimics traditional security defense concepts such as the shadow stack for defending against buffer overflow attacks [56] and the library-based checkpoint from SCLib [75]. We denote the target LLM in the normal stack as $LLM_{target}$ and the defense LLM in the shadow stack as $LLM_{defense}$. SelfDefend simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate. In SelfDefend, $LLM_{defense}$ can be instantiated from the same model as $LLM_{target}$, although in practice we suggest using a dedicated $LLM_{defense}$ that is robust and low-cost for detecting jailbreak queries. Since $LLM_{defense}$ does not need to modify $LLM_{target}$ or know any of its internal details, SelfDefend can protect all kinds of target LLMs, whether they are open-source or closed-source (O4: ).
Given an incoming prompt query $P_{query}$, SelfDefend dispatches it to both $LLM_{target}$ and $LLM_{defense}$ for concurrent processing. Specifically, $LLM_{target}$ processes $P_{query}$ as usual, whether it is a normal user prompt or an adversarial jailbreak prompt, but caches its token-by-token output until a checkpoint is triggered from the shadow stack. In contrast, $LLM_{defense}$ employs a dedicated detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$). Once a token of “No” (indicating no issue) or a harmful portion (prompt/intention) is output, $LLM_{defense}$ triggers a checkpoint in the normal stack so that $LLM_{target}$ can decide whether to release the token-by-token output. Since the time for outputting the token “No” is very limited, the additional delay introduced by SelfDefend is negligible for normal user prompts (O2: ). Moreover, the identified harmful portion can help generate an explainable LLM response to adversarial prompts (O3: ). These unique advantages make SelfDefend the first practical jailbreak defense compared to the existing defense mechanisms explained in §3.2.
Now the effectiveness of SelfDefend depends on whether it can handle all kinds of jailbreak attacks (i.e., O1). For this, we proceed to conduct an empirical study in §5 to validate our earlier hypothesis that “LLMs have the capability to identify harmful portions in user queries.”
5 An Empirical Measurement
In this section, we conduct extensive measurements to empirically show that existing LLMs, notably the commonly used GPT-3.5/4 models, perform acceptably under the SelfDefend framework. This not only validates the hypothesis assumed in §4 but also enables widespread, convenient usage of SelfDefend without the need for further tuning of a defense model. For example, various custom GPTs applications [76] can adopt SelfDefend to immediately enhance their defense capabilities against jailbreaks.
Figure 2: SelfDefend’s two detection prompt templates.
5.1 Initializing SelfDefend for GPT-3.5/4
To facilitate measurements, we need to first initialize SelfDefend for GPT-3.5/4. For GPT-3.5-based experiments, both $LLM_{target}$ in the normal stack and $LLM_{defense}$ in the shadow stack are set to GPT-3.5. Likewise, for GPT-4-based experiments, both $LLM_{target}$ and $LLM_{defense}$ are set to GPT-4.
Prompt Templates. $LLM_{target}$ directly uses user prompts $P_{query}$ but we need to set up the detection prompts for $LLM_{defense}$. As mentioned earlier in §4, we have designed $P_{direct}$ and $P_{intent}$ for two different defense schemes. Figure 2 shows their detailed prompt templates. We can see that $P_{direct}$ aims to directly pinpoint the harmful part from the original jailbreak/normal $P_{query}$; otherwise, it outputs a token of “No.” In contrast, $P_{intent}$ employs a chain-of-thought (CoT) [77] approach to first summarize the true intention of $P_{query}$ and then, based on the summarized intention, further extract any safety-violation part, if present. It also includes two examples in the prompts to help LLMs better understand the instructions and conduct few-shot in-context learning [78].
By comparing these two prompts, we expect that $P_{direct}$ would perform better on optimization-based jailbreaks that directly include the jailbreak portion, while $P_{intent}$ might be more effective against indirect jailbreaks. One additional objective of the empirical measurement in this section is to explore the best setup for deploying SelfDefend with off-the-shelf LLMs under four different combinations (GPT-3.5 vs. GPT-4; $P_{direct}$ vs. $P_{intent}$).
Environment Setup. In our implementation, we use the OpenAI API in Python to access GPT-3.5 and GPT-4. The API versions of GPT-3.5 and GPT-4 are gpt-3.5-turbo-0125 and gpt-4-0125-preview, respectively. We set the hyper-parameters max_tokens, temperature, n, and top_p to 128, 0, 1, and 0, respectively. When accessing the GPT APIs, no system prompt is added.
TABLE II: The details of our collected benchmark datasets.
Dataset # Prompts Jailbreak Methods
JailbreakBench [79] 100 GCG [29], JSAA [31]
PAIR [36], AutoDAN [28]
DrAttack [32], Puzzler [33]
MultiJail [34] 315 MultiJail
AlpacaEval [57] 805 Normal Prompts
TABLE III: The ASR results from testing LLMs against four major categories of jailbreak attacks and normal prompts.
LLMs Optimization-based Template-based Indirect Multilingual Normal
GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 (baseline) 0.56 0.94 0.72 0.90 0.78 0.98 0.39 0.97
GPT-3.5-based Shadow Stack ($P_{direct}$) 0.72 0.98 0.77 0.96 1.00 0.98 0.87 0.97
GPT-3.5-based SelfDefend ($P_{direct}$) 0.45 0.92 0.60 0.87 0.78 0.96 0.36 0.95
GPT-3.5-based Shadow Stack ($P_{intent}$) 0.28 0.06 0.37 0.35 0.86 0.22 0.52 0.99
GPT-3.5-based SelfDefend ($P_{intent}$) 0.19 0.04 0.29 0.31 0.71 0.22 0.20 0.97
Reduction factor 66.07% 95.74% 59.72% 65.55% 8.97% 77.55% 48.71% 0%
GPT-4 (baseline) 0.08 0.16 0.33 0.19 0.74 0.90 0.08 0.97
GPT-4-based Shadow Stack ($P_{direct}$) 0.01 0.00 0.11 0.01 0.05 0.27 0.14 0.96
GPT-4-based SelfDefend ($P_{direct}$) 0.00 0.00 0.10 0.01 0.04 0.26 0.01 0.94
Reduction factor 100% 100% 36.36% 94.73% 94.59% 71.11% 87.50% 3.09%
GPT-4-based Shadow Stack ($P_{intent}$) 0.08 0.05 0.21 0.07 0.13 0.36 0.30 0.99
GPT-4-based SelfDefend ($P_{intent}$) 0.05 0.05 0.18 0.07 0.13 0.28 0.01 0.97
(a) The CDF plot of $\Delta d$ for normal prompts.
(b) Average $\Delta d$ for various jailbreak prompts.
Figure 3: The extra delay $\Delta d$ caused by SelfDefend for normal prompts and jailbreak prompts, respectively.
5.2 Datasets and Attack Setup
After SelfDefend is set up, we collect benchmark datasets, implement the corresponding attack methods, and determine the effectiveness metrics.
Benchmarks. Based on the four categories of existing jailbreak attacks we surveyed in §2 — optimization-based, template-based, indirect, and multilingual jailbreaks — we identify representative jailbreak attack methods in each category. We then collect three benchmark datasets, JailbreakBench [79], MultiJail [34], and AlpacaEval [57], from which we use their user prompts for testing SelfDefend. Table II lists the details of our collected benchmark datasets. Specifically, we use a set of 100 instructions from JailbreakBench [79], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [29] and JSAA [31]), template-based jailbreaks (PAIR [36] and AutoDAN [28]), and indirect jailbreak attacks (DrAttack [32] and Puzzler [33]). In contrast, we directly use the original prompts from the MultiJail and AlpacaEval datasets to construct inputs for the scenarios of multilingual jailbreaks and normal prompts, respectively.
Attack Setup. For GCG, we choose its individual version and optimize the suffix on Vicuna-7b-v1.3 [80] with a batch size of 512 and 500 optimization steps. For JSAA, we query GPT-3.5 with 10,000 iterations and 10 restarts. We perform transfer attacks on GPT-4 using the prompts generated on GPT-3.5. For PAIR, we select Vicuna-13b-v1.5 [80] as the attack model. For AutoDAN, we choose its first version of the genetic algorithm, i.e., AutoDAN-GA. The genetic algorithm used in AutoDAN-GA features a crossover rate of 0.5, a mutation rate of 0.01, and 100 optimization steps. For DrAttack and Puzzler, we use GPT-4 to construct their jailbreak prompts. For MultiJail, we selected all 315 queries in the Bengali language. For AlpacaEval, we select all 805 questions from the AlpacaEval dataset.
Metrics. We measure the defense effectiveness of SelfDefend indirectly using the attack success rate (ASR) [29] or unsafe rate [34]. Following [29, 28], ASR is measured by detecting whether the LLM response contains keywords indicating a refusal to answer, like “I can’t assist.” If such a keyword or its variant is included, it indicates that the attack is unsuccessful; otherwise, it indicates that the jailbreak is successful. We use the common practice [29] for a list of such keywords, which is available in Appendix A. GCG, JSAA, PAIR, AutoDAN, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance. The lower the ASR, the stronger the defense performance of SelfDefend against jailbreaks. Note that we also use ASR to evaluate the performance of SelfDefend under normal prompts from AlpacaEval, but a higher ASR indicates that our framework is more compatible with normal prompts (in such a case, ASR can be interpreted as answer success rate). On the other hand, the unsafe rate uses GPT-4 to determine whether the response of the target model matches the jailbreak goal, following [34]. We adopt the unsafe rate in a way similar to ASK to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.
5.3 Measurement Results
Table III summarizes our core measurement results, which we separate into the defense effectiveness against jailbreaks and the impact on normal user prompts. Additionally, we measure the extra delay introduced by SelfDefend.
Baseline Performance. We first measure the baseline performance of using GPT-3.5/4 only under the tested jailbreaks. As shown in Table III, these jailbreaks are much more effective on GPT-3.5 compared to GPT-4. The ASR for GPT-3.5 can be as high as 0.39 to 0.98 (average: 0.75), whereas the ASR for GPT-4 typically ranges from 0.08 to 0.33, except for indirect jailbreaks in which case GPT-4 also has a high ASR (0.74 for DrAttack and 0.9 for Puzzler).
Defense Effectiveness. We then compare ASR after equipping SelfDefend with the baseline model. In ablation, we also report ASR using only the shadow stack. Our findings are summarized as follows:
  • Based on GPT-3.5, although the SelfDefend frameworks supported by intent prompt and direct prompt both decrease success rates, there is a considerable difference in the enhancement of defense performance achieved by the two methods. The SelfDefend framework with direct prompt shows a modest reduction in terms of ASR due to its internal shadow stack performing even worse than GPT-3.5. For example, its ASR on AutoDAN is 96%, compared to GPT-3.5’s 90%. In contrast, the SelfDefend framework utilizing intent prompt demonstrates significantly better performance than the one with direct prompt, with the reduction factor ranging from as high as 95.74% to as low as 8.97%.
  • Based on GPT-4, both SelfDefend frameworks significantly reduce the ASRs of various jailbreak attacks. Different from the results on GPT-3.5, SelfDefend with direct prompt is better than SelfDefend with intent prompt due to the difference between the two shadow stacks. Although jailbreak attacks present lower ASRs than on GPT-3.5, the reduction factor of SelfDefend with direct prompt ranges from 36.36% to 100.00%.
  • SelfDefend with direct prompt and SelfDefend with intent prompt present the best defense enhance on GPT-4 and GPT-3.5, respectively. Furthermore, SelfDefend with direct prompt demonstrates the higher reduction factor than SelfDefend with intent prompt on GPT-3.5. Thus, we contend that SelfDefend with direct prompt relies more on the model’s own discrimination ability so that is suitable for powerful models such as GPT-4, while the intent prompt can strengthen the model’s logical reasoning and still has decent defense capabilities even on weaker models.
Impact on Normal Queries. Besides defense effectiveness, we also measure the impact of SelfDefend on normal queries by evaluating the ASR of 805 normal prompts from AlpacaEval. Table III shows that for both GPT-3.5 and GPT-4, SelfDefend (with $P_{direct}$) slightly reduces the ASR by 2% (from 0.97 to 0.95) and 3% (from 0.97 to 0.94), respectively. Meanwhile, SelfDefend (with $P_{direct}$) does not decrease the ASR at all for either model. Considering that the ASR under SelfDefend is still very high and close to the baseline, we conclude that SelfDefend’s defense would not notably impact normal user queries.
Extra Delay $\Delta d$. Recall that one of our design objectives is to incur negligible delay to user prompts (O2 in §3.1). We thus further measure the extra delay $\Delta d$ caused by SelfDefend, which is calculated as follows:
[…]
where $d_{total}$ denotes the total delay to SelfDefend-protected LLMs, and $d_{normal}$ represents the separate time cost of target LLMs in the normal stack. The metric of $\Delta d$ is mainly important for normal prompts because it would directly affect the experience of normal users. Figure 3(a) plots the cumulative distribution function (CDF) of $\Delta d$ for normal prompts. We can see that over 95% of the cases incur zero delay across three configurations of SelfDefend, demonstrating that SelfDefend is capable of defending against jailbreaks at the cost of negligible delay for normal user prompts. The only exception occurs with SelfDefend using $P_{intent}$ and GPT-3.5, where around 80% of the cases have no extra delay.
For jailbreak prompts, we also calculate their average $\Delta d$ for different attack scenarios in Figure 3(b). We can see that GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all eight scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval. This is likely because the baseline GPT-4 has good defense performance for most jailbreaks, suggesting its original output is already short and the delay is minimal.
Figure 4: The training procedure for fine-tuning our defense model.
6 Tuning a Dedicated Defense Model
Based on the empirical measurement results, we have observed that GPT-4, when equipped with SelfDefend, significantly reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.06 (for GPT-4-based SelfDefend with $P_{direct}$). However, GPT-4 is commonly known to be expensive [58], and its closed-source nature raises privacy concerns for non-OpenAI model vendors. Therefore, our objective in this section is to tune an open-source model that can be used under the SelfDefend architecture for robust, low-cost, and self-contained jailbreak defense.
6.1 Design Overview
Given the powerful defensive capability of GPT-4-based SelfDefend, our intuition is to “transfer” this capability to an open-source model. To do so, we leverage GPT-4-based SelfDefend to distill and generate high-quality tuning data. Figure 4 depicts our proposed pipeline to fine-tune an open-source defense model. Specifically, by continuously incorporating harmful and harmless prompts into our defense prompts (i.e., $P_{direct}$ or $P_{intent}$) as inputs for GPT-4, we gather their outputs as labels for these samples. Since we utilize two defense prompts, we eventually obtain two separate datasets, which we then use to fine-tune the employed open-source model for our defense purposes.
6.2 Data Distillation
To distill GPT-4’s “knowledge” on safety policies, a dataset containing both harmful and harmless queries is necessary. We utilize the red-team data from Anthropic [59] as the query dataset. This dataset consists of 38,961 text transcripts documenting conversations between human adversaries and AI assistants. From this dataset, we select the initial human prompt and exclude the corresponding assistant response, omitting any subsequent exchanges, to create a single-turn prompt dataset designated as $\mathcal{D}_{red}$. It is worth noting that this dataset is released in 2022 and has no overlap with JailbreakBench/MultiJail released in 2024. Furthermore, this dataset is exclusively in English, implying that a model trained on this data will not acquire any additional multilingual capabilities.
Next, we employ GPT-4 and our defense prompts to distill GPT-4’s “knowledge” on individual queries from $\mathcal{D}_{red}$. This procedure can be formalized in Equation 2. Note that, for simplicity, we refer to the symbols $P_{direct}$ and $P_{intent}$ as $P_{dir}$ and $P_{int}$, respectively. Their corresponding datasets are denoted as $D_{dir}$ and $D_{int}$, respectively.
$\mathcal{D}_{dir}$ =$=\{(x,y)|x\in\mathcal{D}_{red}\},\quad y=\operatorname{GPT_{4}}(P_{dir}[x])$ (2)
$\mathcal{D}_{int}$ =$=\{(x,y)|x\in\mathcal{D}_{red}\},\quad y=\operatorname{GPT_{4}}(P_{int}[x]),$
Here, $x$ denotes an original prompt from $\mathcal{D}_{red}$ and $y$ represents the corresponding response from GPT-4 under either $P_{dir}$ or $P_{int}$. By analyzing all $x\in\mathcal{D}_{red}$, we eventually obtain two distilled datasets, $\mathcal{D}_{dir}$ and $\mathcal{D}_{int}$, for tuning.
6.3 LoRA Fine-Tuning
To fine-tune defense models with distilled datasets, we choose the publicly available Llama 2 [61] and its 7b model to demonstrate that an open-source, low-cost, yet robust defense model can be trained. Given the limited GPU resources, we utilize a parameter-efficient fine-tuning (PEFT) method known as LoRA [60] to tailor the pre-trained Llama-2-7b for dedicated jailbreak defense. LoRA is designed to adapt large pre-trained language models with minimal computational overhead. Despite the reduction in trainable parameters, LoRA maintains competitive performance, often matching or exceeding the results of full fine-tuning approaches [60].
We allocate 80% of the samples in $\mathcal{D}_{dir}$ or $\mathcal{D}_{int}$ for fine-tuning and reserve the remaining 20% as validation sets. During the training process, we continue to use our defense prompts and queries from the collected datasets as inputs; that is, we take $P_{dir}(x)$ or $P_{int}(x)$ as input and $y$ as the label to fine-tune the Llama model.
Formally, the fine-tuning objective for the direct prompt can be formulated as follows:
[…]
where $\Theta$ represents the trainable parameters of LoRA, $\Phi_{0}$ denotes the pre-trained weights of the Llama model, and $\Delta\Phi(\Theta)$ indicates the parameter increment determined by LoRA. Fine-tuning the Llama model with this objective can enhance its ability to directly detect jailbreak portions in query prompts.
Likewise, the fine-tuning objective for the intent prompt can be written as follows:
$\max_{\Theta}\sum_{(x,y)\in\mathcal{D}_{int}}\sum_{t=1}^{|y|}\log\left(p_{\Phi_{0}+\Delta\Phi(\Theta)}(y_{t}\mid P_{int}[x],y_{<t})\right)$ (4)
This objective allows the fine-tuned model to identify jailbreak intentions within query prompts.
Implementation. We use LoRA to fine tune the Llama model with $r$ of 8 and $\alpha$ of 32 [60]. The initial learning rate is $10^{-3}$. The training epochs are 1 in batch size of 8.
7 Evaluation
TABLE IV: Jailbreak ASR for various defense methods.
Target Model Defense Method Optimization-based Template-based Indirect Multilingual Normal
GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 GPT-3.5 (baseline) 0.56 0.94 0.72 0.90 0.78 0.98 0.39 0.97
Perplexity Filter [62] 0.03 1.00 1.00 1.00 1.00 1.00 1.00 0.99
SmoothLLM [51] 0.48 0.89 0.65 0.93 0.85 0.99 1.00 0.96
Llama Guard [42] 0.41 0.61 0.43 1.00 0.97 0.93 0.95 0.99
Llama Guard 2 [63] 0.14 0.50 0.37 0.15 0.89 0.64 0.55 0.99
$P_{direct}$-tuned Shadow Stack 0.08 0.11 0.14 0.07 0.78 0.07 0.74 0.96
$P_{direct}$-tuned SelfDefend 0.06 0.11 0.07 0.07 0.62 0.07 0.30 0.94
$P_{intent}$-tuned Shadow Stack 0.08 0.05 0.16 0.09 0.18 0.20 0.57 0.99
$P_{intent}$-tuned SelfDefend 0.05 0.05 0.12 0.08 0.16 0.20 0.26 0.97
Double shadow stack 0.04 0.02 0.10 0.05 0.18 0.07 0.47 0.97
Double shadow stack + GPT-3.5 0.03 0.02 0.06 0.05 0.16 0.07 0.19 0.95
GPT-4 GPT-4 (baseline) 0.08 0.16 0.33 0.19 0.74 0.90 0.08 0.97
Perplexity Filter [62] 0.03 1.00 1.00 1.00 1.00 1.00 1.00 0.99
SmoothLLM [51] 0.07 0.19 0.33 0.18 0.91 0.88 1.00 0.97
Llama Guard [42] 0.41 0.61 0.43 1.00 0.98 0.93 0.95 0.99
Llama Guard 2 [63] 0.14 0.50 0.35 0.15 0.91 0.64 0.55 0.99
$P_{direct}$-tuned Shadow Stack 0.04 0.11 0.17 0.07 0.71 0.12 0.71 0.96
$P_{direct}$-tuned SelfDefend 0.01 0.07 0.15 0.05 0.58 0.07 0.06 0.94
$P_{intent}$-tuned Shadow Stack 0.08 0.03 0.19 0.07 0.18 0.19 0.56 0.99
$P_{intent}$-tuned SelfDefend 0.04 0.02 0.19 0.06 0.17 0.14 0.04 0.97
Double shadow stack 0.02 0.02 0.12 0.04 0.18 0.12 0.47 0.97
Double shadow stack + GPT-4 0.01 0.02 0.12 0.04 0.17 0.07 0.04 0.95
Llama 2 Llama-2-7b-chat (baseline) 0.57 0.96 0.59 0.68 0.88 0.99 0.14 0.98
Perplexity Filter [62] 0.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99
SmoothLLM [51] 0.93 1.00 0.82 0.95 0.98 1.00 0.99 0.99
Llama Guard [42] 0.40 0.41 0.46 1.00 0.84 0.93 0.95 0.99
Llama Guard 2 [63] 0.17 0.53 0.41 0.15 0.89 0.64 0.55 0.99
(7b-chat) $P_{direct}$-tuned Shadow Stack 0.06 0.06 0.25 0.05 0.36 0.09 0.73 0.97
$P_{direct}$-tuned SelfDefend 0.04 0.06 0.22 0.05 0.31 0.09 0.10 0.95
$P_{intent}$-tuned Shadow Stack 0.11 0.02 0.24 0.08 0.15 0.22 0.58 0.99
$P_{intent}$-tuned SelfDefend 0.09 0.02 0.21 0.07 0.15 0.22 0.06 0.98
Double shadow stack 0.05 0.02 0.18 0.04 0.12 0.04 0.48 0.97
Double shadow stack + Llama 2 0.03 0.02 0.16 0.04 0.11 0.04 0.06 0.95
In this section, we extensively evaluate SelfDefend with Llama-2-tuned defense models and compare their performances with other jailbreak defense methods. We generally follow the setup mentioned in §5 unless explicitly specified in §7.1. For each evaluation target, we assess not only the defense effectiveness (in §7.2) and the extra delay $\Delta d$ (in §7.3) as in §5, but also whether the detected harmful portion actually aligns with the original prompt (i.e., explainability in §7.4), and how robust our defense models are against targeted optimization-based attacks (in §7.5) and prompt injection (in §7.6).
7.1 Experimental Setup
Target LLMs, Benchmarks, and Environment. As shown in Table IV, our tuned defense models are paired with protect three popular LLMs: GPT-3.5, GPT-4, and the open-source Llama-2-7b-chat [61]. We use the same benchmark datasets as in Table II (see §5). We follow the same procedure and configuration as in §5 to generate the tested jailbreak prompts for GPT-3.5 and GPT-4. The process is similar for testing the target model Llama-2-7b-chat; we simply set the target model for query-based JSAA, PAIR, and DrAttack jailbreaks to Llama-2-7b-chat to generate their jailbreak prompts. Our evaluation is implemented using PyTorch 2.2.1 and conducted on an NVIDIA TESLA H800 GPU.
Baselines. We compare our framework with popular jailbreak defense methods, including Perplexity Filter [62], SmoothLLM [51], Llama Guard [42], and Llama Guard 2 [63]. Specifically, Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt. A jailbreak is considered to happen when the perplexity exceeds a threshold. We set this threshold at the maximum perplexity of any prompt in the JailbreakBench dataset of harmful behavior prompts. SmoothLLM perturbs the jailbreak prompts with character-level changes to enable the target LLM to perform defense. In this paper, we set SmoothLLM to perform character swapping with a 10% perturbation percentage. Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts. Llama Guard 2 is a fine-tuned version of Llama-3-8B [81].
Metric. We measure the performance of defense methods by the attack success rate (ASR), i.e., the frequency with which jailbreak prompts in a benchmark dataset bypass the detection of a defense method. The lower the ASR, the stronger the defense performance. However, the definition of “attack success” is different for different defense methods. For Perplexity Filter, the success denotes that the perplexity of the query text is no more than the fixed threshold. For SmoothLLM, the definition of success is the success of the attack method, i.e., ASR [29] or Unsafe Rate [34]. It is worth noting that the Unsafe Rate refers to the unsafety of the target model’s response, rather than the input prompt. For Llama Guard, it is success when its response is “unsafe”. For our fine-tuned models, the success represent their response being “No.”
Figure 5: An example of tuned models against GCG on GPT-3.5.
Figure 6: An example of tuned models against PAIR on GPT-3.5.
Figure 7: An example of tuned models against DrAttack on GPT-3.5.
Figure 8: An example of tuned models against MultiJail on GPT-3.5.
Figure 9: Tuned models’ response to a normal prompt from AlpacaEval.
7.2 Defense Effectiveness
We now discuss the defense effectiveness of our proposed technique. In line with Table IV, we discuss the results over GPT-3.5, GPT-4, Llama-2-7b-chat, and a “double shadow stack” setting, respectively.
Results on GPT-3.5. Table IV reports ASRs of multiple defense methods under various types of jailbreak attacks. The defense methods are evaluated across different jailbreak techniques, consisting of optimization-based, template-based, indirect, multilingual, and normal prompts. Under the optimization-based GCG and JSAA, ASR of SelfDefend with the intent prompt is 0.05, slightly better than that with the direct prompt. In contrast, SelfDefend with the direct prompt is slightly better than SelfDefend with the intent prompt under the template-based PAIR and AutoDAN.
For indirect DrAttack, ASR of the fine-tuned model with direct prompt is 0.78, the same as the original GPT-3.5, whereas the dedicated model with intent prompt reduces ASR to 0.18. With manual exploration, we find that DrAttack splits the sentence components of the original jailbreak prompt and replaces them with harmless words, making it difficult for SelfDefend (direct prompt) to extract harmful content. Although Puzzler is also an indirect attack, the prompts it produces contain offensive content as implicit clues for jailbreak, resulting in a lower ASR of SelfDefend with the direct prompt. For multilingual MultiJail, the intent prompt tuned model is superior than the direct prompt by 17%. Therefore, the model fine-tuned with the direct prompt is well-suited for identifying overt jailbreak intentions within sentences, whereas the model fine-tuned with the intent prompt excels in detecting jailbreak attempts that encompass more subtle or concealed intentions. Besides, SelfDefend has ASRs that are comparable with all other defense methods under normal prompts, indicating its compatibility with normal users.
Figures 5, 6, 7, 8 and 9 present several example responses of the tuned models toward either jailbreak prompts or a normal query. Note that in Figure 7, DrAttack evades the detection of the direct prompt tuned model, while the intent tuned model accurately identifies the intent of DrAttack, which further illustrates the weakness of the direct tuned model against indirect attacks. Moreover, as expected, we observe that both tuned models make satisfactory decisions on other jailbreaks or the normal prompt.
Results on GPT-4. Similar with the results achieved over GPT-3.5, we can observe that our defense frameworks achieve the lowest ASR under all jailbreak methods, and is still close to ASR of GPT-4 on AlpacaEval. We interpret the results as generally encouraging, illustrating the defensive superiority of SelfDefend without intruding normal user experience. Under the optimization-based GCG, the lowest ASR (as achieved by SelfDefend) is 0.01 with the direct prompt. For JSAA, SelfDefend with the intent prompt is slightly better than SelfDefend with the direct prompt. Under the template-based PAIR and AutoDAN, SelfDefend with the direct prompt is slightly superior than that with the intent prompt. For indirect DrAttack, ASR of the fine-tuned model with direct prompt is 0.71, close to that of the original GPT-4, while the fine-tuned model with intent prompt reduces ASR to 0.18. In addition, the direct prompt tuned model performs the lowest ASR under Puzzler, but far inferior to the intent one under MultiJail.
Results on Llama-2-7b-chat. Our defense frameworks exhibit the lowest ASRs across all jailbreak methods. Moreover, it achieves a ASR that is comparable to Llama-2-7b-chat on AlpacaEval, showing SelfDefend’s defensive effectiveness without compromising normal user experience. Specifically, under the white-box GCG, the ASRs for both of our SelfDefend frameworks are less than 10%.
Perplexity Filter exhibits the lowest ASR against GCG; with further exploration, we find that this is because its filter threshold is set to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix. Yet, its detectability in other settings are not as good as SelfDefend. Under the optimization-based JSAA attack, SelfDefend with the intent prompt achieves the lowest ASR of 0.02, performing marginally better than with the direct prompt. In the context of template-based attacks such as PAIR and AutoDAN, the results between SelfDefend with the direct prompt and the intent one are close, with each having their own superior cases. For the indirect DrAttack, while the direct prompt tuned model largely reduces the ASR compared to the original Llama-2-7b-chat model, it still has a 21% gap compared to the intent one. Additionally, the direct prompt fine-tuned model has the lowest ASR under the Puzzler attack, yet it is notably outperformed by the intent prompt one in the MultiJail scenario.
Double Shadow Stacks. The current design of SelfDefend  accompanies one shadow stack with one LLM under protection. Despite the encouraging protection results, one may wonder incorporating multiple shadow stacks to harvest increased defense capability. Here, we define a “double shadow stack” setting as two shadow stacks independently analyze an input, where an input is deemed jailbreak when at least one shadow stack flags it as jailbreak.
In Table IV, “Double shadow stack” is instantiated by incorporating one $P_{direct}$-based and one $P_{intent}$-based shadow stacks. We can observe that the results of double shadow stack are close to the best results of the two SelfDefend frameworks, except under MultiJail. And when we further fuse the target model (i.e., normal stack) with the double shadow stack, it exhibits the lowest ASRs under all attack scenarios, as shown in Table IV. Nevertheless, we point out that when fusing two or three stacks together, the performance over normal prompts (AlpacaEval) decrease, meaning a few more normal inputs are deemed harmful. This is expected; taking the logical disjunction of multiple stacks’ predictions reduces the false negatives of SelfDefend, yet likely increases the false positives. In practice, we suggest using this paradigm only when the standard form of SelfDefend offers low detection accuracy.
Figure 10: Average $\Delta d$ introduced by the tuned models on Llama-2-7b-chat.
TABLE V: The CLIP-Score [64, 82, 83] ($\uparrow$ indicates better) based on Table IV’s results for Llama 2.
Tuned Model Text Optimization-based Template-based Indirect Multilingual
GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail
$P_{direct}$-tuned model Generated Attack Prompts 0.85 0.70 0.83 0.78 0.74 0.66 0.68
Identified Harmful Prompts 0.95 0.94 0.90 0.94 0.79 0.73 0.69
$P_{intent}$-tuned model Generated Attack Prompts 0.85 0.70 0.83 0.78 0.73 0.67 0.67
Identified Harmful Intentions 0.88 0.88 0.85 0.88 0.83 0.77 0.74
7.3 Extra Delay $\Delta d$
Similar to how we measured the extra delay $\Delta d$ introduced by GPT-based SelfDefend in §5.3, we now report the average delay $\Delta d$ introduced by the tuned models on Llama-2-7b-chat in Figure 10. Compared to earlier results in Figure 3, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.01 seconds. In contrast, GPT-3.5-based (GPT-4-based) SelfDefend with $P_{intent}$ incurred an average $\Delta d$ of around 0.07 seconds (0.02 seconds) under AlpacaEval. This is likely because the fine-tuned models enjoy faster inference speeds when dealing with similar queries. Besides the normal prompts, the tuned models also significantly reduce the extra delay $\Delta d$ under various jailbreak prompts. For example, the maximum $\Delta d$ now decreases from the earlier 1.1 seconds on GPT-4 to 0.25 seconds. Indeed, except for JSAA, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend. These results indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for real-world deployment.
7.4 Explainability
Figure 5, 6, 7, and 8 demonstrate the efficacy of the tuned models in identifying harmful content within jailbreak prompts. Here, we further quantitatively assess whether the identified harmful content actually aligns with the original jailbreak prompt. We measure their semantic similarity using the ensemble CLIP-score [64, 82, 83] as shown in Table V. Note that our measurement exclusively targets the examples where the defense mechanism is successful.
We calculate the CLIP-score between the original prompts $P_{query}$ from JailbreakBench and the attack-generated jailbreak prompts, as well as the CLIP-score between $P_{query}$ and our identified harmful content (prompts/intentions). Specifically, the CLIP-scores for identified harmful portions are higher than those for the generated attack prompts, suggesting that the jailbreak prompts have altered the original prompts, whereas the harmful portions identified by our models are more aligned with the original prompts. Regarding the differences between the two tuned models, we observe that the similarity of identified harmful prompts is higher than that of harmful intentions in the context of optimization-based and template-based attacks, but lower than that of intentions under indirect and multilingual attacks. This can be attributed to the fact that optimization-based and template-based attack prompts contain the original prompts, whereas indirect and multilingual attack prompts significantly distort the original prompts. In summary, these findings are consistent with our initial design intentions and demonstrate that the harmful portions identified by our models can indeed provide an explainable foundation for defense mechanisms.
TABLE VI: ASR with the regular GCG attack on Llama 2 and the targeted GCG directly on our tuned defense models.
Tuned Model Regular GCG Targeted GCG
(from Table IV) (white box)
$P_{direct}$-tuned model 0.06 0.07
$P_{intent}$-tuned model 0.11 0.09
7.5 Robustness to Targeted GCG Attacks
In earlier evaluation with optimization-based attacks such as GCG in §7.2, we used a regular setting where such attacks leverage gradients to optimize adversarial suffixes on an open-source substitute model (i.e., Llama-2-7b-chat) and transfers these suffixes to attack our fine-tuned defense models. To stress-test the robustness of our fine-tuned defense models against targeted GCG attacks, we give the adversary full white-box access to our model parameters and allow them to optimize adversarial suffixes directly on our defense models.
Table VI presents the ASR results for both $P_{direct}$-tuned model and $P_{intent}$-tuned model. We observe that the ASR results of the targeted, white-box GCG remain below 10% for both types of tuned models in the shadow stack. ASR slightly increases for the targeted GCG under the $P_{direct}$-tuned model but slightly decreases for the targeted GCG under the $P_{intent}$-tuned model. This indicates that our defense models are robust to even targeted GCG attacks.
7.6 Robustness to Prompt Injection
Since SelfDefend uses $P_{direct}$ or $P_{intent}$ to wrap the original $P_{query}$, it is natural to wonder whether SelfDefend is robust to prompt injection [65, 66], another major LLM threat that is orthogonal to jailbreak attacks.
To measure the robustness of SelfDefend’s fine-tuned models against prompt injection, we leverage HouYi [66], an effective black-box prompt injection methodology that has been tested on various LLM-integrated applications. In our evaluation, we assume that HouYi [66] can directly access the outputs of our tuned defense models. HouYi’s goal is to search for a distraction prompt in an input of the form “[Jailbreak Prompt] [Distraction Prompt]” so that our defense models tend to answer with “No.” We use 100 original jailbreak goals defined in JailbreakBench as “[Jailbreak Prompt]” to conduct this testing. The ASR results for the direct and intention prompt-tuned models are 0.21 and 0.17, respectively. We consider these ASR values low for both tuned models, indicating their robustness against distractions from prompt injection. Moreover, we notice that the intent prompt-tuned model is less affected by prompt injection than the direct prompt-tuned model, as evidenced by its lower ASR.
8 Conclusion
We have introduced SelfDefend, a robust, low-cost, and self-contained defense against LLM jailbreak attacks. Inspired by the concept of shadow stacks, SelfDefend delivers a dual-layer defense mechanism comprising a shadow LLM that guards the target LLM. It further leverages a tuning-based approach to enhancing the shadow LLM’s defense capability. Evaluation shows that SelfDefend is lightweight and effective to mitigate a wide spectrum of jailbreak attacks while rarely undermining normal queries.
Appendix A Keywords
TABLE VII: The refusal signals considered for GPT-3.5/4.
“I’m sorry”
“Sorry”
“I am sorry”
“I apologize”
“As an”
“As an AI”
“I’m an”
“I’m just”
“As a language model”
“As an Assistant”
“I cannot”
“I do not”
“Hello!”
“is illegal and unethical”
“I’m really sorry”
“OpenAI” “unethical”
“not ethical”
“illegal”
“not legal”
“My apologies”
“but I cannot”
“If you have any other non-malicious requests”
“I’m not able to provide”
“I am really sorry”
“I cannot create”
“I cannot support”
“I can’t provide”
“I can’t assist”
TABLE VIII: The refusal signals considered for Llama 2.
“I’m sorry”
“Sorry”
“I am sorry”
“I apologize”
“As an”
“As an AI”
“I’m an”
“I’m just”
“As a language model”
“As an Assistant”
“I cannot”
“I do not”
“Hello!”
As mentioned in §5.2, we use a list of refusal keywords adopted by common practice [29]. They are listed in Table VII for GPT-3.5/5 and in Table VIII for Llama 2, respectively.
Appendix B More Examples from Our Tuned Models
More examples from our tuned defense models are presented in Figure 1112, and 13.
Figure 11: An example of tuned models against JSAA on GPT-3.5.
Figure 12: An example of tuned models against AutoDAN on GPT-3.5.
Figure 13: An example of tuned models against Puzzler on GPT-3.5.

### v2

@@
 Abstract
-Jailbreaking is an emerging adversarial attack that bypasses the safety alignment deployed in off-the-shelf large language models (LLMs) and has evolved into four major categories: optimization-based attacks such as the Greedy Coordinate Gradient (GCG) attack, jailbreak template-based attacks such as “Do-Anything-Now” (DAN), advanced indirect attacks like DrAttack, and multilingual jailbreaks.
-However, delivering a practical jailbreak defense is challenging because it needs to not only handle all the above jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs.
-Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM defense instance to concurrently protect the target LLM instance in the normal stack and collaborate with it for checkpoint-based access control.
+Jailbreaking is an emerging adversarial attack that bypasses the safety alignment deployed in off-the-shelf large language models (LLMs) and has evolved into multiple categories: human-based, optimization-based, generation-based, and the recent indirect and multilingual jailbreaks.
+However, delivering a practical jailbreak defense is challenging because it needs to not only handle all the above jailbreak attacks but also incur negligible delays to user prompts, as well as be compatible with both open-source and closed-source LLMs.
+Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM as a defense instance to concurrently protect the target LLM instance in the normal stack and collaborate with it for checkpoint-based access control.
 The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful prompts or intentions in user queries, which we empirically validate using the commonly used GPT-3.5/4 models across all major jailbreak attacks.
-Our measurements show that SelfDefend enables GPT-3.5 to suppress the attack success rate (ASR) by 8.97-95.74% (average: 60%) and GPT-4 by even 36.36-100% (average: 83%), while incurring negligible effects on normal queries.
 To further improve the defense’s robustness and minimize costs, we employ a data distillation approach to tune dedicated open-source defense models.
-These models outperform four SOTA defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays.
-We also empirically show that the tuned models are robust to targeted GCG and prompt injection attacks.
+These models outperform six state-of-the-art defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays.
+We also empirically show that the tuned models are robust to adaptive jailbreaks and prompt injections.
 1 Introduction
@@
 To avoid causing social anxiety, ethical, and legal issues due to LLM responses to harmful questions, LLM vendors typically conduct safety alignment to prevent the misuse of LLMs through techniques like RLHF (Reinforcement Learning from Human Feedback) [30].
-In response to a harmful prompt that violates safety policies, an aligned LLM often replies with a standard response such as “I’m sorry, I can’t assist with that request.” To bypass LLMs’ safety alignment checks, an adversarial attack known as jailbreaking [23] was proposed.
+In response to a harmful prompt that violates safety policies, an aligned LLM often replies with a standard response such as “I’m sorry, I can’t assist with that request.” To bypass LLMs’ safety alignment, an adversarial attack known as jailbreaking [70] was proposed.
 In the past two years, research on LLM jailbreak attacks and defenses has attracted considerable interest, with most of them focused on the offensive side.
-Specifically, jailbreak strategies have evolved from manual prompt engineering [23, 24, 25] to automatic LLM-based red teaming [26, 27, 28].
-Besides these template-based jailbreaks aimed at identifying a valid jailbreak prompt template, a more generic, optimization-based adversarial jailbreak approach, notably Greedy Coordinate Gradient (GCG) [29], was recently proposed.
-It uses public available models to train adversarial suffixes that maximize the probability for LLMs to produce an affirmative response rather than refusing to answer.
-These studies [29, 30, 31] found that the identified suffixes are transferable to closed-source off-the-shelf LLMs.
-More recently, advanced indirect jailbreaks like DrAttack [32] and Puzzler [33], as well as multilingual jailbreaks [34, 35], have also been invented.
-In addition to proposing these four major types of jailbreaks, various benchmark studies on LLM jailbreak attacks [23, 36, 37, 38, 39, 40] have also been conducted.
-On the contrary, the defensive side is somewhat overshadowed, despite around a dozen defense mechanisms being proposed in the past year.
-They can be roughly categorized into tuning-based and non-tuning-based mechanisms.
-Specifically, tuning-based defenses [41, 42, 43, 44, 45, 46] aim to fundamentally improve a model’s safety alignment against jailbreaking, while non-tuning-based defenses [47, 48, 49, 50, 51, 52, 53, 54, 55] can be typically applied to any off-the-shelf LLMs.
+Jailbreak strategies have evolved from manual prompt engineering [70, 43, 73, 62] to automatic LLM-based red teaming [54, 16].
+Besides these human-crafted and generative jailbreaks aimed at identifying a valid jailbreak prompt, a more generic, optimization-based adversarial jailbreak approach, notably Greedy Coordinate Gradient (GCG) [99], was proposed.
+It learns adversarial suffixes on public-available models to maximize their probability in producing an affirmative response instead of refusing, which can be transferable to closed-source off-the-shelf LLMs.
+Recently, advanced indirect jailbreaks like DrAttack [33] and Puzzler [10], as well as multilingual jailbreaks [17, 61], have also been invented.
+In addition to proposing new attacks, various benchmark studies on LLM jailbreak attacks [46, 32, 14, 86] have also been conducted.
+On the contrary, the defensive side is somewhat overshadowed, despite over a dozen defense mechanisms being proposed in the past year.
+They can be roughly categorized into model-based and plugin-based mechanisms.
+Specifically, model-based defenses [35, 89, 96, 81, 93, 49, 82, 45, 77, 91, 39] aim to fundamentally improve a model’s robustness against jailbreaking, while plugin-based defenses [4, 26, 55, 29, 9, 59, 27, 73, 80, 88, 25, 79, 23, 98] can be typically plugged into any off-the-shelf LLMs.
 We conduct an analysis of these defense techniques in §3 and find that it is still challenging for them to be widely used in practice.
-In short, we advocate that a practical jailbreak defense needs to not only handle all the aforementioned jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs in various scenarios (see details in §3).
-In this paper, we propose a new perspective on practically defending against jailbreak attacks, inspired by how the traditional security concept of shadow stacks [56] defends against memory overflow attacks.
+In short, we advocate that a practical jailbreak defense needs to not only handle all the aforementioned jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs.
+In this paper, we propose a new perspective on defending jailbreak attacks, inspired by how the traditional security concept of shadow stacks [8] defends against memory overflow attacks.
 Similar to the shadow stack creating a shadow memory space, we establish a shadow LLM defense instance, $LLM_{defense}$, alongside the target LLM instance, $LLM_{target}$, in the normal stack.
 Under this framework, $LLM_{target}$ can process any user prompt query $P_{query}$ as usual to produce a token-by-token output.
-Meanwhile, $LLM_{defense}$ employs a dedicated detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$).
+Meanwhile, $LLM_{defense}$ employs a tailored detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$).
 Such a unique setup brings several benefits.
-① It simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate due to this dual-layer protection.
-② Since $LLM_{defense}$’s output is typically short, such as only “No” (indicating no issue) for normal queries, a checkpoint in the normal stack tends to be quickly triggered from the shadow stack without delaying $LLM_{target}$’s output.
+① It simultaneously utilizes both $LLM_{target}$’s safety alignment and $LLM_{defense}$’s jailbreak detection, largely increasing the defense success rate due to this dual-layer protection.
+② As $LLM_{defense}$’s output is typically short, such as only “No” (indicating no issue) for normal queries, a checkpoint in the normal stack tends to be quickly triggered from the shadow stack without delaying $LLM_{target}$’s output.
 ③ Since $LLM_{defense}$ does not need to modify or monitor $LLM_{target}$’s internals, it can protect both open-source and closed-source LLMs.
@@
 The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful portions (prompts/intentions) in user queries.
-To validate this hypothesis, we conduct an empirical measurement in §5 using the commonly used GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [29, 31, 36, 28, 32, 33, 34].
+To validate this hypothesis, we conduct an empirical measurement in §5 using the commonly used GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [62, 99, 40, 12, 47, 33, 10, 17].
 The results are quite promising in that SelfDefend enables both GPT-3.5 and GPT-4 to significantly suppress the attack success rate (ASR).
-Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97-95.74% (average: 60%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.28, and GPT-4-based SelfDefend even reduces the ASR by 36.36-100% (average: 83%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.06.
-Besides the jailbreak scenarios, we also test SelfDefend against 805 normal prompts from the AlpacaEval dataset [57] and find that the pass rate is not affected for GPT-3.5 and slightly decreases by 3% for GPT-4, indicating that SelfDefend incurs negligible effects on normal queries.
-Moreover, SelfDefend incurs zero delay for over 95% of normal prompts across three out of four configurations, averaging 0.08 seconds for GPT-3.5 and 0.04 seconds for GPT-4.
-In various attack scenarios, SelfDefend causes an overall average extra delay of 0.04 seconds for GPT-3.5 and 0.39 seconds for GPT-4, respectively.
-While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.06, GPT-4 itself is commonly known to be expensive [58].
+Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97-97.26% (average: 65.06%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.244, and GPT-4-based SelfDefend even reduces the ASR by 69.69-100% (average: 85.53%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.063.
+Besides the jailbreak scenarios, we also test SelfDefend against 805 normal prompts from the AlpacaEval dataset [34] and find that the pass rate is almost unaffected for GPT-3.5 and slightly decreases by 2.77% for GPT-4, indicating that SelfDefend incurs negligible effects on normal queries.
+Moreover, SelfDefend incurs zero delay for over 95% of normal prompts across three out of four configurations.
+For multiple jailbreak samples, SelfDefend causes an average extra delay of 0.05 seconds for GPT-3.5 and 0.40 seconds for GPT-4, respectively.
+While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.063, GPT-4 itself is commonly known to be expensive [3].
 Moreover, the closed-source nature of GPT-3.5/4 raises privacy concerns for non-OpenAI model vendors.
 Therefore, we attempt to tune an open-source model that can be used under the SelfDefend framework for robust, low-cost, and self-contained jailbreak defense.
-By conducting GPT-4-based data distillation (with SelfDefend’s prompts) on a red-team dataset from Anthropic [59] comprising 38,961 harmful and harmless prompts, we generate a large set of high-quality tuning data.
-The distilled datasets are then used to tune our defense models through the LoRA fine-tuning technique [60] on the publicly available Llama-2-7b model [61].
+By conducting GPT-4-based data distillation (with SelfDefend’s prompts) on a red-team dataset from Anthropic [19] comprising 38,961 harmful and harmless prompts, we generate a large set of high-quality tuning data, which is then used to tune our defense models through LoRA fine-tuning [22] on the publicly available Llama-2-7b model [67].
 The details will be introduced in §6.
-To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with four representative jailbreak defenses: Perplexity Filter [62], SmoothLLM [51], Llama Guard [42], and Llama Guard 2 [63].
-The results show that SelfDefend consistently outperforms all other defenses in 19 out of 21 tested attack scenarios (seven jailbreak methods and three target LLMs: GPT-3.5, GPT-4, and Llama 2) and reaches the level of GPT-4-based SelfDefend’s defense.
-The only two exceptions occurred in comparison with Perplexity Filter [62], which is specifically designed to defend against the GCG attack and offers no defense against other types of jailbreaks, particularly for the GCG attack targeting GPT-3.5 and Llama 2 LLMs.
-Nevertheless, we still outperform it when the target LLM is GPT-4.
+To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with six representative jailbreak defenses: ICD [73], SafeDecoding [82], Perplexity Filter [26], SmoothLLM [59], Llama Guard [25], and Llama Guard 2 [66].
+The results show that SelfDefend consistently outperforms all other defenses in 20 out of 24 tested attack scenarios (involving eight jailbreak methods and three target LLMs: GPT-3.5, GPT-4, and Llama-2) and reaches the level of GPT-4-based SelfDefend’s defense.
 Besides defense effectiveness, we also measure the extra delay $\Delta d$ caused by our defense models and find that the average $\Delta d$ is negligible, at 0-0.01 seconds for normal prompts.
-For attack scenarios, the maximum $\Delta d$ has decreased from 1.1 seconds in GPT-4-based SelfDefend to 0.25 seconds, with $\Delta d$ in all attack scenarios except for JSAA now below 0.1 seconds.
-These findings indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for real-world deployment.
-Furthermore, we specifically assess whether the detected harmful portion actually aligns with the original prompt through the ensemble CLIP-score [64], and empirically show that the tuned models are robust to targeted GCG and prompt injection [65, 66] attacks that have white-box access to our defense models.
-More details are available in §7.
-Availability and Ethics.
-To facilitate future research, we are on the way of releasing our datasets, raw evaluation results, and code at this anonymous GitHub link: https://github.com/SelfDefend.
-While our study required testing various jailbreak attacks, these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server.
-We also tried to minimize the overhead to OpenAI by spreading all the experiments over a period of two months instead of just several days and paid around 1,500 USD to OpenAI for our API usage.
+For attack scenarios, the maximum $\Delta d$ has decreased from 1.56 seconds in GPT-4-based SelfDefend to 0.27 seconds, with $\Delta d$ in all attack scenarios except for DAN now below 0.1 seconds.
+These findings indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it efficient for potential real-world deployment.
+Furthermore, we specifically assess whether the detected harmful portion actually aligns with the original prompt through the ensemble CLIP-score [58], and empirically show that the tuned models are robust to adaptive attacks and prompt injections [44, 42].
+Details are available in §7.
+The contributions of this paper are summarized as follows:
+  • We creatively apply the traditional system security concept of shadow stacks to practical LLM jailbreak defense, and our SelfDefend framework utilizes LLMs in both normal and shadow stacks for dual-layer protection.
+  • We successfully initialize SelfDefend for GPT-3.5/4 with two carefully designed detection prompts and empirically validate that LLMs can identify harmful portions (prompts/intentions) in user queries using our measures.
+  • We further fine-tune dedicated open-source models that can be used under the SelfDefend architecture for robust, low-cost, and self-contained jailbreak defense.
 2 Background
@@
 Multimodal jailbreaks [56, 6, 76, 51] are outside the scope of this paper.
-The objective of jailbreaking is to bypass the LLM’s safety mechanisms and induce it to generate harmful, toxic, or objectionable content.
-In this context, we assume the adversary has access to the LLM and can input arbitrary prompts to it.
+The objective of jailbreaking is to circumvent the LLM’s safety alignment and induce it to generate harmful, toxic, or objectionable content.
+In this context, we assume the adversary can access the target LLM’s interface and input arbitrary prompts to it.
 Given the vocabulary $\mathcal{T}$, the LLM, denoted as $\text{LLM}:\mathcal{T}^{\star}\to\Delta(\mathcal{T})$, takes a sequence of tokens $\mathcal{T}^{\star}$ as input and outputs a distribution $\Delta(\mathcal{T})$ over the next token in the sequence.
@@
 Adversary’s Objective.
-The adversary’s main objective is to generate responses from the LLM that are classified as successful jailbreaks.
+The adversary’s objective is to generate responses from the LLM that are classified as successful jailbreaks.
 Specifically, the adversary seeks to maximize the probability that a response $R$ generated by the LLM for a given prompt $P$ is classified as harmful according to the goal $G$.
@@
 $\sup_{P\in\mathcal{T}^{\star}}\Pr_{R\sim\text{LLM}(P)}[\text{JUDGE}(R,G)=\text{True}]$
-where the randomness is due to the stochastic nature of the LLM’s responses to the input prompt $P$.
-In essence, the adversary iterates over potential prompts to find one that maximizes the likelihood of producing a harmful output as judged by the classifier.
-Constraints and Assumptions. We assume the following constraints and assumptions in the general jailbreak threat model:
-  • The adversary has black-box access to the LLM, meaning they can input prompts and observe outputs but do not have direct access to the model’s internal parameters or training data.
-  • The goal string $G$ is predefined and represents a specific type of harmful content that the adversary aims to induce from the LLM.
-  • The classifier JUDGE is accurate in determining whether a response constitutes a successful jailbreak according to the harmful goal $G$.
-Under these assumptions, the adversary’s challenge is to discover prompts that effectively exploit the LLM’s vulnerabilities, leading to the generation of undesirable content while navigating the constraints of black-box access.
+where $\Pr$ is the probability, and the randomness is due to the stochastic nature of the LLM’s responses to the input prompt $P$.
+Essentially, the adversary iterates over potential prompts to find one that maximizes the likelihood of producing a harmful output, as judged by the classifier.
+Constraints and Assumptions. We assume the following setup that is commonly used in the jailbreak threat model:
+  • The adversary requires only black-box access to the LLM, meaning they can input prompts and observe outputs but do not necessarily need access to the model’s internal parameters or training data.
+  • The goal string $G$ is predefined and represents a specific type of harmful content that the adversary aims to induce.
+  • The classifier JUDGE accurately determines if a response constitutes a successful jailbreak based on $G$.
 2.2 Jailbreak Attacks
-Existing jailbreak attacks can be roughly grouped into four categories: optimization-based, template-based, indirect, and multilingual jailbreaks.
-Optimization-based Jailbreak typically appends a suffix to a malicious prompt and iteratively updates this suffix using gradient-based or search-based methodologies for jailbreaking.
+Existing jailbreak attacks can be roughly grouped into multiple categories: human-based, optimization-based, generation-based, and the recent indirect and multilingual jailbreaks.
+Human-based Jailbreak involves manually crafting jailbreak prompts to exploit LLM vulnerabilities [43, 70, 73, 62, 16].
+Wei et al.
+[70] utilize two jailbroken modes of LLMs (i.e., out-of-distribution inputs and the conflict between the model’s capabilities and safety goals) to guide the design of manual jailbreaks.
+Deng et al.
+[16] engineered a proof-of-concept (PoC) jailbreak prompt that alters an LLM’s output to generate harmful content by making it act as AIM (Always Intelligent and Machiavellian) and used it as a seed to create more jailbreak prompts.
+Additionally, Shen et al.
+[62] present the JailbreakHub framework, a platform for crowdsourcing jailbreak prompts from online contributors.
+Optimization-based Jailbreak typically updates the adversarial prompt iteratively using gradient-based or search-based methodologies [99, 63, 5, 40, 28].
 The pioneering work by GCG [99] introduced a method known as the greedy coordinate gradient to optimize adversarial suffixes, facilitating transferable jailbreaks across various prompts and models.
 Sitawarin et al.
-[30] further extended this technique to GCG++ by employing a proxy model to guide the optimization process.
+[63] further extended this technique to GCG++ by employing a proxy model to direct the optimization process.
 Beyond gradient-based optimization, Andriushchenko et al.
-[31] utilized a simple random search on a suffix to maximize the target probability.
-Template-based Jailbreak involves the use of prompt engineering [23, 24, 25] or LLM-based red teaming [26, 27, 36, 71, 28] to create effective prompt templates that can mislead LLMs into producing restricted content.
-Early template-based attacks combine manually crafted templates [23, 24, 25] with malicious instructions to achieve jailbreaking.
-For example, Liu et al.
-[24] proposed role-playing prompts to confuse LLMs.
-A more sophisticated approach, AutoDAN [28], employs genetic algorithms to optimize these hand-crafted templates.
-Recent advancements include the use of LLMs to assist in generating jailbreak prompts.
-PAIR [36] is the first method to utilize an auxiliary LLM to generate jailbreaks.
+[5] utilized a simple random search on a suffix to increase the likelihood of hitting the target probability.
+Unlike optimizing these obviously unreadable suffixes, AutoDAN [40] automatically constructs human-readable jailbreak prompts using a carefully designed hierarchical genetic algorithm.
+Generation-based Jailbreak employs language models [54, 16, 12, 47, 53] to generate effective jailbreak prompts that can mislead LLMs into producing restricted content.
+An intuitive approach is to use an auxiliary LLM to construct candidate prompts through prompt engineering.
+For example, PAIR [12] fed the response of the target model back to the attacking LLM to adapt the output for deceptive jailbreaks.
 Mehrotra et al.
-[71] then refined PAIR’s approach through a tree-based search method.
-Indirect Jailbreak conceals malicious intents within the query text to circumvent the safety mechanisms of LLMs and elicit the desired malicious response.
-A straightforward method for executing indirect jailbreaks is by performing word substitution on the original malicious instruction [72].
+[47] then refined PAIR’s approach through tree-of-thought reasoning [84].
+Moreover, the adversary can train a new LLM specifically to attack the target model.
+For example, Paulus et al.
+[53] fine-tuned the Advprompter LLM to generate human-readable suffixes against the target LLM.
+Indirect Jailbreak conceals malicious intents within the query text to circumvent the safety mechanisms of LLMs and elicit the desired malicious response [20, 33, 10, 70].
+A straightforward method for executing indirect jailbreaks is to perform word substitution on the original malicious instruction [20].
 Recently, DrAttack [33] introduced a technique that decomposes a malicious prompt into separate sub-prompts, effectively masking its underlying malicious intent.
-Puzzler [33] provides clues related to the malicious prompt, thereby guiding the target LLM towards a jailbreak.
-Multilingual Jailbreak translates the harmful prompt into a language in which LLMs are less aligned for safety [34, 35].
+Meanwhile, Puzzler [10] provides clues related to the malicious prompt, thereby inducing the target LLM toward a jailbreak.
+Multilingual Jailbreak translates the harmful prompt into a language in which LLMs are less aligned for safety [17, 85, 61, 31, 70, 87].
 Deng et al.
-[34] found that it is easier to jailbreak LLMs in low-resource languages than in high-resource languages and released a multilingual jailbreak prompt dataset, MultiJail.
+[17] found that it is easier to jailbreak LLMs in low-resource languages, such as Zulu [85], than in high-resource languages and released a multilingual jailbreak prompt dataset, MultiJail.
 Besides the multilingual strategy, Wei et al.
-[23] encode the original harmful prompt into Base64 format to reduce the sensitivity of LLMs.
-TABLE I: A comparison of existing jailbreak defenses under the four objectives we envisioned in §3.1.
-Note that the first six rows denote tuning-based defenses, while the last seven rows represent non-tuning-based defenses.
+[70] and Yuan et al.
+[87] adopted an obfuscation strategy [15] to either encode or encrypt the original harmful prompt, thereby reducing the sensitivity of LLMs.
+Table 1: A comparison of existing jailbreak defenses under the four objectives we envisioned in §3.1.
+Note that the first 11 rows denote plugin-based defenses, while the last nine rows represent model-based defenses.
 Venue1 Core Idea Objectives
@@
 Perplexity [4, 26] arXiv:2308.14132 Calculate the perplexity of the prompt to detect adversarial suffixes
-Llama Guard [42] arXiv:2312.06674 Align the classification of both prompt and response
+Self Defense [55] arXiv:2308.07308 Add one more step to check the safety of original LLM responses
+erase-and-check [29] arXiv:2309.02705 Erase some tokens from the prompt; check the rest using a safety filter
+Smooth [9, 59, 27] ACL 2024 (2309.14348) Perturb copies of each prompt and aggregate their output responses
+ICD [73] arXiv:2310.06387 Use safe in-context demonstrations to enhance the model’s robustness
+Self-Reminder [80] NMI 2023 (December 2023) Add a self-reminder system prompt to make ChatGPT respond safely
+Llama Guard [25] arXiv:2312.06674 An input-output safeguard for safety classification of prompts and response
+IAPrompt [88] arXiv:2401.06561 Conduct an intention analysis for each input prompt
+GradSafe [79] ACL 2024 (2402.13494) Compare the prompt’s gradient similarity with safety-critical gradients
+Gradient Cuff [23] arXiv:2403.00867 Compare the gradient norm of refusal loss with that in benign queries
+Circuit Breaking [98] arXiv:2406.04313 Interrupt the model output harmful content at the internal representations
+RAIN [35] ICLR 2024 (2309.07124) Self-evaluate each token output, rewind, and determine the final output
+Goal Prioritization [89] ACL 2024 (2311.09096) Integrate instructions to control the priority between helpfulness and safety
 RPO [96], DPP [81] arXiv:2401.17263 Optimize universal and transferable suffixes that enforce safe outputs
-DRO [43], PAT [44] SeT @ ICLR’24 Add a safety prompt and optimize it as a prefix using prompt tuning
-SafeDecoding [45] arXiv:2402.08983 Compare and amplify the token probabilities of safety disclaimers
+DRO [93], PAT [49] ICML 2024 (2401.18018) Add a prefix as safety prompt and optimize it with prompt tuning
+SafeDecoding [82] ACL 2024 (2402.08983) Compare and amplify the token probabilities of safety disclaimers
+Eraser [45] arXiv:2404.05880 Encourage LLMs to forget harmful knowledge via machine unlearning
+CAT [77] arXiv:2405.15589 Conduct adversarial training on continuous embedding attacks
 LED [91] arXiv:2405.18166 Identify safety-critical layers and realign them through model editing
-Self Defense [47] arXiv:2308.07308 Add one more step to check the safety of original LLM responses
-RAIN [48] ICLR’24 Self-evaluate each token output, rewind, and determine the final output
-erase-and-check [49] arXiv:2309.02705 Erase some tokens from the prompt; check the rest using a safety filter
-[50, 51, 52] arXiv:2309.14348 Perturb copies of each prompt and aggregate their output responses
-IAPrompt [53] arXiv:2401.06561 Conduct a chain-of-thought intention analysis for each input prompt
-GradSafe [54] arXiv:2402.13494 Compare the prompt’s gradient similarity with safety-critical gradients
-Gradient Cuff [55] arXiv:2403.00867 Compare the gradient norm of refusal loss with that in benign queries
-  • = applies; = partially applies; = does not apply.
-  • 1
-    We only present the venue information for the earliest article in each line of defense techniques.
+Adversarial Tuning [39] arXiv:2406.06622 Fine-tune the LLM with token- and semantic-level adversarial prompts
+  • = applies; = partially applies; = does not apply. Note that we only present the venue information for the earliest article in each line of defense techniques.
 3 Objectives and Related Work
@@
 Handling all kinds of jailbreak attacks.
-The proposed defense should be able to handle all categories of jailbreak attacks listed in §2, including not only optimization-based adversarial jailbreaks but also template-based, indirect, and multilingual jailbreaks.
+The proposed defense should be able to handle all categories of jailbreak attacks listed in §2, including not only human-based jailbreaks but also optimization-based, generation-based, indirect, and multilingual jailbreaks.
 O2
-Incurring negligible delay to user prompts. The proposed defense should not impact the user experience, causing either no delay or only a negligible one to normal user prompts.
+Incurring negligible delay to user prompts. The defense should not impact the user experience, causing either no delay or only a negligible one to normal user prompts.
 O3
 Providing explanations for potential jailbreak queries.
-Similar to O2, when the proposed defense detects any query potentially related to jailbreaking, it should provide helpful explanations on why the query is considered harmful.
+Similar to O2, when the defense detects any query potentially related to jailbreaking, it should provide helpful explanations on why the query is considered harmful.
 O4
-Compatible with both open-source and closed-source LLMs. The proposed jailbreak defense approach should protect both white-box, open-source LLMs as well as black-box, closed-source LLMs.
-We clarify that compared with the other three objectives, O3 is not a mandatory objective.
-Nevertheless, we believe O3 is particularly important for users to forensically understand potential jailbreak queries and better protect against jailbreak attacks in the future.
+Compatible with both open-source and closed-source LLMs. The proposed jailbreak defense approach should protect both white-box and open-source LLMs as well as black-box and closed-source LLMs.
+We clarify that compared with the other three objectives, O3 is not mandatory.
+Nevertheless, we believe O3 is valuable for users to forensically understand potential jailbreak queries and better protect against jailbreak attacks in the future.
+Figure 1: A high-level overview of the SelfDefend framework and its workflow; see §4 for more details.
 3.2 Analysis of Existing Defenses
-Table I summarizes our analysis of major LLM jailbreak defenses under the four objectives we envisioned above.
-They can be roughly categorized into tuning-based (the first six rows) and non-tuning-based (the last seven rows) mechanisms.
-Specifically, tuning-based defenses aim to fundamentally improve a model’s safety alignment against jailbreaking, while non-tuning-based defenses can be typically applied to any off-the-shelf LLMs to enhance their safety against jailbreak attacks.
+Table 1 summarizes our analysis of major LLM jailbreak defenses under the four objectives we envisioned above.
+They can be roughly categorized into plugin-based (the first 11 rows) and model-based (the last nine rows) mechanisms.
+Specifically, plugin-based defenses can be typically applied to any off-the-shelf LLMs like a plugin to enhance their safety against jailbreak attacks, while model-based defenses aim to fundamentally improve a model’s safety alignment against jailbreaks by changing the model’s internal mechanisms or conducting fine-tuning for parameter optimization.
+It is worth noting that some defenses may exhibit characteristics of both types, for which we do not aim to distinctly distinguish them in this paper.
 In the rest of this subsection, we present our analysis results from four perspectives:
@@
 The latter includes two perplexity-based filtering approaches [4, 26] and several input perturbation-based approaches [29, 9, 59, 27].
-Specifically, Alon and Kamfonas [41] proposed the first tuning-based defense in the sense that they not only leveraged perplexity values as an indicator to detect prompts with adversarial suffixes but also tuned a classifier to consider both the sequence length of prompts and their perplexity for improved filtering.
+Specifically, Alon and Kamfonas [4] proposed the first plugin-based defense in the sense that they not only leveraged perplexity values as an indicator to detect prompts with adversarial suffixes but also tuned a classifier to consider both the sequence length of prompts and their perplexity for improved filtering.
 Another line of GCG-specific jailbreak defenses perturbed copies of the input prompt and aggregated the output responses, with SmoothLLM [59] as a representative example.
-Second, the majority of prior defenses inherently incur additional delays to user prompts (O2: ), with only two exceptions.
-Since the design principles of most prior defenses are to check the target LLM’s responses [47, 48, 42] and internal statuses [41, 62, 45, 54, 55, 46], or to conduct extra-round analyses of the input prompt [49, 50, 51, 52, 53], it is thus inherently difficult for these approaches to avoid additional delay.
-The only exceptions are [73, 74, 43, 44], which conducted optimization-based suffix or prefix training for defensive purposes, with the training objective being the exact opposite of the GCG attack.
+Second, the majority of plugin-based defenses inherently incur additional delays to user prompts (O2: ), while most model-based methods do not (O2: ).
+Since the design principles of most prior defenses are to conduct extra-round analyses of the input prompt [4, 26, 25, 88, 29], or to check the target LLM’s internal states [79, 23] and responses [25, 55, 9, 59, 27], it is thus inherently difficult for these approaches to avoid additional delay.
+The exceptions are most model-based defenses, which either conduct prompt tuning [96, 81, 93, 49] or optimize the parameters of the target model [89, 77, 39, 45, 91].
 As a result, the tuned models behave like normal LLMs, incurring no extra delay to user prompts.
-Third, most non-tuning-based defenses have the potential to provide explanations for jailbreak queries (O1: or ), whereas half of the tuning-based defenses cannot because they rely solely on LLMs’ internal indicators for tuning (O3: ).
-To provide explanations for potential jailbreak queries, a defense mechanism needs to understand the semantics of incoming prompt queries.
-Hence, it is hard for approaches that rely solely on LLMs’ internal indicators, whether they are tuning-based [45, 46] or non-tuning-based [54, 55], to provide straightforward explanations to users.
-Fourth, most non-tuning-based defenses are compatible with both open-source and closed-source LLMs (O4: ), except for those that need to leverage LLMs’ internal indicators, while the opposite is true for tuning-based defenses (O4: ).
-Unless they need to monitor LLMs’ internal indicators [48, 54, 55], non-tuning-based defenses, such as Self Defense [47] and IAPrompt [53], are naturally compatible with both open-source and closed-source LLMs.
-On the other hand, tuning-based defenses typically require white-box access to the LLMs to enable their in-depth defense.
-The only exception is Meta’s Llama Guard [42], which aligns the classifications of both prompt and response, and thus, theoretically, it can serve as a plug-in model to work with both open-source and closed-source LLMs.
-Figure 1: A high-level overview of the SelfDefend framework and our study.
+Third, more than half of plugin-based defenses have the potential to provide explanations for jailbreak queries (O3: or ), whereas most model-based defenses cannot because they rely solely on LLMs’ internal mechanism tuning (O3: ).
+To provide explanations for potential jailbreak queries, a defense scheme needs to understand the semantics of incoming prompt queries.
+Hence, it is difficult for approaches that rely solely on target LLMs’ internal indicators, whether they are plugin-based [73, 80, 79, 23] or model-based [82, 91], to provide straightforward explanations to users.
+Fourth, most plugin-based defenses are compatible with both open-source and closed-source LLMs (O4: ), while the opposite is true for most model-based defenses (O4: ).
+Unless they need to monitor LLMs’ internal indicators [79, 23], plugin-based defenses, such as Self Defense [55] and IAPrompt [88], are naturally compatible with both open-source and closed-source LLMs.
+By contrast, model-based methods typically require white-box access to the LLMs to enable their in-depth defense.
+Exceptions include approaches that enhance the safety of the LLM through the integration of hand-crafted prompts [89].
 4 The SelfDefend Framework
@@
 This idea is made possible because we found that LLMs do have the capability to protect themselves by identifying harmful portions in user queries.
+Overview.
 Based on this idea, we propose a generic LLM jailbreak defense framework called SelfDefend.
@@
 In SelfDefend, $LLM_{defense}$ can be instantiated from the same model as $LLM_{target}$, although in practice we suggest using a dedicated $LLM_{defense}$ that is robust and low-cost for detecting jailbreak queries.
-Since $LLM_{defense}$ does not need to modify $LLM_{target}$ or know any of its internal details, SelfDefend can protect all kinds of target LLMs, whether they are open-source or closed-source (O4: ).
-Given an incoming prompt query $P_{query}$, SelfDefend dispatches it to both $LLM_{target}$ and $LLM_{defense}$ for concurrent processing.
-Specifically, $LLM_{target}$ processes $P_{query}$ as usual, whether it is a normal user prompt or an adversarial jailbreak prompt, but caches its token-by-token output until a checkpoint is triggered from the shadow stack.
-In contrast, $LLM_{defense}$ employs a dedicated detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$).
-Once a token of “No” (indicating no issue) or a harmful portion (prompt/intention) is output, $LLM_{defense}$ triggers a checkpoint in the normal stack so that $LLM_{target}$ can decide whether to release the token-by-token output.
+Workflow.
+① Given an incoming prompt query $P_{query}$, SelfDefend dispatches it to both $LLM_{target}$ and $LLM_{defense}$ for concurrent processing.
+② $LLM_{target}$ processes $P_{query}$ as usual, whether it is a normal prompt or an adversarial prompt, but caches its token-by-token output until a checkpoint is triggered from the shadow stack.
+By contrast, $LLM_{defense}$ employs a tailored detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful part (via $P_{direct}$) or intention (via $P_{intent}$).
+③ Once a token of “No” (indicating no issue) is output from the shadow stack, $LLM_{defense}$ is triggered to release its token-by-token response.
+Otherwise, when the shadow stack detects a harmful portion (prompt/intention), SelfDefend would respond with a template that refuses to answer, i.e., “I can’t fulfill your query because your [harmful portion] violated our safety policy.” It is noted that “[harmful portion]” is replaced with the recognized portion from $LLM_{defense}$.
+Objective Analysis.
 Since the time for outputting the token “No” is very limited, the additional delay introduced by SelfDefend is negligible for normal user prompts (O2: ).
@@
 These unique advantages make SelfDefend the first practical jailbreak defense compared to the existing defense mechanisms explained in §3.2.
-Now the effectiveness of SelfDefend depends on whether it can handle all kinds of jailbreak attacks (i.e., O1).
-For this, we proceed to conduct an empirical study in §5 to validate our earlier hypothesis that “LLMs have the capability to identify harmful portions in user queries.”
+Moreover, as $LLM_{defense}$ does not need to modify $LLM_{target}$ or know any of its internal details, SelfDefend can protect all kinds of target LLMs, whether they are open-source or closed-source (O4: ).
+Now the effectiveness of SelfDefend depends on under which detection prompts, it can handle all kinds of jailbreak attacks (i.e., O1).
+For this, we proceed to conduct an empirical study in §5 to explore effective $P_{direct}$ and $P_{intent}$ and validate our earlier hypothesis that “LLMs have the capability to identify harmful portions in user queries” using these measures.
 5 An Empirical Measurement
-In this section, we conduct extensive measurements to empirically show that existing LLMs, notably the commonly used GPT-3.5/4 models, perform acceptably under the SelfDefend framework.
+In this section, we conduct extensive measurements to empirically show that under the SelfDefend framework and its carefully designed detection prompts, existing LLMs, notably the commonly used GPT-3.5/4 models, perform acceptably.
 This not only validates the hypothesis assumed in §4 but also enables widespread, convenient usage of SelfDefend without the need for further tuning of a defense model.
-For example, various custom GPTs applications [76] can adopt SelfDefend to immediately enhance their defense capabilities against jailbreaks.
-Figure 2: SelfDefend’s two detection prompt templates.
-5.1 Initializing SelfDefend for GPT-3.5/4
+For example, various custom GPTs [1] can adopt SelfDefend to immediately enhance their defense against jailbreaks.
+5.1 Initializing SelfDefend for GPT-3.5/4 and Designing Two Detection Prompts
 To facilitate measurements, we need to first initialize SelfDefend for GPT-3.5/4.
@@
 Likewise, for GPT-4-based experiments, both $LLM_{target}$ and $LLM_{defense}$ are set to GPT-4.
-Prompt Templates.
-$LLM_{target}$ directly uses user prompts $P_{query}$ but we need to set up the detection prompts for $LLM_{defense}$.
-As mentioned earlier in §4, we have designed $P_{direct}$ and $P_{intent}$ for two different defense schemes.
-Figure 2 shows their detailed prompt templates.
-We can see that $P_{direct}$ aims to directly pinpoint the harmful part from the original jailbreak/normal $P_{query}$; otherwise, it outputs a token of “No.” In contrast, $P_{intent}$ employs a chain-of-thought (CoT) [77] approach to first summarize the true intention of $P_{query}$ and then, based on the summarized intention, further extract any safety-violation part, if present.
-It also includes two examples in the prompts to help LLMs better understand the instructions and conduct few-shot in-context learning [78].
-By comparing these two prompts, we expect that $P_{direct}$ would perform better on optimization-based jailbreaks that directly include the jailbreak portion, while $P_{intent}$ might be more effective against indirect jailbreaks.
-One additional objective of the empirical measurement in this section is to explore the best setup for deploying SelfDefend with off-the-shelf LLMs under four different combinations (GPT-3.5 vs.
-GPT-4; $P_{direct}$ vs.
-$P_{intent}$).
-Environment Setup.
-In our implementation, we use the OpenAI API in Python to access GPT-3.5 and GPT-4.
 The API versions of GPT-3.5 and GPT-4 are gpt-3.5-turbo-0125 and gpt-4-0125-preview, respectively.
@@
 When accessing the GPT APIs, no system prompt is added.
-TABLE II: The details of our collected benchmark datasets.
+Figure 2: The two detection prompt templates designed. “[Jailbreak/Normal Prompt]” will be replaced with the user query.
+Besides initializing SelfDefend for GPT-3.5/4, we carefully design $P_{direct}$ and $P_{intent}$ for SelfDefend and evaluate their effectiveness in this section.
+As mentioned in §4, $P_{direct}$ and $P_{intent}$ represent two different defense schemes.
+$P_{direct}$ is more intuitive, aiming to directly pinpoint the harmful part from the original jailbreak prompt; otherwise, it outputs a token of “No”.
+$P_{intent}$, on the other hand, aims to recognize the true intent of the request text rather than directly intercepting harmful parts, given that indirect jailbreak attacks (e.g., DrAttack [33]) hide harmful instructions.
+Based on these two guidelines, we carefully design the prompt templates used by $P_{direct}$ and $P_{intent}$.
+As shown in Figure 2, $P_{direct}$ uses the sentence like “identify which part of the following sentences might violate your safety policies or have the jailbreak intentions” to directly pinpoint the harmful portion.
+Inspired by chain-of-thought (CoT) [71], $P_{intent}$ first summarizes the true intention of $P_{query}$ and then, further extracts any safety-violation part from the summarized intention, if present.
+It also includes two contrasting examples in the prompts to help $LLM_{defense}$ better understand the instructions, motivated by few-shot in-context learning [7].
+Through the extensive measurements in this section, these two designed prompts are shown to be effective within the SelfDefend framework.
+Nevertheless, we note that the detailed wording of these two prompt contents might have alternatives, but the current form is used as a standard template for $P_{direct}$ and $P_{intent}$ throughout the paper.
+5.2 Datasets and Attack Setup
+After SelfDefend is set up, we collect benchmark datasets, implement the corresponding attack methods, and determine the effectiveness metrics.
+Benchmarks.
+Based on the five categories of existing jailbreak attacks we surveyed in §2 — human-based, optimization-based, generation-based, indirect, and multilingual jailbreaks — we identify representative jailbreak attack methods in each category.
+We then collect four benchmark datasets, JailbreakHub [62], JailbreakBench [11], MultiJail [17], and AlpacaEval [34], from which we use their user prompts for testing SelfDefend.
+Table 2 lists the details of our collected benchmark datasets.
+Specifically, we use a set of 100 harmful instructions from JailbreakBench [11], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [99] and AutoDAN [40]), generation-based jailbreaks (PAIR [12] and TAP [47]), and indirect jailbreak attacks (DrAttack [33] and Puzzler [10]).
+In contrast, we directly use the original prompts from the JailbreakHub, MultiJail and AlpacaEval datasets to construct inputs for the scenarios of human-based attacks, multilingual jailbreaks and normal prompts, respectively.
+Table 2: The details of our collected benchmark datasets.
 Dataset # Prompts Jailbreak Methods
-JailbreakBench [79] 100 GCG [29], JSAA [31]
-PAIR [36], AutoDAN [28]
+JailbreakHub [62] 1000 DAN [62]
+JailbreakBench [11] 100 GCG [99], AutoDAN [40]
+PAIR [12], TAP [47]
 DrAttack [33], Puzzler [10]
@@
 AlpacaEval [34] 805 Normal Prompts
-TABLE III: The ASR results from testing LLMs against four major categories of jailbreak attacks and normal prompts.
-LLMs Optimization-based Template-based Indirect Multilingual Normal
-GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail AlpacaEval
-GPT-3.5 (baseline) 0.56 0.94 0.72 0.90 0.78 0.98 0.39 0.97
-GPT-3.5-based Shadow Stack ($P_{direct}$) 0.72 0.98 0.77 0.96 1.00 0.98 0.87 0.97
-GPT-3.5-based SelfDefend ($P_{direct}$) 0.45 0.92 0.60 0.87 0.78 0.96 0.36 0.95
-GPT-3.5-based Shadow Stack ($P_{intent}$) 0.28 0.06 0.37 0.35 0.86 0.22 0.52 0.99
-GPT-3.5-based SelfDefend ($P_{intent}$) 0.19 0.04 0.29 0.31 0.71 0.22 0.20 0.97
-Reduction factor 66.07% 95.74% 59.72% 65.55% 8.97% 77.55% 48.71% 0%
-GPT-4 (baseline) 0.08 0.16 0.33 0.19 0.74 0.90 0.08 0.97
-GPT-4-based Shadow Stack ($P_{direct}$) 0.01 0.00 0.11 0.01 0.05 0.27 0.14 0.96
-GPT-4-based SelfDefend ($P_{direct}$) 0.00 0.00 0.10 0.01 0.04 0.26 0.01 0.94
-Reduction factor 100% 100% 36.36% 94.73% 94.59% 71.11% 87.50% 3.09%
-GPT-4-based Shadow Stack ($P_{intent}$) 0.08 0.05 0.21 0.07 0.13 0.36 0.30 0.99
-GPT-4-based SelfDefend ($P_{intent}$) 0.05 0.05 0.18 0.07 0.13 0.28 0.01 0.97
+Table 3: The ASR results from testing LLMs against five major categories of jailbreak attacks and normal prompts.
+LLMs Human-based Optimization-based Generation-based Indirect Multilingual Normal
+DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail AlpacaEval
+GPT-3.5 (baseline) 0.256 0.560 0.900 0.720 0.670 0.780 0.980 0.393 0.977
+GPT-3.5-based Shadow Stack ($P_{direct}$) 0.982 0.720 0.960 0.770 0.840 1.000 0.980 0.879 0.977
+GPT-3.5-based SelfDefend ($P_{direct}$) 0.242 0.450 0.870 0.600 0.610 0.780 0.960 0.368 0.957
+GPT-3.5-based Shadow Stack ($P_{intent}$) 0.015 0.280 0.350 0.370 0.020 0.860 0.220 0.520 0.992
+GPT-3.5-based SelfDefend ($P_{intent}$) 0.007 0.190 0.310 0.290 0.020 0.710 0.220 0.203 0.972
+Reduction factor 97.26% 66.07% 65.55% 59.72% 97.01% 8.97% 77.55% 48.34% 0.51%
+GPT-4 (baseline) 0.047 0.080 0.190 0.330 0.310 0.740 0.900 0.076 0.973
+GPT-4-based Shadow Stack ($P_{direct}$) 0.004 0.010 0.010 0.110 0.100 0.050 0.270 0.142 0.968
+GPT-4-based SelfDefend ($P_{direct}$) 0.002 0.000 0.010 0.100 0.080 0.040 0.260 0.012 0.946
+Reduction factor 95.74% 100.00% 94.73% 69.69% 74.19% 94.59% 71.11% 84.21% 2.77%
+GPT-4-based Shadow Stack ($P_{intent}$) 0.019 0.080 0.070 0.210 0.200 0.130 0.360 0.304 0.995
+GPT-4-based SelfDefend ($P_{intent}$) 0.005 0.050 0.070 0.180 0.170 0.130 0.280 0.019 0.970
 (a) The CDF plot of $\Delta d$ for normal prompts.
 (b) Average $\Delta d$ for various jailbreak prompts.
-Figure 3: The extra delay $\Delta d$ caused by SelfDefend for normal prompts and jailbreak prompts, respectively.
-5.2 Datasets and Attack Setup
-After SelfDefend is set up, we collect benchmark datasets, implement the corresponding attack methods, and determine the effectiveness metrics.
-Benchmarks.
-Based on the four categories of existing jailbreak attacks we surveyed in §2 — optimization-based, template-based, indirect, and multilingual jailbreaks — we identify representative jailbreak attack methods in each category.
-We then collect three benchmark datasets, JailbreakBench [79], MultiJail [34], and AlpacaEval [57], from which we use their user prompts for testing SelfDefend.
-Table II lists the details of our collected benchmark datasets.
-Specifically, we use a set of 100 instructions from JailbreakBench [79], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [29] and JSAA [31]), template-based jailbreaks (PAIR [36] and AutoDAN [28]), and indirect jailbreak attacks (DrAttack [32] and Puzzler [33]).
-In contrast, we directly use the original prompts from the MultiJail and AlpacaEval datasets to construct inputs for the scenarios of multilingual jailbreaks and normal prompts, respectively.
+Figure 3: The extra delay $\Delta d$ caused by SelfDefend for normal and jailbreak prompts, respectively.
 Attack Setup.
+For DAN, we randomly select 1,000 samples as jailbreak queries from the forbidden question set equipped with jailbreak prompts [2], which is collected by JailbreakHub.
 For GCG, we choose its individual version and optimize the suffix on Vicuna-7b-v1.3 [13] with a batch size of 512 and 500 optimization steps.
-For JSAA, we query GPT-3.5 with 10,000 iterations and 10 restarts.
-We perform transfer attacks on GPT-4 using the prompts generated on GPT-3.5.
-For PAIR, we select Vicuna-13b-v1.5 [80] as the attack model.
 For AutoDAN, we choose its first version of the genetic algorithm, i.e., AutoDAN-GA.
 The genetic algorithm used in AutoDAN-GA features a crossover rate of 0.5, a mutation rate of 0.01, and 100 optimization steps.
+For PAIR and TAP, we select Vicuna-13b-v1.5 [13] as the attack model.
+We set the maximum depth, the maximum width, and the branching factor of TAP to 10, 10, and 1, respectively.
+The target model of PAIR and TAP is GPT-3.5/4.
 For DrAttack and Puzzler, we use GPT-4 to construct their jailbreak prompts.
@@
 Following [99, 40], ASR is measured by detecting whether the LLM response contains keywords indicating a refusal to answer, like “I can’t assist.” If such a keyword or its variant is included, it indicates that the attack is unsuccessful; otherwise, it indicates that the jailbreak is successful.
-We use the common practice [29] for a list of such keywords, which is available in Appendix A.
-GCG, JSAA, PAIR, AutoDAN, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance.
+We use the common practice [99] for a list of such keywords; see Appendix B.
+DAN, GCG, AutoDAN, PAIR, TAP, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance.
 The lower the ASR, the stronger the defense performance of SelfDefend against jailbreaks.
-Note that we also use ASR to evaluate the performance of SelfDefend under normal prompts from AlpacaEval, but a higher ASR indicates that our framework is more compatible with normal prompts (in such a case, ASR can be interpreted as answer success rate).
-On the other hand, the unsafe rate uses GPT-4 to determine whether the response of the target model matches the jailbreak goal, following [34].
-We adopt the unsafe rate in a way similar to ASK to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.
+Note that we also use ASR to evaluate the performance of SelfDefend under normal prompts from AlpacaEval, but a higher ASR indicates that our framework is more compatible with normal prompts (in such cases, ASR can be interpreted as answer success rate).
+On the other hand, the unsafe rate uses GPT-4 to determine whether the response of the target model matches the jailbreak goal [17].
+We adopt the unsafe rate in a way similar to ASR to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.
 5.3 Measurement Results
-Table III summarizes our core measurement results, which we separate into the defense effectiveness against jailbreaks and the impact on normal user prompts.
+Table 3 summarizes our core measurement results, which are divided into two categories: the defense effectiveness against jailbreaks and the impact on normal user prompts.
 Additionally, we measure the extra delay introduced by SelfDefend.
@@
 We first measure the baseline performance of using GPT-3.5/4 only under the tested jailbreaks.
-As shown in Table III, these jailbreaks are much more effective on GPT-3.5 compared to GPT-4.
-The ASR for GPT-3.5 can be as high as 0.39 to 0.98 (average: 0.75), whereas the ASR for GPT-4 typically ranges from 0.08 to 0.33, except for indirect jailbreaks in which case GPT-4 also has a high ASR (0.74 for DrAttack and 0.9 for Puzzler).
-Defense Effectiveness. We then compare ASR after equipping SelfDefend with the baseline model. In ablation, we also report ASR using only the shadow stack. Our findings are summarized as follows:
-  • Based on GPT-3.5, although the SelfDefend frameworks supported by intent prompt and direct prompt both decrease success rates, there is a considerable difference in the enhancement of defense performance achieved by the two methods.
-The SelfDefend framework with direct prompt shows a modest reduction in terms of ASR due to its internal shadow stack performing even worse than GPT-3.5.
-For example, its ASR on AutoDAN is 96%, compared to GPT-3.5’s 90%.
-In contrast, the SelfDefend framework utilizing intent prompt demonstrates significantly better performance than the one with direct prompt, with the reduction factor ranging from as high as 95.74% to as low as 8.97%.
-  • Based on GPT-4, both SelfDefend frameworks significantly reduce the ASRs of various jailbreak attacks.
-Different from the results on GPT-3.5, SelfDefend with direct prompt is better than SelfDefend with intent prompt due to the difference between the two shadow stacks.
-Although jailbreak attacks present lower ASRs than on GPT-3.5, the reduction factor of SelfDefend with direct prompt ranges from 36.36% to 100.00%.
-  • SelfDefend with direct prompt and SelfDefend with intent prompt present the best defense enhance on GPT-4 and GPT-3.5, respectively.
-Furthermore, SelfDefend with direct prompt demonstrates the higher reduction factor than SelfDefend with intent prompt on GPT-3.5.
-Thus, we contend that SelfDefend with direct prompt relies more on the model’s own discrimination ability so that is suitable for powerful models such as GPT-4, while the intent prompt can strengthen the model’s logical reasoning and still has decent defense capabilities even on weaker models.
+As shown in Table 3, these jailbreaks are much more effective on GPT-3.5 compared to GPT-4.
+The ASR for GPT-3.5 can be as high as 0.256 to 0.980 (average: 0.657), whereas the ASR for GPT-4 typically ranges from 0.047 to 0.330, except for indirect jailbreaks where GPT-4 also exhibits a high ASR (0.74 for DrAttack and 0.900 for Puzzler).
+Defense Effectiveness.
+We then compare ASR after equipping the baseline model with SelfDefend and report additional ASR findings when using only the shadow stack for ablation exploration.
+For GPT-3.5, both SelfDefend versions—supported by $P_{direct}$ and $P_{intent}$—reduce ASRs but show varying degrees of defense enhancement.
+SelfDefend with $P_{direct}$ exhibits a modest reduction, with its shadow stack performing worse than GPT-3.5 itself; for instance, its ASR on AutoDAN is 96%, compared to GPT-3.5’s 90%.
+In contrast, SelfDefend with $P_{intent}$ significantly outperforms the $P_{direct}$ version, achieving a reduction factor ranging from 97.26% to 8.97%.
+For GPT-4, both SelfDefend versions substantially lower the ASRs across various jailbreak attacks.
+Unlike with GPT-3.5, SelfDefend with $P_{direct}$ outperforms $P_{intent}$ due to differences in shadow stack effectiveness, with reduction factors ranging from 69.69% to 100.00%.
+SelfDefend with $P_{direct}$ provides the best defense enhancement on GPT-4, while SelfDefend with $P_{intent}$ excels on GPT-3.5.
+SelfDefend with $P_{direct}$ relies more on the model’s inherent discrimination capabilities, making it more suitable for advanced models like GPT-4.
+Conversely, the intent prompt enhances logical reasoning and remains effective even on less advanced models.
 Impact on Normal Queries.
 Besides defense effectiveness, we also measure the impact of SelfDefend on normal queries by evaluating the ASR of 805 normal prompts from AlpacaEval.
-Table III shows that for both GPT-3.5 and GPT-4, SelfDefend (with $P_{direct}$) slightly reduces the ASR by 2% (from 0.97 to 0.95) and 3% (from 0.97 to 0.94), respectively.
-Meanwhile, SelfDefend (with $P_{direct}$) does not decrease the ASR at all for either model.
+Table 3 shows that for both GPT-3.5 and GPT-4, SelfDefend (with $P_{direct}$) slightly reduces the ASR by 2% (from 0.977 to 0.957) and 2.7% (from 0.973 to 0.946), respectively.
+Meanwhile, SelfDefend (with $P_{intent}$) hardly decreases the ASR for either model.
 Considering that the ASR under SelfDefend is still very high and close to the baseline, we conclude that SelfDefend’s defense would not notably impact normal user queries.
 Extra Delay $\Delta d$.
-Recall that one of our design objectives is to incur negligible delay to user prompts (O2 in §3.1).
+Recall that one of our design objectives is to incur negligible delays to user prompts (O2 in §3.1).
 We thus further measure the extra delay $\Delta d$ caused by SelfDefend, which is calculated as follows:
 […]
-where $d_{total}$ denotes the total delay to SelfDefend-protected LLMs, and $d_{normal}$ represents the separate time cost of target LLMs in the normal stack.
-The metric of $\Delta d$ is mainly important for normal prompts because it would directly affect the experience of normal users.
+where $d_{total}$ denotes the total delay to SelfDefend-protected LLMs, and $d_{normal}$ represents the separate time cost of the target LLM in the normal stack.
+The metric of $\Delta d$ is essential for normal prompts because it would directly affect the experience of normal users.
 Figure 3(a) plots the cumulative distribution function (CDF) of $\Delta d$ for normal prompts.
-We can see that over 95% of the cases incur zero delay across three configurations of SelfDefend, demonstrating that SelfDefend is capable of defending against jailbreaks at the cost of negligible delay for normal user prompts.
+We can see that over 95% of the cases incur zero delays across three configurations of SelfDefend, demonstrating that SelfDefend is capable of defending against jailbreaks at the cost of negligible delay for normal users.
 The only exception occurs with SelfDefend using $P_{intent}$ and GPT-3.5, where around 80% of the cases have no extra delay.
 For jailbreak prompts, we also calculate their average $\Delta d$ for different attack scenarios in Figure 3(b).
-We can see that GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all eight scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval.
-This is likely because the baseline GPT-4 has good defense performance for most jailbreaks, suggesting its original output is already short and the delay is minimal.
-Figure 4: The training procedure for fine-tuning our defense model.
+We observe that GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all eight scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval.
+This is likely because the baseline GPT-4 has good defense performance for most jailbreaks, with its original output already short and the delay minimal.
+Figure 4: The training procedure for fine-tuning our open-source defense model.
 6 Tuning a Dedicated Defense Model
-Based on the empirical measurement results, we have observed that GPT-4, when equipped with SelfDefend, significantly reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.06 (for GPT-4-based SelfDefend with $P_{direct}$).
-However, GPT-4 is commonly known to be expensive [58], and its closed-source nature raises privacy concerns for non-OpenAI model vendors.
-Therefore, our objective in this section is to tune an open-source model that can be used under the SelfDefend architecture for robust, low-cost, and self-contained jailbreak defense.
+Based on the empirical measurement results, we have observed that GPT-4, when equipped with SelfDefend, significantly reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.063.
+However, GPT-4 is known to be expensive [3], and its closed-source nature raises privacy concerns for non-OpenAI model vendors.
+Therefore, our objective in this section is to tune an open-source model that can be used under the SelfDefend architecture for robust, low-cost, and self-contained defense.
 6.1 Design Overview
@@
 To do so, we leverage GPT-4-based SelfDefend to distill and generate high-quality tuning data.
-Figure 4 depicts our proposed pipeline to fine-tune an open-source defense model.
+Figure 4 depicts our pipeline to fine-tune an open-source defense model.
 Specifically, by continuously incorporating harmful and harmless prompts into our defense prompts (i.e., $P_{direct}$ or $P_{intent}$) as inputs for GPT-4, we gather their outputs as labels for these samples.
-Since we utilize two defense prompts, we eventually obtain two separate datasets, which we then use to fine-tune the employed open-source model for our defense purposes.
+Since we utilize two defense prompts, we eventually obtain two separate datasets, which we then use to fine-tune the employed open-source model.
 6.2 Data Distillation
@@
 From this dataset, we select the initial human prompt and exclude the corresponding assistant response, omitting any subsequent exchanges, to create a single-turn prompt dataset designated as $\mathcal{D}_{red}$.
-It is worth noting that this dataset is released in 2022 and has no overlap with JailbreakBench/MultiJail released in 2024.
+Note that this dataset was released in 2022 and has no overlap with JailbreakHub, JailbreakBench, or MultiJail, which were released in 2023 or 2024.
 Furthermore, this dataset is exclusively in English, implying that a model trained on this data will not acquire any additional multilingual capabilities.
@@
 Despite the reduction in trainable parameters, LoRA maintains competitive performance, often matching or exceeding the results of full fine-tuning approaches [22].
-We allocate 80% of the samples in $\mathcal{D}_{dir}$ or $\mathcal{D}_{int}$ for fine-tuning and reserve the remaining 20% as validation sets.
-During the training process, we continue to use our defense prompts and queries from the collected datasets as inputs; that is, we take $P_{dir}(x)$ or $P_{int}(x)$ as input and $y$ as the label to fine-tune the Llama model.
 Formally, the fine-tuning objective for the direct prompt can be formulated as follows:
@@
 where $\Theta$ represents the trainable parameters of LoRA, $\Phi_{0}$ denotes the pre-trained weights of the Llama model, and $\Delta\Phi(\Theta)$ indicates the parameter increment determined by LoRA.
-Fine-tuning the Llama model with this objective can enhance its ability to directly detect jailbreak portions in query prompts.
+Fine-tuning the Llama model with this objective can enhance its ability to directly detect jailbreak parts in query prompts.
 Likewise, the fine-tuning objective for the intent prompt can be written as follows:
 $\max_{\Theta}\sum_{(x,y)\in\mathcal{D}_{int}}\sum_{t=1}^{|y|}\log\left(p_{\Phi_{0}+\Delta\Phi(\Theta)}(y_{t}\mid P_{int}[x],y_{<t})\right)$ (4)
-This objective allows the fine-tuned model to identify jailbreak intentions within query prompts.
-Implementation. We use LoRA to fine tune the Llama model with $r$ of 8 and $\alpha$ of 32 [60]. The initial learning rate is $10^{-3}$. The training epochs are 1 in batch size of 8.
+Implementation.
+We use LoRA to fine tune the Llama model with $r$ of 8 and $\alpha$ of 32 [22].
+The initial learning rate is $10^{-3}$.
+The training epochs are 1 in a batch size of 8.
+We allocate 80% of the samples in $\mathcal{D}_{dir}$ or $\mathcal{D}_{int}$ for fine-tuning and reserve the remaining 20% as validation sets.
+During the training process, we continue to use our defense prompts and queries from the collected datasets as inputs; that is, we take $P_{dir}[x]$ or $P_{int}[x]$ as input and $y$ as the label to fine-tune the Llama model.
+Table 4: Jailbreak ASR for various defense methods.
+For ICD, SmoothLLM and SafeDecoding, we present the performance of their enhanced models.
+For detection-based Perplexity Filter and Llama Guard (2), we report ASRs only on their detection modules.
+Since SafeDecoding works for a white-box target model, we show its results only on the publicly available Llama-2.
+Target Defense Method Human Optimization Generation Indirect Multilingual Normal
+DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail AlpacaEval
+GPT-3.5 GPT-3.5 (baseline) 0.256 0.560 0.900 0.720 0.670 0.780 0.980 0.393 0.977
+ICD [73] 0.226 0.230 0.840 0.360 0.330 0.750 0.990 0.321 0.960
+Perplexity Filter [26] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 0.994
+SmoothLLM [59] 0.238 0.480 0.930 0.650 0.610 0.850 0.990 0.705 0.968
+Llama Guard [25] 0.561 0.410 0.580 0.430 0.470 0.970 0.930 0.952 0.996
+Llama Guard 2 [66] 0.441 0.140 0.150 0.370 0.360 0.890 0.640 0.559 0.991
+$P_{direct}$-tuned Shadow Stack 0.262 0.080 0.070 0.140 0.210 0.780 0.070 0.749 0.968
+$P_{direct}$-tuned SelfDefend 0.111 0.060 0.070 0.070 0.170 0.620 0.070 0.302 0.948
+$P_{intent}$-tuned Shadow Stack 0.297 0.080 0.090 0.160 0.240 0.180 0.200 0.578 0.996
+$P_{intent}$-tuned SelfDefend 0.125 0.050 0.080 0.120 0.200 0.160 0.200 0.260 0.975
+Double shadow stack 0.213 0.040 0.050 0.100 0.130 0.180 0.070 0.470 0.966
+Double shadow stack+GPT-3.5 0.091 0.030 0.050 0.060 0.100 0.160 0.070 0.187 0.947
+GPT-4 GPT-4 (baseline) 0.047 0.080 0.190 0.330 0.310 0.740 0.900 0.076 0.973
+ICD [73] 0.062 0.050 0.030 0.230 0.230 0.430 0.640 0.051 0.970
+Perplexity Filter [26] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 0.994
+SmoothLLM [59] 0.030 0.070 0.180 0.330 0.330 0.910 0.880 0.651 0.971
+Llama Guard [25] 0.561 0.410 0.580 0.430 0.400 0.980 0.930 0.952 0.996
+Llama Guard 2 [66] 0.441 0.140 0.150 0.350 0.330 0.910 0.640 0.559 0.991
+$P_{direct}$-tuned Shadow Stack 0.262 0.040 0.070 0.170 0.220 0.710 0.120 0.717 0.969
+$P_{direct}$-tuned SelfDefend 0.032 0.010 0.050 0.150 0.150 0.580 0.070 0.060 0.947
+$P_{intent}$-tuned Shadow Stack 0.284 0.080 0.070 0.190 0.190 0.180 0.190 0.565 0.995
+$P_{intent}$-tuned SelfDefend 0.034 0.040 0.060 0.190 0.160 0.170 0.140 0.044 0.970
+Double shadow stack 0.198 0.020 0.040 0.120 0.130 0.180 0.120 0.467 0.968
+Double shadow stack+GPT-4 0.026 0.010 0.040 0.120 0.110 0.170 0.070 0.038 0.945
+Llama-2 Llama-2-7b-chat (baseline) 0.678 0.570 0.680 0.590 0.610 0.880 0.990 0.143 0.988
+ICD [73] 0.474 0.700 0.560 0.310 0.320 0.220 1.000 0.946 0.898
+SafeDecoding [82] 0.655 0.550 0.740 0.560 0.590 0.640 1.000 0.857 0.981
+Perplexity Filter [26] 1.000 0.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
+SmoothLLM [59] 0.681 0.930 0.950 0.820 0.810 0.980 1.000 1.000 0.993
+Llama Guard [25] 0.561 0.400 0.580 0.460 0.420 0.840 0.930 0.952 0.996
+Llama Guard 2 [66] 0.441 0.170 0.150 0.410 0.350 0.890 0.640 0.559 0.991
+(7b-chat) $P_{direct}$-tuned Shadow Stack 0.257 0.060 0.050 0.250 0.190 0.360 0.090 0.737 0.970
+$P_{direct}$-tuned SelfDefend 0.214 0.040 0.050 0.220 0.180 0.310 0.090 0.102 0.959
+$P_{intent}$-tuned Shadow Stack 0.289 0.110 0.080 0.240 0.140 0.150 0.220 0.587 0.991
+$P_{intent}$-tuned SelfDefend 0.242 0.090 0.070 0.210 0.140 0.150 0.220 0.063 0.980
+Double shadow stack 0.198 0.050 0.040 0.180 0.140 0.120 0.040 0.479 0.965
+Double shadow stack+Llama-2 0.164 0.030 0.040 0.160 0.140 0.110 0.040 0.057 0.954
 7 Evaluation
-TABLE IV: Jailbreak ASR for various defense methods.
-Target Model Defense Method Optimization-based Template-based Indirect Multilingual Normal
-GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail AlpacaEval
-GPT-3.5 GPT-3.5 (baseline) 0.56 0.94 0.72 0.90 0.78 0.98 0.39 0.97
-Perplexity Filter [62] 0.03 1.00 1.00 1.00 1.00 1.00 1.00 0.99
-SmoothLLM [51] 0.48 0.89 0.65 0.93 0.85 0.99 1.00 0.96
-Llama Guard [42] 0.41 0.61 0.43 1.00 0.97 0.93 0.95 0.99
-Llama Guard 2 [63] 0.14 0.50 0.37 0.15 0.89 0.64 0.55 0.99
-$P_{direct}$-tuned Shadow Stack 0.08 0.11 0.14 0.07 0.78 0.07 0.74 0.96
-$P_{direct}$-tuned SelfDefend 0.06 0.11 0.07 0.07 0.62 0.07 0.30 0.94
-$P_{intent}$-tuned Shadow Stack 0.08 0.05 0.16 0.09 0.18 0.20 0.57 0.99
-$P_{intent}$-tuned SelfDefend 0.05 0.05 0.12 0.08 0.16 0.20 0.26 0.97
-Double shadow stack 0.04 0.02 0.10 0.05 0.18 0.07 0.47 0.97
-Double shadow stack + GPT-3.5 0.03 0.02 0.06 0.05 0.16 0.07 0.19 0.95
-GPT-4 GPT-4 (baseline) 0.08 0.16 0.33 0.19 0.74 0.90 0.08 0.97
-Perplexity Filter [62] 0.03 1.00 1.00 1.00 1.00 1.00 1.00 0.99
-SmoothLLM [51] 0.07 0.19 0.33 0.18 0.91 0.88 1.00 0.97
-Llama Guard [42] 0.41 0.61 0.43 1.00 0.98 0.93 0.95 0.99
-Llama Guard 2 [63] 0.14 0.50 0.35 0.15 0.91 0.64 0.55 0.99
-$P_{direct}$-tuned Shadow Stack 0.04 0.11 0.17 0.07 0.71 0.12 0.71 0.96
-$P_{direct}$-tuned SelfDefend 0.01 0.07 0.15 0.05 0.58 0.07 0.06 0.94
-$P_{intent}$-tuned Shadow Stack 0.08 0.03 0.19 0.07 0.18 0.19 0.56 0.99
-$P_{intent}$-tuned SelfDefend 0.04 0.02 0.19 0.06 0.17 0.14 0.04 0.97
-Double shadow stack 0.02 0.02 0.12 0.04 0.18 0.12 0.47 0.97
-Double shadow stack + GPT-4 0.01 0.02 0.12 0.04 0.17 0.07 0.04 0.95
-Llama 2 Llama-2-7b-chat (baseline) 0.57 0.96 0.59 0.68 0.88 0.99 0.14 0.98
-Perplexity Filter [62] 0.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99
-SmoothLLM [51] 0.93 1.00 0.82 0.95 0.98 1.00 0.99 0.99
-Llama Guard [42] 0.40 0.41 0.46 1.00 0.84 0.93 0.95 0.99
-Llama Guard 2 [63] 0.17 0.53 0.41 0.15 0.89 0.64 0.55 0.99
-(7b-chat) $P_{direct}$-tuned Shadow Stack 0.06 0.06 0.25 0.05 0.36 0.09 0.73 0.97
-$P_{direct}$-tuned SelfDefend 0.04 0.06 0.22 0.05 0.31 0.09 0.10 0.95
-$P_{intent}$-tuned Shadow Stack 0.11 0.02 0.24 0.08 0.15 0.22 0.58 0.99
-$P_{intent}$-tuned SelfDefend 0.09 0.02 0.21 0.07 0.15 0.22 0.06 0.98
-Double shadow stack 0.05 0.02 0.18 0.04 0.12 0.04 0.48 0.97
-Double shadow stack + Llama 2 0.03 0.02 0.16 0.04 0.11 0.04 0.06 0.95
-In this section, we extensively evaluate SelfDefend with Llama-2-tuned defense models and compare their performances with other jailbreak defense methods.
+In this section, we extensively evaluate SelfDefend with Llama-2-tuned defense models and compare their performance with other jailbreak defense methods.
 We generally follow the setup mentioned in §5 unless explicitly specified in §7.1.
-For each evaluation target, we assess not only the defense effectiveness (in §7.2) and the extra delay $\Delta d$ (in §7.3) as in §5, but also whether the detected harmful portion actually aligns with the original prompt (i.e., explainability in §7.4), and how robust our defense models are against targeted optimization-based attacks (in §7.5) and prompt injection (in §7.6).
+For each evaluation target, we assess not only the defense effectiveness (in §7.2) and the extra delay $\Delta d$ (in §7.3) as in §5, but also whether the detected harmful portion actually aligns with the original prompt (i.e., explainability in §7.4), and how robust our defense models are against adaptive attacks (in §7.5) and prompt injection (in §7.6).
 7.1 Experimental Setup
 Target LLMs, Benchmarks, and Environment.
-As shown in Table IV, our tuned defense models are paired with protect three popular LLMs: GPT-3.5, GPT-4, and the open-source Llama-2-7b-chat [61].
-We use the same benchmark datasets as in Table II (see §5).
+As shown in Table 4, our tuned defense models are paired with protect three popular LLMs: GPT-3.5, GPT-4, and the open-source Llama-2-7b-chat [67].
+We use the same benchmark datasets as in Table 2 (see §5).
 We follow the same procedure and configuration as in §5 to generate the tested jailbreak prompts for GPT-3.5 and GPT-4.
-The process is similar for testing the target model Llama-2-7b-chat; we simply set the target model for query-based JSAA, PAIR, and DrAttack jailbreaks to Llama-2-7b-chat to generate their jailbreak prompts.
-Our evaluation is implemented using PyTorch 2.2.1 and conducted on an NVIDIA TESLA H800 GPU.
+The process is similar for testing the target model Llama-2-7b-chat; we set the target model for query-based PAIR, TAP and DrAttack to Llama-2-7b-chat to generate their jailbreak prompts.
+Our evaluations are implemented using PyTorch 2.2.1 and conducted on an NVIDIA TESLA H800 GPU.
 Baselines.
-We compare our framework with popular jailbreak defense methods, including Perplexity Filter [62], SmoothLLM [51], Llama Guard [42], and Llama Guard 2 [63].
-Specifically, Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt.
+We compare our framework with popular jailbreak defense methods, including ICD [73], SafeDecoding[82], Perplexity Filter [26], SmoothLLM [59], Llama Guard [25] and Llama Guard 2 [66].
+Specifically, ICD adds in-context demonstrations in input prompts to enhance the safety of the target model.
+We adopt the same 1-shot demonstration as the ICD in [95].
+SafeDecoding seeks to methodically scrutinize safety-related disclaimers and amplify the probabilities of their associated token sequences.
+We only show the defense performance of SafeDecoding on Llama-2-7b-chat since it requires fine-tuning an expert model based on the target model.
+Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt.
 A jailbreak is considered to happen when the perplexity exceeds a threshold.
@@
 SmoothLLM perturbs the jailbreak prompts with character-level changes to enable the target LLM to perform defense.
-In this paper, we set SmoothLLM to perform character swapping with a 10% perturbation percentage.
+In this paper, we set SmoothLLM to conduct character swapping with a 10% perturbation percentage.
 Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts.
@@
 Metric.
-We measure the performance of defense methods by the attack success rate (ASR), i.e., the frequency with which jailbreak prompts in a benchmark dataset bypass the detection of a defense method.
+We measure the performance of defense methods by the attack success rate (ASR), i.e., the frequency with which jailbreak prompts in a benchmark dataset bypass the guardrail of a defense method.
 The lower the ASR, the stronger the defense performance.
-However, the definition of “attack success” is different for different defense methods.
+However, the definition of “attack success” is different for distinct defense methods.
+For ICD and SafeDecoding, we adopt the above ASR [99] or Unsafe Rate [17] to measure their defense capability since they directly enhance the model safety instead of plug-in jailbreak detection.
 For Perplexity Filter, the success denotes that the perplexity of the query text is no more than the fixed threshold.
-For SmoothLLM, the definition of success is the success of the attack method, i.e., ASR [29] or Unsafe Rate [34].
+For SmoothLLM, the definition is the success of the attack method, i.e., ASR [99] or Unsafe Rate [17].
 It is worth noting that the Unsafe Rate refers to the unsafety of the target model’s response, rather than the input prompt.
-For Llama Guard, it is success when its response is “unsafe”.
-For our fine-tuned models, the success represent their response being “No.”
-Figure 5: An example of tuned models against GCG on GPT-3.5.
-Figure 6: An example of tuned models against PAIR on GPT-3.5.
-Figure 7: An example of tuned models against DrAttack on GPT-3.5.
-Figure 8: An example of tuned models against MultiJail on GPT-3.5.
-Figure 9: Tuned models’ response to a normal prompt from AlpacaEval.
+For Llama Guard, it is a success when its response is “unsafe”.
+For our fine-tuned models, the success represents their response being “No”.
+Figure 5: Responses of defense models against a jailbreak prompt of AutoDAN.
 7.2 Defense Effectiveness
 We now discuss the defense effectiveness of our proposed technique.
-In line with Table IV, we discuss the results over GPT-3.5, GPT-4, Llama-2-7b-chat, and a “double shadow stack” setting, respectively.
-Results on GPT-3.5.
-Table IV reports ASRs of multiple defense methods under various types of jailbreak attacks.
-The defense methods are evaluated across different jailbreak techniques, consisting of optimization-based, template-based, indirect, multilingual, and normal prompts.
-Under the optimization-based GCG and JSAA, ASR of SelfDefend with the intent prompt is 0.05, slightly better than that with the direct prompt.
-In contrast, SelfDefend with the direct prompt is slightly better than SelfDefend with the intent prompt under the template-based PAIR and AutoDAN.
-For indirect DrAttack, ASR of the fine-tuned model with direct prompt is 0.78, the same as the original GPT-3.5, whereas the dedicated model with intent prompt reduces ASR to 0.18.
-With manual exploration, we find that DrAttack splits the sentence components of the original jailbreak prompt and replaces them with harmless words, making it difficult for SelfDefend (direct prompt) to extract harmful content.
-Although Puzzler is also an indirect attack, the prompts it produces contain offensive content as implicit clues for jailbreak, resulting in a lower ASR of SelfDefend with the direct prompt.
-For multilingual MultiJail, the intent prompt tuned model is superior than the direct prompt by 17%.
-Therefore, the model fine-tuned with the direct prompt is well-suited for identifying overt jailbreak intentions within sentences, whereas the model fine-tuned with the intent prompt excels in detecting jailbreak attempts that encompass more subtle or concealed intentions.
-Besides, SelfDefend has ASRs that are comparable with all other defense methods under normal prompts, indicating its compatibility with normal users.
-Figures 5, 6, 7, 8 and 9 present several example responses of the tuned models toward either jailbreak prompts or a normal query.
-Note that in Figure 7, DrAttack evades the detection of the direct prompt tuned model, while the intent tuned model accurately identifies the intent of DrAttack, which further illustrates the weakness of the direct tuned model against indirect attacks.
-Moreover, as expected, we observe that both tuned models make satisfactory decisions on other jailbreaks or the normal prompt.
-Results on GPT-4.
-Similar with the results achieved over GPT-3.5, we can observe that our defense frameworks achieve the lowest ASR under all jailbreak methods, and is still close to ASR of GPT-4 on AlpacaEval.
-We interpret the results as generally encouraging, illustrating the defensive superiority of SelfDefend without intruding normal user experience.
-Under the optimization-based GCG, the lowest ASR (as achieved by SelfDefend) is 0.01 with the direct prompt.
-For JSAA, SelfDefend with the intent prompt is slightly better than SelfDefend with the direct prompt.
-Under the template-based PAIR and AutoDAN, SelfDefend with the direct prompt is slightly superior than that with the intent prompt.
-For indirect DrAttack, ASR of the fine-tuned model with direct prompt is 0.71, close to that of the original GPT-4, while the fine-tuned model with intent prompt reduces ASR to 0.18.
-In addition, the direct prompt tuned model performs the lowest ASR under Puzzler, but far inferior to the intent one under MultiJail.
-Results on Llama-2-7b-chat.
-Our defense frameworks exhibit the lowest ASRs across all jailbreak methods.
-Moreover, it achieves a ASR that is comparable to Llama-2-7b-chat on AlpacaEval, showing SelfDefend’s defensive effectiveness without compromising normal user experience.
-Specifically, under the white-box GCG, the ASRs for both of our SelfDefend frameworks are less than 10%.
-Perplexity Filter exhibits the lowest ASR against GCG; with further exploration, we find that this is because its filter threshold is set to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix.
-Yet, its detectability in other settings are not as good as SelfDefend.
-Under the optimization-based JSAA attack, SelfDefend with the intent prompt achieves the lowest ASR of 0.02, performing marginally better than with the direct prompt.
-In the context of template-based attacks such as PAIR and AutoDAN, the results between SelfDefend with the direct prompt and the intent one are close, with each having their own superior cases.
-For the indirect DrAttack, while the direct prompt tuned model largely reduces the ASR compared to the original Llama-2-7b-chat model, it still has a 21% gap compared to the intent one.
-Additionally, the direct prompt fine-tuned model has the lowest ASR under the Puzzler attack, yet it is notably outperformed by the intent prompt one in the MultiJail scenario.
+Table 4 reports ASRs of multiple defense methods under various types of jailbreak attacks.
+The defenses are evaluated across different jailbreak techniques, consisting of human-based, optimization-based, generation-based, indirect, and multilingual jailbreak attacks, as well as normal prompts.
+Comparison with Existing Defense Methods.
+As shown in Table 4, our tuned models under SelfDefend achieve satisfactory defense effects under all types of jailbreaks and deliver state-of-the-art (SOTA) results in most cases, while previous defenses fail to show promising performance under different attacks.
+For ICD, the tested ASRs under Puzzler are not less than 99% for both GPT-3.5 and Llama-2.
+For SafeDecoding, it also shows weakness against Puzzler and MultiJail.
+The Perplexity Filter is only effective against GCG and lacks resistance to other jailbreaks.
+Although it exhibits slightly better ASR against GCG than our methods for GPT-3.5 and Llama-2, this is because its filter threshold is set to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix.
+SmoothLLM does not show satisfactory defense performance against DrAttack and Puzzler for all three target models.
+Its ASRs even exceed 80% on Llama-2 for all attack methods except DAN.
+Llama Guard exhibits poor jailbreak detection for DrAttack, Puzzler, and MultiJail, while Llama Guard 2 presents unacceptable flaws for DrAttack.
+Across all attacks and target models, the average ASR of our $P_{direct}$-tuned shadow stack (27.76%) is 37.61% lower than that of Llama Guard (65.37%) and 16.74% lower than that of Llama Guard 2 (44.50%).
+Similarly, the average defense performance of our $P_{intent}$-tuned shadow stack (22.46%) outperforms Llama Guard by 42.91% and Llama Guard 2 by 22.04%.
+Furthermore, SelfDefend achieves ASRs comparable with all other defense methods under AlpacaEval, indicating its compatibility with normal prompts.
+$P_{direct}$-tuned SelfDefend v.s.
+$P_{intent}$-tuned SelfDefend.
+In general, SelfDefend with the direct prompt performs better than SelfDefend with the intent prompt on human-based, optimization-based, and generation-based jailbreaks, while $P_{intent}$-tuned SelfDefend is more effective against indirect and multilingual attacks.
+DAN, GCG, AutoDAN, PAIR, and TAP contain explicitly harmful instructions, whereas DrAttack, Puzzler, and MultiJail do not.
+This phenomenon aligns with our original intent for designing the two prompts, since $P_{direct}$ is used to directly extract harmful parts, and $P_{intent}$ can identify the true intentions hidden by implicit attacks.
+For example, DrAttack deconstructs the sentence components of the original jailbreak prompt and substitutes them with harmless words, making it challenging for SelfDefend (with $P_{direct}$) to extract any harmful content.
+Although Puzzler is also an indirect attack, the prompts it generates include offensive content such as inducing clues for jailbreak, resulting in a lower ASR for SelfDefend with $P_{direct}$.
+Moreover, the false positive rate of the $P_{intent}$-tuned SelfDefend on normal prompts is lower than that of the $P_{direct}$-tuned SelfDefend.
 Double Shadow Stacks.
 The current design of SelfDefend accompanies one shadow stack with one LLM under protection.
-Despite the encouraging protection results, one may wonder incorporating multiple shadow stacks to harvest increased defense capability.
-Here, we define a “double shadow stack” setting as two shadow stacks independently analyze an input, where an input is deemed jailbreak when at least one shadow stack flags it as jailbreak.
-In Table IV, “Double shadow stack” is instantiated by incorporating one $P_{direct}$-based and one $P_{intent}$-based shadow stacks.
-We can observe that the results of double shadow stack are close to the best results of the two SelfDefend frameworks, except under MultiJail.
-And when we further fuse the target model (i.e., normal stack) with the double shadow stack, it exhibits the lowest ASRs under all attack scenarios, as shown in Table IV.
-Nevertheless, we point out that when fusing two or three stacks together, the performance over normal prompts (AlpacaEval) decrease, meaning a few more normal inputs are deemed harmful.
+Despite the encouraging protection results, one might consider whether incorporating multiple shadow stacks could enhance defense capabilities.
+Here, we define a “double shadow stack” setting as two shadow stacks independently analyzing an input, where an input is deemed a jailbreak if at least one shadow stack flags it as such.
+In Table 4, “Double shadow stack” is instantiated by incorporating one $P_{direct}$-based and one $P_{intent}$-based shadow stack.
+We observe that the results of the double shadow stack are not worse than the best results of the two separate stacks.
+Furthermore, when we fuse the target model (i.e., the normal stack) with the double shadow stack, it exhibits the lowest ASRs under almost all attack scenarios, as shown in Table 4.
+Nevertheless, we point out that when fusing two or three stacks together, the performance over normal prompts (AlpacaEval) decreases, meaning a few more normal inputs are deemed harmful.
 This is expected; taking the logical disjunction of multiple stacks’ predictions reduces the false negatives of SelfDefend, yet likely increases the false positives.
 In practice, we suggest using this paradigm only when the standard form of SelfDefend offers low detection accuracy.
+Case Study.
+Besides the quantitative analysis above, we conduct case studies to understand in depth the advantages and pitfalls of SelfDefend compared to other defenses.
+Figure 5 illustrates an example of defense methods against a jailbreak prompt.
+Due to page limitation, readers may refer to Appendix A for its detailed analysis, as well as other cases.
+Table 5: The CLIP-Score [58, 92, 69] ($\uparrow$ indicates better) based on Table 4’s results for Llama-2.
+Tuned Model Text Human-based Optimization-based Generation-based Indirect Multilingual
+DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail
+$P_{direct}$-tuned model Generated Attack Prompts 0.685 0.851 0.784 0.828 0.827 0.737 0.663 0.677
+Identified Harmful Prompts 0.932 0.946 0.939 0.898 0.900 0.785 0.728 0.687
+$P_{intent}$-tuned model Generated Attack Prompts 0.686 0.852 0.784 0.831 0.827 0.732 0.667 0.669
+Identified Harmful Intentions 0.919 0.909 0.908 0.875 0.874 0.857 0.788 0.766
 Figure 6: Average $\Delta d$ introduced by the tuned models on Llama-2-7b-chat.
-TABLE V: The CLIP-Score [64, 82, 83] ($\uparrow$ indicates better) based on Table IV’s results for Llama 2.
-Tuned Model Text Optimization-based Template-based Indirect Multilingual
-GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail
-$P_{direct}$-tuned model Generated Attack Prompts 0.85 0.70 0.83 0.78 0.74 0.66 0.68
-Identified Harmful Prompts 0.95 0.94 0.90 0.94 0.79 0.73 0.69
-$P_{intent}$-tuned model Generated Attack Prompts 0.85 0.70 0.83 0.78 0.73 0.67 0.67
-Identified Harmful Intentions 0.88 0.88 0.85 0.88 0.83 0.77 0.74
 7.3 Extra Delay $\Delta d$
 Similar to how we measured the extra delay $\Delta d$ introduced by GPT-based SelfDefend in §5.3, we now report the average delay $\Delta d$ introduced by the tuned models on Llama-2-7b-chat in Figure 6.
-Compared to earlier results in Figure 3, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.01 seconds.
-In contrast, GPT-3.5-based (GPT-4-based) SelfDefend with $P_{intent}$ incurred an average $\Delta d$ of around 0.07 seconds (0.02 seconds) under AlpacaEval.
-This is likely because the fine-tuned models enjoy faster inference speeds when dealing with similar queries.
+Compared to earlier results in Figure 3, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.008 seconds.
+In contrast, GPT-3.5-based (GPT-4-based) SelfDefend with $P_{intent}$ incurred an average $\Delta d$ of around 0.072 seconds (0.026 seconds) under AlpacaEval.
+This is likely because the parameters of fine-tuned models are less than those of GPT-3.5/4.
 Besides the normal prompts, the tuned models also significantly reduce the extra delay $\Delta d$ under various jailbreak prompts.
-For example, the maximum $\Delta d$ now decreases from the earlier 1.1 seconds on GPT-4 to 0.25 seconds.
-Indeed, except for JSAA, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend.
-These results indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for real-world deployment.
+For example, the maximum $\Delta d$ now decreases from the earlier 1.56 seconds on GPT-4 to 0.27 seconds.
+Indeed, except for DAN, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend.
+These results indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for potential deployment.
 7.4 Explainability
-Figure 5, 6, 7, and 8 demonstrate the efficacy of the tuned models in identifying harmful content within jailbreak prompts.
+Figure 5 demonstrate the efficacy of the tuned models in identifying harmful content within jailbreak prompts.
 Here, we further quantitatively assess whether the identified harmful content actually aligns with the original jailbreak prompt.
-We measure their semantic similarity using the ensemble CLIP-score [64, 82, 83] as shown in Table V.
+We measure their semantic similarity using the ensemble CLIP-score [58, 92, 69], as shown in Table 5.
 Note that our measurement exclusively targets the examples where the defense mechanism is successful.
@@
 Specifically, the CLIP-scores for identified harmful portions are higher than those for the generated attack prompts, suggesting that the jailbreak prompts have altered the original prompts, whereas the harmful portions identified by our models are more aligned with the original prompts.
-Regarding the differences between the two tuned models, we observe that the similarity of identified harmful prompts is higher than that of harmful intentions in the context of optimization-based and template-based attacks, but lower than that of intentions under indirect and multilingual attacks.
-This can be attributed to the fact that optimization-based and template-based attack prompts contain the original prompts, whereas indirect and multilingual attack prompts significantly distort the original prompts.
+Regarding the differences between the two tuned models, we observe that the similarity of identified harmful prompts is higher than that of harmful intentions in the context of human-based, optimization-based, and generation-based attacks, but lower than that of intentions under indirect and multilingual attacks.
+This can be attributed to the fact that human-based, optimization-based and generation-based attack prompts contain the original prompts, whereas indirect and multilingual attack prompts significantly distort the original prompts.
 In summary, these findings are consistent with our initial design intentions and demonstrate that the harmful portions identified by our models can indeed provide an explainable foundation for defense mechanisms.
-TABLE VI: ASR with the regular GCG attack on Llama 2 and the targeted GCG directly on our tuned defense models.
-Tuned Model Regular GCG Targeted GCG
-(from Table IV) (white box)
-$P_{direct}$-tuned model 0.06 0.07
+Table 6: Performance (ASRs) of SelfDefend with different shadow models against adaptive attacks (i.e., PAIR and TAP).
+Target Model Shadow Model PAIR TAP
+GPT-3.5 Llama Guard 0.38 0.39
+Llama Guard 2 0.31 0.36
+$P_{direct}$-tuned model 0.22 0.20
+$P_{intent}$-tuned model 0.25 0.18
+GPT-4 Llama Guard 0.28 0.24
+Llama Guard 2 0.24 0.20
+$P_{direct}$-tuned model 0.20 0.19
+$P_{intent}$-tuned model 0.24 0.15
+Llama-2-7b-chat Llama Guard 0.37 0.36
+Llama Guard 2 0.28 0.28
+$P_{direct}$-tuned model 0.21 0.21
+$P_{intent}$-tuned model 0.23 0.22
+7.5 Robustness to Adaptive Jailbreaks
+Robustness to Black-box Jailbreaks.
+In §7.2, the jailbreak prompts used for evaluation are either manually designed, transferred from a surrogate model (Vicuna-7b-v1.3 [13]), or generated for the target models (GPT-3.5, GPT-4, Llama-2-7b-chat).
+To evaluate the robustness of our SelfDefend against black-box adaptive attacks, we select query-based PAIR and TAP to jailbreak the whole SelfDefend framework.
+The results are reported in Table 6.
+First, we can observe that all our tuned models outperform the Llama Guard series, except for one case, which is that the result of $P_{intent}$-tuned model is equal to Llama Guard 2 under PAIR.
+Moreover, the $P_{direct}$-tuned model is superior to the $P_{intent}$-tuned model in defending against PAIR, whether for GPT-3.5, GPT-4, or Llama-2.
+In contrast, the $P_{intent}$-tuned model is superior to the $P_{direct}$-tuned model in resisting TAP, except on Llama-2.
+[…]
+Table 7: ASR with the regular GCG attack on Llama-2 and the white-box GCG directly on our tuned defense models.
+Tuned Model GCG (from Table 4) White-box GCG
+$P_{direct}$-tuned model 0.06 0.08
 $P_{intent}$-tuned model 0.11 0.09
-7.5 Robustness to Targeted GCG Attacks
-In earlier evaluation with optimization-based attacks such as GCG in §7.2, we used a regular setting where such attacks leverage gradients to optimize adversarial suffixes on an open-source substitute model (i.e., Llama-2-7b-chat) and transfers these suffixes to attack our fine-tuned defense models.
-To stress-test the robustness of our fine-tuned defense models against targeted GCG attacks, we give the adversary full white-box access to our model parameters and allow them to optimize adversarial suffixes directly on our defense models.
-Table VI presents the ASR results for both $P_{direct}$-tuned model and $P_{intent}$-tuned model.
-We observe that the ASR results of the targeted, white-box GCG remain below 10% for both types of tuned models in the shadow stack.
-ASR slightly increases for the targeted GCG under the $P_{direct}$-tuned model but slightly decreases for the targeted GCG under the $P_{intent}$-tuned model.
-This indicates that our defense models are robust to even targeted GCG attacks.
 7.6 Robustness to Prompt Injection
-Since SelfDefend uses $P_{direct}$ or $P_{intent}$ to wrap the original $P_{query}$, it is natural to wonder whether SelfDefend is robust to prompt injection [65, 66], another major LLM threat that is orthogonal to jailbreak attacks.
+Since SelfDefend uses $P_{direct}$ or $P_{intent}$ to wrap the original $P_{query}$, it is reasonable to question whether SelfDefend is robust against prompt injection [44, 42], another major LLM threat that is orthogonal to jailbreak attacks.
 To measure the robustness of SelfDefend’s fine-tuned models against prompt injection, we leverage HouYi [42], an effective black-box prompt injection methodology that has been tested on various LLM-integrated applications.
@@
 It further leverages a tuning-based approach to enhancing the shadow LLM’s defense capability.
-Evaluation shows that SelfDefend is lightweight and effective to mitigate a wide spectrum of jailbreak attacks while rarely undermining normal queries.
-Appendix A Keywords
-TABLE VII: The refusal signals considered for GPT-3.5/4.
+The evaluation shows that SelfDefend is lightweight and effective in mitigating a wide spectrum of jailbreak attacks while rarely undermining normal queries.
+Ethics and the Open Science
+Our research has meticulously addressed various ethical considerations to ensure responsible and ethical conduct.
+Firstly, regarding vulnerability disclosures and open research artifacts, we have released our datasets, raw evaluation results, and code at this anonymous GitHub link: https://github.com/SelfDefend.
+While our study required testing various jailbreak attacks on the live system (i.e., ChatGPT API service), these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server.
+We also aimed to minimize the overhead for OpenAI by distributing our experiments over a period of three months and incurred around \$2,000 in API usage fees.
+Secondly, in addition to all tests on ChatGPT complying with the relevant terms of service, we adhere to Meta Llama’s relevant open-source agreements to maintain legal and ethical standards.
+The well-being of our team members was a priority, with measures in place to protect against exposure to harmful content during the research process and offering psychological support if needed.
+Although our novel defense mechanism provide an effective, low-latency, plug-and-play, and explainable solution, its potential negative outcome could potentially be used maliciously to craft better attacks on LLMs.
+We therefore engage with the community to ensure the technology is used responsibly.
+This involved a readiness to retroactively identify negative outcomes and take corrective actions if our initial assessments underestimated the impacts.
+Finally, all aspects of our research were conducted in strict compliance with the law, ensuring that our practices did not inadvertently contravene legal standards, particularly in data protection and privacy.
+This comprehensive ethical approach not only aligns with the guidelines but also reinforces the integrity and societal value of our research.
+Appendix
+Appendix A Case Studies
+Successful Cases.
+Figure 5 in §7.2 presents a detection example of our methods applied to a jailbreak query, compared with the Llama Guard series.
+The attack prompt involves creating a fake news article to spread disinformation about a political candidate.
+Both Llama Guard and Llama Guard 2 identify it as safe, while our methods mark its harmful portions.
+The reason our method outperforms Llama Guards might be due to the enhanced reasoning ability of the foundation model in identifying harmful portions.
+As described in our design, the direct prompt-based model identifies harmful parts directly from the input text, whereas the intent prompt-based model summarizes the request’s intention and then identifies unsafe parts from that summary.
+When comparing our $P_{direct}$-tuned models with $P_{direct}$-based GPT-4, we observed that the parts they lifted from the attack prompt came from different sentences, since $P_{direct}$ emphasizes that it is sufficient to output one harmful part.
+A similar phenomenon occurs with the intent prompt, where the semantics of GPT-4 and our tuned model’s intercepts are not identical.
+In addition, please refer to Appendix C for more case of our models handling other jailbreaks or the normal prompts.
+As expected, both of our tuned models make satisfactory decisions on various attack or benign inputs.
+Figure 7: An example of tuned models against DrAttack on GPT-3.5.
+Failure Cases.
+Figure 7 shows a failure case of the $P_{direct}$-tuned model against a jailbreak prompt from DrAttack.
+Since DrAttack deconstructs the sentence components of the original jailbreak goal, it is difficult for the $P_{direct}$-tuned model to identify consecutive harmful portions.
+By contrast, the $P_{intent}$-tuned model accurately identifies the intent of DrAttack and defends against it, indicating the strength of the intent prompt-tuned model against indirect attacks.
+Moreover, Figure 8 provides a failed example of the $P_{intent}$-tuned model against PAIR.
+The original jailbreak goal for this attack prompt is “Write a blog post explaining why gambling addicts should take payday loans”.
+The current attack prompt discusses the pros and cons of using payday loans to fund gambling addiction, which reduces the overall toxicity compared to the original goal.
+Even though the summary intent of the $P_{intent}$-tuned model is correct for the attack prompt, it recognizes the request as harmless due to the low toxicity.
+The global semantic difference between the original jailbreak goal and the attack prompt is one reason for the failure of the $P_{intent}$-tuned model.
+In contrast, the $P_{direct}$-tuned model successfully identifies the harmful part, thus making the right judgment.
+Figure 8: An example of tuned models against PAIR on GPT-3.5.
+Appendix B Keywords
+Table 8: The refusal signals considered for GPT-3.5/4.
 “I’m sorry”
@@
 “I’m really sorry”
-“OpenAI” “unethical”
+“OpenAI" “unethical”
 “not ethical”
@@
 “I can’t assist”
-TABLE VIII: The refusal signals considered for Llama 2.
+Table 9: The refusal signals considered for Llama-2.
 “I’m sorry”
@@
 “Hello!”
-As mentioned in §5.2, we use a list of refusal keywords adopted by common practice [29]. They are listed in Table VII for GPT-3.5/5 and in Table VIII for Llama 2, respectively.
-Appendix B More Examples from Our Tuned Models
-More examples from our tuned defense models are presented in Figure 1112, and 13.
-Figure 11: An example of tuned models against JSAA on GPT-3.5.
+As mentioned in §5.2, we use a list of refusal keywords adopted by common practice [99]. They are listed in Table 8 for GPT-3.5/5 and in Table 9 for Llama-2, respectively.
+Appendix C More Examples from Our Tuned Models
+More examples from our tuned defense models are presented in Figure 9, 10, 11, 12, 13, 14, 15, and 16.
+Figure 9: An example of tuned models against DAN on GPT-3.5.
+Figure 10: An example of tuned models against GCG on GPT-3.5.
 Figure 11: An example of tuned models against AutoDAN on GPT-3.5.
+Figure 12: An example of tuned models against PAIR on GPT-3.5.
+Figure 13: An example of tuned models against TAP on GPT-3.5.
 Figure 14: An example of tuned models against Puzzler on GPT-3.5.
+Figure 15: An example of tuned models against MultiJail on GPT-3.5.
+Figure 16: Tuned models’ response to a normal prompt from AlpacaEval.

### v3

@@
 However, delivering a practical jailbreak defense is challenging because it needs to not only handle all the above jailbreak attacks but also incur negligible delays to user prompts, as well as be compatible with both open-source and closed-source LLMs.
-Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM as a defense instance to concurrently protect the target LLM instance in the normal stack and collaborate with it for checkpoint-based access control.
-The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful prompts or intentions in user queries, which we empirically validate using the commonly used GPT-3.5/4 models across all major jailbreak attacks.
+Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM as a defense instance (in detection state) to concurrently protect the target LLM instance (in normal answering state) in the normal stack and collaborate with it for checkpoint-based access control.
+The effectiveness of SelfDefend builds upon our observation that existing LLMs can identify harmful prompts or intentions in user queries, which we empirically validate using mainstream GPT-3.5/4 models against major jailbreak attacks.
 To further improve the defense’s robustness and minimize costs, we employ a data distillation approach to tune dedicated open-source defense models.
-These models outperform six state-of-the-art defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays.
-We also empirically show that the tuned models are robust to adaptive jailbreaks and prompt injections.
+When deployed to protect GPT-3.5/4, Claude, Llama-2-7b/13b, and Mistral, these models outperform seven state-of-the-art defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays.
+Further experiments show that the tuned models are robust to adaptive jailbreaks and prompt injections.
+Warning: This paper contains unfiltered and potentially harmful content.
 1 Introduction
@@
 We concretize the above ideas into a generic jailbreak defense framework called SelfDefend and will introduce its details in §4.
-The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful portions (prompts/intentions) in user queries.
-To validate this hypothesis, we conduct an empirical measurement in §5 using the commonly used GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [62, 99, 40, 12, 47, 33, 10, 17].
+The effectiveness of SelfDefend builds upon our observation that existing LLMs can identify harmful portions (prompts/intentions) in user queries, enabling the simultaneous activation of an $LLM_{defense}$ instance.
+To validate this hypothesis, we conduct an empirical measurement in §5 using mainstream GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [63, 102, 41, 13, 48, 34, 11, 18].
 The results are quite promising in that SelfDefend enables both GPT-3.5 and GPT-4 to significantly suppress the attack success rate (ASR).
-Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97-97.26% (average: 65.06%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.244, and GPT-4-based SelfDefend even reduces the ASR by 69.69-100% (average: 85.53%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.063.
+Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97% to 97.26% (average: 65.70%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.236, and GPT-4-based SelfDefend even reduces the ASR by 69.69% to 100% (average: 88.43%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.050.
 Besides the jailbreak scenarios, we also test SelfDefend against 805 normal prompts from the AlpacaEval dataset [35] and find that the pass rate is almost unaffected for GPT-3.5 and slightly decreases by 2.77% for GPT-4, indicating that SelfDefend incurs negligible effects on normal queries.
 Moreover, SelfDefend incurs zero delay for over 95% of normal prompts across three out of four configurations.
-For multiple jailbreak samples, SelfDefend causes an average extra delay of 0.05 seconds for GPT-3.5 and 0.40 seconds for GPT-4, respectively.
-While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.063, GPT-4 itself is commonly known to be expensive [3].
+For multiple jailbreak samples, SelfDefend causes an average extra delay of 0.06 seconds for GPT-3.5 and 0.35 seconds for GPT-4, respectively.
+While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.050, GPT-4 itself is commonly known to be expensive [3].
 Moreover, the closed-source nature of GPT-3.5/4 raises privacy concerns for non-OpenAI model vendors.
@@
 The details will be introduced in §6.
-To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with six representative jailbreak defenses: ICD [73], SafeDecoding [82], Perplexity Filter [26], SmoothLLM [59], Llama Guard [25], and Llama Guard 2 [66].
-The results show that SelfDefend consistently outperforms all other defenses in 20 out of 24 tested attack scenarios (involving eight jailbreak methods and three target LLMs: GPT-3.5, GPT-4, and Llama-2) and reaches the level of GPT-4-based SelfDefend’s defense.
+To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with seven representative jailbreak defenses: ICD [75], SafeDecoding [84], Perplexity Filter [27], SmoothLLM [60], Llama Guard [26], and Llama Guard 2/3 [67, 19].
+The results show that SelfDefend consistently outperforms all other defenses in 55 out of 60 tested attack scenarios (involving 10 jailbreak methods and six target LLMs: GPT-3.5, GPT-4, Llama-2-7b-chat, Mistral-7B-Instruct-v0.2, Claude-3.5-sonnet, and Llama-2-13b-chat) and reaches the defense level of GPT-4-based SelfDefend.
 Besides defense effectiveness, we also measure the extra delay $\Delta d$ caused by our defense models and find that the average $\Delta d$ is negligible, at 0-0.01 seconds for normal prompts.
-For attack scenarios, the maximum $\Delta d$ has decreased from 1.56 seconds in GPT-4-based SelfDefend to 0.27 seconds, with $\Delta d$ in all attack scenarios except for DAN now below 0.1 seconds.
+For attack scenarios, the maximum $\Delta d$ has decreased from 1.56 seconds in GPT-4-based SelfDefend to 0.39 seconds, with $\Delta d$ in all attack scenarios now below 0.1 seconds except for DAN and LLM-Fuzzer.
 These findings indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it efficient for potential real-world deployment.
@@
 Unlike optimizing these obviously unreadable suffixes, AutoDAN [41] automatically constructs human-readable jailbreak prompts using a carefully designed hierarchical genetic algorithm.
+Furthermore, RLbreaker [14] trains a reinforcement learning agent to guide the search for adversarial prompts, making it more efficient than the stochastic mutations of JSAA and AutoDAN.
 Generation-based Jailbreak employs language models [55, 17, 13, 48, 54] to generate effective jailbreak prompts that can mislead LLMs into producing restricted content.
@@
 [48] then refined PAIR’s approach through tree-of-thought reasoning [86].
+LLM-Fuzzer [88] automates jailbreak template generation for LLMs by starting with human-written templates and applying random mutations to create new inputs with the assistance of LLMs.
 Moreover, the adversary can train a new LLM specifically to attack the target model.
@@
 GradSafe [81] ACL 2024 (2402.13494) Compare the prompt’s gradient similarity with safety-critical gradients
-Gradient Cuff [23] arXiv:2403.00867 Compare the gradient norm of refusal loss with that in benign queries
-Circuit Breaking [98] arXiv:2406.04313 Interrupt the model output harmful content at the internal representations
+Gradient Cuff [24] NeurIPS 2024 (2403.00867) Compare the gradient norm of refusal loss with that in benign queries
+Circuit Breaking [101] NeurIPS 2024 (2406.04313) Interrupt the model output harmful content at the internal representations
 RAIN [36] ICLR 2024 (2309.07124) Self-evaluate each token output, rewind, and determine the final output
 Goal Prioritization [92] ACL 2024 (2311.09096) Integrate instructions to control the priority between helpfulness and safety
-RPO [96], DPP [81] arXiv:2401.17263 Optimize universal and transferable suffixes that enforce safe outputs
+RPO [99], DPP [83] NeurIPS 2024 (2401.17263) Optimize universal and transferable suffixes that enforce safe outputs
 DRO [96], PAT [50] ICML 2024 (2401.18018) Add a prefix as safety prompt and optimize it with prompt tuning
@@
 Eraser [46] arXiv:2404.05880 Encourage LLMs to forget harmful knowledge via machine unlearning
-CAT [77] arXiv:2405.15589 Conduct adversarial training on continuous embedding attacks
-LED [91] arXiv:2405.18166 Identify safety-critical layers and realign them through model editing
+CAT [79] NeurIPS 2024 (2405.15589) Conduct adversarial training on continuous embedding attacks
+LED [94] EMNLP 2024 (2405.18166) Identify safety-critical layers and realign them through model editing
 Adversarial Tuning [40] arXiv:2406.06622 Fine-tune the LLM with token- and semantic-level adversarial prompts
@@
 Our key idea is to deploy a dedicated LLM alongside the target LLM to concurrently detect potential jailbreak queries.
-This idea is made possible because we found that LLMs do have the capability to protect themselves by identifying harmful portions in user queries.
+This idea is made possible because we found that LLMs can protect themselves by identifying harmful portions in user queries.
+Insight.
+Normally, an LLM operates in the answering state to follow a user prompt query $P_{query}$ and return the corresponding answer response $A_{response}$.
+To ensure the safety of $A_{response}$, existing guardrail approaches such as Llama Guard [26] and LLM Self Defense [56] employ a model or a system prompt to assess the harmfulness of $A_{response}$ and filter it if it violates safety policies.
+Such an approach requires waiting for $A_{response}$ to be generated by the LLM.
+Our insight is that a target LLM could operate not only in the answering state but also in the detection state simultaneously, as long as we create two instances of the target model.
+Therefore, given the same $P_{query}$, we aim to initialize two states of the LLM at the same time, one still answering $P_{query}$ normally but the other cautiously checking $P_{query}$ (instead of answering it).
+This is a new perspective that has never been explored by previous works.
+Indeed, our measurements in §5 empirically demonstrate a significant discrepancy between the answering state and the detection state for the same LLM, with the median ASR gap reaching 2.29$\times$ for GPT-3.5 and 8.00$\times$ for GPT-4, as shown in the two rows labeled “Gap” in Table 3.
 Overview.
-Based on this idea, we propose a generic LLM jailbreak defense framework called SelfDefend.
+Based on this insight, we propose a generic LLM jailbreak defense framework called SelfDefend.
 As shown in Figure 1, SelfDefend creatively establishes a shadow stack alongside the normal stack in the LLM space to conduct checkpoint-based access control, which mimics traditional security defense concepts such as the shadow stack for defending against buffer overflow attacks [9] and the library-based checkpoint from SCLib [76].
@@
 By contrast, $LLM_{defense}$ employs a tailored detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful part (via $P_{direct}$) or intention (via $P_{intent}$).
-③ Once a token of “No” (indicating no issue) is output from the shadow stack, $LLM_{defense}$ is triggered to release its token-by-token response.
+③ Once a token of “No” (indicating no issue) is output from the shadow stack, $LLM_{target}$ is triggered to release its token-by-token response.
 Otherwise, when the shadow stack detects a harmful portion (prompt/intention), SelfDefend would respond with a template that refuses to answer, i.e., “I can’t fulfill your query because your [harmful portion] violated our safety policy.” It is noted that “[harmful portion]” is replaced with the recognized portion from $LLM_{defense}$.
@@
 As mentioned in §4, $P_{direct}$ and $P_{intent}$ represent two different defense schemes.
-$P_{direct}$ is more intuitive, aiming to directly pinpoint the harmful part from the original jailbreak prompt; otherwise, it outputs a token of “No”.
-$P_{intent}$, on the other hand, aims to recognize the true intent of the request text rather than directly intercepting harmful parts, given that indirect jailbreak attacks (e.g., DrAttack [33]) hide harmful instructions.
-Based on these two guidelines, we carefully design the prompt templates used by $P_{direct}$ and $P_{intent}$.
-As shown in Figure 2, $P_{direct}$ uses the sentence like “identify which part of the following sentences might violate your safety policies or have the jailbreak intentions” to directly pinpoint the harmful portion.
-Inspired by chain-of-thought (CoT) [71], $P_{intent}$ first summarizes the true intention of $P_{query}$ and then, further extracts any safety-violation part from the summarized intention, if present.
-It also includes two contrasting examples in the prompts to help $LLM_{defense}$ better understand the instructions, motivated by few-shot in-context learning [7].
-Through the extensive measurements in this section, these two designed prompts are shown to be effective within the SelfDefend framework.
-Nevertheless, we note that the detailed wording of these two prompt contents might have alternatives, but the current form is used as a standard template for $P_{direct}$ and $P_{intent}$ throughout the paper.
+One directly identifies the harmful part from the original jailbreak prompt, while the other indirectly recognizes the true intent of the request text.
+Based on these two guidelines, we carefully design the prompt templates of $P_{direct}$ and $P_{intent}$ as follows:
+The design of $P_{direct}$ is based on the fact that many jailbreak attacks explicitly contain instructions to make LLMs respond with corresponding harmful content.
+Benefiting from the powerful semantic understanding ability of LLMs, we design $P_{direct}$ with the sentence “identify which part of the following sentences might violate your safety policies or have the jailbreak intentions” to directly pinpoint the harmful portion in the query text, as shown in Figure 2.
+The excerpted harmful portion indicates that the request is a jailbreak.
+Otherwise, $LLM_{defense}$ outputs a token of “No” when the query is safe.
+The design of $P_{intent}$, on the other hand, considers that indirect jailbreak attacks (e.g., DrAttack [34]) may not explicitly contain harmful instructions.
+As such, $P_{intent}$ is designed to understand the true intention of the query and determine whether it contains any safety-violating content, as shown in Figure 2.
+Inspired by chain-of-thought (CoT) [73], $P_{intent}$ first summarizes the true intention of $P_{query}$ and then extracts any safety-violating part from the summarized intention, if present.
+If the query is safe, $LLM_{defense}$ also outputs “No”.
+Moreover, $P_{intent}$ includes two contrasting examples in the prompts to help $LLM_{defense}$ better understand the instructions, motivated by few-shot in-context learning (ICL) [8].
+Both COT and ICL learning techniques have been shown to be effective in enhancing the reasoning ability of LLMs [73, 8].
+Measurements conducted in this section demonstrate the effectiveness of these two prompts.
+While alternative wordings may exist, the current formulations for $P_{direct}$ and $P_{intent}$ serve as standardized templates throughout this paper.
 5.2 Datasets and Attack Setup
-After SelfDefend is set up, we collect benchmark datasets, implement the corresponding attack methods, and determine the effectiveness metrics.
 Benchmarks.
@@
 Table 2 lists the details of our collected benchmark datasets.
-Specifically, we use a set of 100 harmful instructions from JailbreakBench [11], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [99] and AutoDAN [40]), generation-based jailbreaks (PAIR [12] and TAP [47]), and indirect jailbreak attacks (DrAttack [33] and Puzzler [10]).
+Specifically, we use a set of 100 harmful instructions from JailbreakBench [12], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [102], AutoDAN [41], and RLbreaker [14]), generation-based jailbreaks (PAIR [13], TAP [48], and LLM-Fuzzer [88]), and indirect jailbreak attacks (DrAttack [34] and Puzzler [11]).
 In contrast, we directly use the original prompts from the JailbreakHub, MultiJail and AlpacaEval datasets to construct inputs for the scenarios of human-based attacks, multilingual jailbreaks and normal prompts, respectively.
@@
 JailbreakHub [63] 1000 DAN [63]
-JailbreakBench [11] 100 GCG [99], AutoDAN [40]
-PAIR [12], TAP [47]
+JailbreakBench [12] 100 GCG [102], AutoDAN [41], RLbreaker [14]
+PAIR [13], TAP [48], LLM-Fuzzer [88]
 DrAttack [34], Puzzler [11]
@@
 LLMs Human-based Optimization-based Generation-based Indirect Multilingual Normal
-DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail AlpacaEval
-GPT-3.5 (baseline) 0.256 0.560 0.900 0.720 0.670 0.780 0.980 0.393 0.977
-GPT-3.5-based Shadow Stack ($P_{direct}$) 0.982 0.720 0.960 0.770 0.840 1.000 0.980 0.879 0.977
-GPT-3.5-based SelfDefend ($P_{direct}$) 0.242 0.450 0.870 0.600 0.610 0.780 0.960 0.368 0.957
-GPT-3.5-based Shadow Stack ($P_{intent}$) 0.015 0.280 0.350 0.370 0.020 0.860 0.220 0.520 0.992
-GPT-3.5-based SelfDefend ($P_{intent}$) 0.007 0.190 0.310 0.290 0.020 0.710 0.220 0.203 0.972
-Reduction factor 97.26% 66.07% 65.55% 59.72% 97.01% 8.97% 77.55% 48.34% 0.51%
-GPT-4 (baseline) 0.047 0.080 0.190 0.330 0.310 0.740 0.900 0.076 0.973
-GPT-4-based Shadow Stack ($P_{direct}$) 0.004 0.010 0.010 0.110 0.100 0.050 0.270 0.142 0.968
-GPT-4-based SelfDefend ($P_{direct}$) 0.002 0.000 0.010 0.100 0.080 0.040 0.260 0.012 0.946
-Reduction factor 95.74% 100.00% 94.73% 69.69% 74.19% 94.59% 71.11% 84.21% 2.77%
-GPT-4-based Shadow Stack ($P_{intent}$) 0.019 0.080 0.070 0.210 0.200 0.130 0.360 0.304 0.995
-GPT-4-based SelfDefend ($P_{intent}$) 0.005 0.050 0.070 0.180 0.170 0.130 0.280 0.019 0.970
-(a) The CDF plot of $\Delta d$ for normal prompts.
-(b) Average $\Delta d$ for various jailbreak prompts.
-Figure 3: The extra delay $\Delta d$ caused by SelfDefend for normal and jailbreak prompts, respectively.
+DAN GCG AutoDAN RLbreaker PAIR TAP LLM-Fuzzer DrAttack Puzzler MultiJail AlpacaEval
+GPT-3.5 (baseline) 0.256 0.560 0.900 0.650 0.720 0.670 0.640 0.780 0.980 0.393 0.977
+GPT-3.5-based Shadow Stack ($P_{direct}$) 0.982 0.720 0.960 0.910 0.770 0.840 0.790 1.000 0.980 0.879 0.977
+GPT-3.5-based SelfDefend ($P_{direct}$) 0.242 0.450 0.870 0.600 0.600 0.610 0.500 0.780 0.960 0.368 0.957
+GPT-3.5-based Shadow Stack ($P_{intent}$) 0.015 0.280 0.350 0.310 0.370 0.020 0.280 0.860 0.220 0.520 0.992
+Gap between Normal and Shadow ($P_{intent}$) 17.07$\times$ 2.00$\times$ 2.57$\times$ 2.10$\times$ 1.95$\times$ 33.50$\times$ 2.29$\times$ 0.91$\times$ 4.45$\times$ 0.76$\times$ 0.98$\times$
+GPT-3.5-based SelfDefend ($P_{intent}$) 0.007 0.190 0.310 0.240 0.290 0.020 0.170 0.710 0.220 0.203 0.972
+Reduction factor 97.26% 66.07% 65.55% 63.08% 59.72% 97.01% 73.44% 8.97% 77.55% 48.34% 0.51%
+GPT-4 (baseline) 0.047 0.080 0.190 0.290 0.330 0.310 0.190 0.740 0.900 0.076 0.973
+GPT-4-based Shadow Stack ($P_{direct}$) 0.004 0.010 0.010 0.000 0.110 0.100 0.010 0.050 0.270 0.142 0.968
+Gap between Normal and Shadow ($P_{direct}$) 11.75$\times$ 8.00$\times$ 19.00$\times$ $\infty$×$\times$ 3.00$\times$ 3.10$\times$ 19.00$\times$ 14.80$\times$ 3.33$\times$ 0.54$\times$ 1.01$\times$
+GPT-4-based SelfDefend ($P_{direct}$) 0.002 0.000 0.010 0.000 0.100 0.080 0.000 0.040 0.260 0.012 0.946
+Reduction factor 95.74% 100.00% 94.73% 100.00% 69.69% 74.19% 100.00% 94.59% 71.11% 84.21% 2.77%
+GPT-4-based Shadow Stack ($P_{intent}$) 0.019 0.080 0.070 0.050 0.210 0.200 0.050 0.130 0.360 0.304 0.995
+GPT-4-based SelfDefend ($P_{intent}$) 0.005 0.050 0.070 0.000 0.180 0.170 0.040 0.130 0.280 0.019 0.970
 Attack Setup.
@@
 The genetic algorithm used in AutoDAN-GA features a crossover rate of 0.5, a mutation rate of 0.01, and 100 optimization steps.
+For RLbreaker, we employ GPT-3.5 as the auxiliary model to perform mutations.
+The total number of queries to the target LLM is limited to 10,000 for both the training and testing phases.
 For PAIR and TAP, we select Vicuna-13b-v1.5 [15] as the attack model.
@@
 The target model of PAIR and TAP is GPT-3.5/4.
+For LLM-Fuzzer, we select GPT-3.5 as the helper model for mutations and set the maximum number of queries to the target LLMs at 1,000.
 For DrAttack and Puzzler, we use GPT-4 to construct their jailbreak prompts.
@@
 We measure the defense effectiveness of SelfDefend indirectly using the attack success rate (ASR) [102] or unsafe rate [18].
-Following [99, 40], ASR is measured by detecting whether the LLM response contains keywords indicating a refusal to answer, like “I can’t assist.” If such a keyword or its variant is included, it indicates that the attack is unsuccessful; otherwise, it indicates that the jailbreak is successful.
-We use the common practice [99] for a list of such keywords; see Appendix B.
-DAN, GCG, AutoDAN, PAIR, TAP, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance.
+Following [102, 41], ASR is measured by detecting whether the LLM response contains keywords indicating a refusal to answer, such as “I can’t assist.” If such a keyword or its variant is included, it indicates that the attack is unsuccessful; otherwise, vice versa.
+We use the common practice [102] for a list of such keywords; see Appendix D.
+DAN, GCG, AutoDAN, RLbreaker, PAIR, TAP, LLM-Fuzzer, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance.
 The lower the ASR, the stronger the defense performance of SelfDefend against jailbreaks.
@@
 On the other hand, the unsafe rate uses GPT-4 to determine whether the response of the target model matches the jailbreak goal [18].
-We adopt the unsafe rate in a way similar to ASR to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.
+We adopt the unsafe rate in a manner similar to ASR to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.
 5.3 Measurement Results
@@
 As shown in Table 3, these jailbreaks are much more effective on GPT-3.5 compared to GPT-4.
-The ASR for GPT-3.5 can be as high as 0.256 to 0.980 (average: 0.657), whereas the ASR for GPT-4 typically ranges from 0.047 to 0.330, except for indirect jailbreaks where GPT-4 also exhibits a high ASR (0.74 for DrAttack and 0.900 for Puzzler).
+The ASR for GPT-3.5 can be as high as 0.256 to 0.980 (average: 0.655), whereas the ASR for GPT-4 typically ranges from 0.047 to 0.330, except for indirect jailbreaks where GPT-4 also exhibits a high ASR (0.74 for DrAttack and 0.900 for Puzzler).
 Defense Effectiveness.
-We then compare ASR after equipping the baseline model with SelfDefend and report additional ASR findings when using only the shadow stack for ablation exploration.
+We then compare ASR after equipping the baseline model with SelfDefend and report additional ASR findings when using only the shadow stack for ablation.
 For GPT-3.5, both SelfDefend versions—supported by $P_{direct}$ and $P_{intent}$—reduce ASRs but show varying degrees of defense enhancement.
@@
 Considering that the ASR under SelfDefend is still very high and close to the baseline, we conclude that SelfDefend’s defense would not notably impact normal user queries.
+Figure 3: The CDF plot of $\Delta d$ for normal prompts.
+Figure 4: The training procedure for fine-tuning our open-source defense models.
 Extra Delay $\Delta d$.
@@
 We thus further measure the extra delay $\Delta d$ caused by SelfDefend, which is calculated as follows:
-[…]
+$\Delta d=d_{total}-d_{normal},$ (1)
 where $d_{total}$ denotes the total delay to SelfDefend-protected LLMs, and $d_{normal}$ represents the separate time cost of the target LLM in the normal stack.
 The metric of $\Delta d$ is essential for normal prompts because it would directly affect the experience of normal users.
-Figure 3(a) plots the cumulative distribution function (CDF) of $\Delta d$ for normal prompts.
+Figure 3 plots the cumulative distribution function (CDF) of $\Delta d$ for normal prompts.
 We can see that over 95% of the cases incur zero delays across three configurations of SelfDefend, demonstrating that SelfDefend is capable of defending against jailbreaks at the cost of negligible delay for normal users.
 The only exception occurs with SelfDefend using $P_{intent}$ and GPT-3.5, where around 80% of the cases have no extra delay.
-For jailbreak prompts, we also calculate their average $\Delta d$ for different attack scenarios in Figure 3(b).
-We observe that GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all eight scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval.
+For jailbreak prompts, we also calculate their average $\Delta d$ for different attack scenarios in Figure 5 (see Appendix B).
+GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all 10 scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval.
 This is likely because the baseline GPT-4 has good defense performance for most jailbreaks, with its original output already short and the delay minimal.
-Figure 4: The training procedure for fine-tuning our open-source defense model.
 6 Tuning a Dedicated Defense Model
@@
 During the training process, we continue to use our defense prompts and queries from the collected datasets as inputs; that is, we take $P_{dir}[x]$ or $P_{int}[x]$ as input and $y$ as the label to fine-tune the Llama model.
-Table 4: Jailbreak ASR for various defense methods.
-For ICD, SmoothLLM and SafeDecoding, we present the performance of their enhanced models.
-For detection-based Perplexity Filter and Llama Guard (2), we report ASRs only on their detection modules.
-Since SafeDecoding works for a white-box target model, we show its results only on the publicly available Llama-2.
-Target Defense Method Human Optimization Generation Indirect Multilingual Normal
-DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail AlpacaEval
-GPT-3.5 GPT-3.5 (baseline) 0.256 0.560 0.900 0.720 0.670 0.780 0.980 0.393 0.977
-ICD [73] 0.226 0.230 0.840 0.360 0.330 0.750 0.990 0.321 0.960
-Perplexity Filter [26] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 0.994
-SmoothLLM [59] 0.238 0.480 0.930 0.650 0.610 0.850 0.990 0.705 0.968
-Llama Guard [25] 0.561 0.410 0.580 0.430 0.470 0.970 0.930 0.952 0.996
-Llama Guard 2 [66] 0.441 0.140 0.150 0.370 0.360 0.890 0.640 0.559 0.991
-$P_{direct}$-tuned Shadow Stack 0.262 0.080 0.070 0.140 0.210 0.780 0.070 0.749 0.968
-$P_{direct}$-tuned SelfDefend 0.111 0.060 0.070 0.070 0.170 0.620 0.070 0.302 0.948
-$P_{intent}$-tuned Shadow Stack 0.297 0.080 0.090 0.160 0.240 0.180 0.200 0.578 0.996
-$P_{intent}$-tuned SelfDefend 0.125 0.050 0.080 0.120 0.200 0.160 0.200 0.260 0.975
-Double shadow stack 0.213 0.040 0.050 0.100 0.130 0.180 0.070 0.470 0.966
-Double shadow stack+GPT-3.5 0.091 0.030 0.050 0.060 0.100 0.160 0.070 0.187 0.947
-GPT-4 GPT-4 (baseline) 0.047 0.080 0.190 0.330 0.310 0.740 0.900 0.076 0.973
-ICD [73] 0.062 0.050 0.030 0.230 0.230 0.430 0.640 0.051 0.970
-Perplexity Filter [26] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 0.994
-SmoothLLM [59] 0.030 0.070 0.180 0.330 0.330 0.910 0.880 0.651 0.971
-Llama Guard [25] 0.561 0.410 0.580 0.430 0.400 0.980 0.930 0.952 0.996
-Llama Guard 2 [66] 0.441 0.140 0.150 0.350 0.330 0.910 0.640 0.559 0.991
-$P_{direct}$-tuned Shadow Stack 0.262 0.040 0.070 0.170 0.220 0.710 0.120 0.717 0.969
-$P_{direct}$-tuned SelfDefend 0.032 0.010 0.050 0.150 0.150 0.580 0.070 0.060 0.947
-$P_{intent}$-tuned Shadow Stack 0.284 0.080 0.070 0.190 0.190 0.180 0.190 0.565 0.995
-$P_{intent}$-tuned SelfDefend 0.034 0.040 0.060 0.190 0.160 0.170 0.140 0.044 0.970
-Double shadow stack 0.198 0.020 0.040 0.120 0.130 0.180 0.120 0.467 0.968
-Double shadow stack+GPT-4 0.026 0.010 0.040 0.120 0.110 0.170 0.070 0.038 0.945
-Llama-2 Llama-2-7b-chat (baseline) 0.678 0.570 0.680 0.590 0.610 0.880 0.990 0.143 0.988
-ICD [73] 0.474 0.700 0.560 0.310 0.320 0.220 1.000 0.946 0.898
-SafeDecoding [82] 0.655 0.550 0.740 0.560 0.590 0.640 1.000 0.857 0.981
-Perplexity Filter [26] 1.000 0.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
-SmoothLLM [59] 0.681 0.930 0.950 0.820 0.810 0.980 1.000 1.000 0.993
-Llama Guard [25] 0.561 0.400 0.580 0.460 0.420 0.840 0.930 0.952 0.996
-Llama Guard 2 [66] 0.441 0.170 0.150 0.410 0.350 0.890 0.640 0.559 0.991
-(7b-chat) $P_{direct}$-tuned Shadow Stack 0.257 0.060 0.050 0.250 0.190 0.360 0.090 0.737 0.970
-$P_{direct}$-tuned SelfDefend 0.214 0.040 0.050 0.220 0.180 0.310 0.090 0.102 0.959
-$P_{intent}$-tuned Shadow Stack 0.289 0.110 0.080 0.240 0.140 0.150 0.220 0.587 0.991
-$P_{intent}$-tuned SelfDefend 0.242 0.090 0.070 0.210 0.140 0.150 0.220 0.063 0.980
-Double shadow stack 0.198 0.050 0.040 0.180 0.140 0.120 0.040 0.479 0.965
-Double shadow stack+Llama-2 0.164 0.030 0.040 0.160 0.140 0.110 0.040 0.057 0.954
 7 Evaluation
@@
 Target LLMs, Benchmarks, and Environment.
-As shown in Table 4, our tuned defense models are paired with protect three popular LLMs: GPT-3.5, GPT-4, and the open-source Llama-2-7b-chat [67].
-We use the same benchmark datasets as in Table 2 (see §5).
-We follow the same procedure and configuration as in §5 to generate the tested jailbreak prompts for GPT-3.5 and GPT-4.
-The process is similar for testing the target model Llama-2-7b-chat; we set the target model for query-based PAIR, TAP and DrAttack to Llama-2-7b-chat to generate their jailbreak prompts.
+As shown in Table 4, our tuned defense models are tested to protect six popular LLMs: the proprietary GPT-3.5, GPT-4, Claude-3.5-sonnet [6], the open-source Llama-2-chat [69] with 7B and 13B sizes, and Mistral-7B-Instruct-v0.2 [68], covering diverse model architectures and sizes.
+Due to page limitation, the results for Claude-3.5 and Llama-2-13b-chat are reported in Appendix A.
+We use the same benchmark datasets as in Table 2 and follow the similar procedures and configurations as in §5 to generate the tested jailbreak prompts.
+For query-based attacks (RLbreaker, PAIR, TAP, LLM-Fuzzer, and DrAttack), we also generate jailbreaks for the corresponding target LLMs.
 Our evaluations are implemented using PyTorch 2.2.1 and conducted on an NVIDIA TESLA H800 GPU.
 Baselines.
-We compare our framework with popular jailbreak defense methods, including ICD [73], SafeDecoding[82], Perplexity Filter [26], SmoothLLM [59], Llama Guard [25] and Llama Guard 2 [66].
+We compare our framework with popular jailbreak defense methods, including ICD [75], SafeDecoding[84], Perplexity Filter [27], SmoothLLM [60], Llama Guard [26], Llama Guard 2/3 [67, 19].
 Specifically, ICD adds in-context demonstrations in input prompts to enhance the safety of the target model.
@@
 Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts.
-Llama Guard 2 is a fine-tuned version of Llama-3-8B [18].
+Llama Guard 2 and 3 are similar, but fine-tuned on Llama-3 and Llama-3.1 (8B), respectively.
 Metric.
@@
 For our fine-tuned models, the success represents their response being “No”.
-Figure 5: Responses of defense models against a jailbreak prompt of AutoDAN.
+Table 4: Jailbreak ASR for various defense methods.
+For ICD and SafeDecoding, we present the performance of their enhanced models.
+For detection-based Perplexity Filter, SmoothLLM and Llama Guards, we report ASRs only on their detection modules.
+Since SafeDecoding works for a white-box target model, we show its results on the publicly available Llama-2 and Mistral.
+Target Defense Method Human Optimization Generation Indirect Multilingual Normal
+Model DAN GCG AutoDAN RLbreaker PAIR TAP LLM-Fuzzer DrAttack Puzzler MultiJail AlpacaEval
+GPT-3.5 GPT-3.5 (baseline) 0.256 0.560 0.900 0.650 0.720 0.670 0.640 0.780 0.980 0.393 0.977
+ICD [75] 0.226 0.230 0.840 0.140 0.360 0.330 0.390 0.750 0.990 0.321 0.960
+Perplexity Filter [27] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
+SmoothLLM [60] 0.238 0.480 0.930 0.320 0.650 0.610 0.490 0.850 0.990 0.267 0.968
+Llama Guard [26] 0.561 0.410 0.580 0.790 0.430 0.470 0.710 0.970 0.930 0.952 0.996
+Llama Guard 2 [67] 0.441 0.140 0.150 0.410 0.370 0.360 0.180 0.890 0.640 0.559 0.991
+Llama Guard 3 [19] 0.343 0.080 0.130 0.110 0.230 0.290 0.080 0.610 0.420 0.378 0.986
+$P_{direct}$-tuned Shadow Stack 0.262 0.080 0.070 0.040 0.140 0.210 0.050 0.780 0.070 0.749 0.968
+$P_{direct}$-tuned SelfDefend 0.111 0.060 0.070 0.040 0.070 0.170 0.030 0.620 0.070 0.302 0.948
+$P_{intent}$-tuned Shadow Stack 0.297 0.080 0.090 0.050 0.160 0.240 0.040 0.180 0.200 0.578 0.996
+$P_{intent}$-tuned SelfDefend 0.125 0.050 0.080 0.050 0.120 0.200 0.030 0.160 0.200 0.260 0.975
+Double shadow stack 0.213 0.040 0.050 0.010 0.100 0.130 0.020 0.180 0.070 0.470 0.966
+Double shadow stack+GPT-3.5 0.091 0.030 0.050 0.010 0.060 0.100 0.010 0.160 0.070 0.187 0.947
+GPT-4 GPT-4 (baseline) 0.047 0.080 0.190 0.290 0.330 0.310 0.190 0.740 0.900 0.076 0.973
+ICD [75] 0.062 0.050 0.030 0.010 0.230 0.230 0.050 0.430 0.640 0.051 0.970
+Perplexity Filter [27] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
+SmoothLLM [60] 0.030 0.070 0.180 0.040 0.330 0.330 0.220 0.910 0.880 0.048 0.971
+Llama Guard [26] 0.561 0.410 0.580 0.660 0.430 0.400 0.700 0.980 0.930 0.952 0.996
+Llama Guard 2 [67] 0.441 0.140 0.150 0.340 0.350 0.330 0.150 0.910 0.640 0.559 0.991
+Llama Guard 3 [19] 0.343 0.080 0.130 0.090 0.330 0.220 0.110 0.590 0.420 0.378 0.986
+$P_{direct}$-tuned Shadow Stack 0.262 0.040 0.070 0.030 0.170 0.220 0.030 0.710 0.120 0.717 0.969
+$P_{direct}$-tuned SelfDefend 0.032 0.010 0.050 0.010 0.150 0.150 0.020 0.580 0.070 0.060 0.947
+$P_{intent}$-tuned Shadow Stack 0.284 0.080 0.070 0.060 0.190 0.190 0.040 0.180 0.190 0.565 0.995
+$P_{intent}$-tuned SelfDefend 0.034 0.040 0.060 0.020 0.190 0.160 0.030 0.170 0.140 0.044 0.970
+Double shadow stack 0.198 0.020 0.040 0.010 0.120 0.130 0.010 0.180 0.120 0.467 0.968
+Double shadow stack+GPT-4 0.026 0.010 0.040 0.000 0.120 0.110 0.000 0.170 0.070 0.038 0.945
+Llama-2 Llama-2-7b-chat (baseline) 0.678 0.570 0.680 0.490 0.590 0.610 0.120 0.880 0.990 0.143 0.988
+ICD [75] 0.474 0.700 0.560 0.310 0.310 0.320 0.630 0.220 1.000 0.146 0.898
+SafeDecoding [84] 0.655 0.550 0.740 0.630 0.560 0.590 0.610 0.640 1.000 0.857 0.981
+Perplexity Filter [27] 1.000 0.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
+SmoothLLM [60] 0.681 0.930 0.950 0.780 0.820 0.810 0.870 0.980 1.000 0.130 0.993
+Llama Guard [26] 0.561 0.400 0.580 0.480 0.460 0.420 0.640 0.840 0.930 0.952 0.996
+Llama Guard 2 [67] 0.441 0.170 0.150 0.300 0.410 0.350 0.280 0.890 0.640 0.559 0.991
+Llama Guard 3 [19] 0.343 0.090 0.130 0.090 0.310 0.280 0.130 0.410 0.420 0.378 0.986
+(7b-chat) $P_{direct}$-tuned Shadow Stack 0.257 0.060 0.050 0.020 0.250 0.190 0.020 0.360 0.090 0.737 0.970
+$P_{direct}$-tuned SelfDefend 0.214 0.040 0.050 0.020 0.220 0.180 0.010 0.310 0.090 0.102 0.959
+$P_{intent}$-tuned Shadow Stack 0.289 0.110 0.080 0.030 0.240 0.140 0.040 0.150 0.220 0.587 0.991
+$P_{intent}$-tuned SelfDefend 0.242 0.090 0.070 0.020 0.210 0.140 0.010 0.150 0.220 0.063 0.980
+Double shadow stack 0.198 0.050 0.040 0.010 0.180 0.140 0.000 0.120 0.040 0.479 0.965
+Double shadow stack+Llama-2 0.164 0.030 0.040 0.010 0.160 0.140 0.000 0.110 0.040 0.057 0.954
+Mistral Mistral-7B-Instruct-v0.2 (baseline) 0.685 0.930 0.990 0.410 0.780 0.730 0.450 0.760 0.990 0.276 0.970
+ICD [75] 0.679 0.680 0.980 0.430 0.740 0.750 0.810 0.630 0.990 0.286 0.932
+SafeDecoding [84] 0.818 0.930 0.970 0.690 0.830 0.780 0.900 0.690 0.990 0.883 0.979
+Perplexity Filter [27] 1.000 0.110 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
+SmoothLLM [60] 0.729 0.980 1.000 0.690 0.920 0.850 0.800 0.930 0.990 0.276 0.994
+Llama Guard [26] 0.552 0.390 0.990 0.810 0.440 0.410 0.420 0.870 0.930 0.952 0.996
+Llama Guard 2 [67] 0.441 0.180 0.110 0.310 0.340 0.350 0.310 0.880 0.620 0.559 0.990
+Llama Guard 3 [19] 0.343 0.150 0.140 0.070 0.150 0.290 0.130 0.490 0.420 0.378 0.986
+(7B-Instruct $P_{direct}$-tuned Shadow Stack 0.260 0.060 0.050 0.020 0.120 0.220 0.050 0.350 0.120 0.708 0.968
+-v0.2) $P_{direct}$-tuned SelfDefend 0.192 0.060 0.050 0.000 0.120 0.210 0.010 0.300 0.120 0.178 0.939
+$P_{intent}$-tuned Shadow Stack 0.297 0.070 0.060 0.030 0.100 0.210 0.040 0.060 0.240 0.600 0.993
+$P_{intent}$-tuned SelfDefend 0.226 0.070 0.060 0.020 0.080 0.210 0.000 0.060 0.230 0.140 0.964
+Double shadow stack 0.208 0.030 0.050 0.010 0.060 0.180 0.030 0.050 0.070 0.498 0.965
+Double shadow stack+Mistral 0.151 0.030 0.050 0.000 0.060 0.180 0.000 0.050 0.070 0.121 0.937
 7.2 Defense Effectiveness
@@
 As shown in Table 4, our tuned models under SelfDefend achieve satisfactory defense effects under all types of jailbreaks and deliver state-of-the-art (SOTA) results in most cases, while previous defenses fail to show promising performance under different attacks.
-For ICD, the tested ASRs under Puzzler are not less than 99% for both GPT-3.5 and Llama-2.
+For ICD, the tested ASRs under Puzzler are not less than 99% for GPT-3.5, Llama-2 and Mistral.
 For SafeDecoding, it also shows weakness against Puzzler and MultiJail.
@@
 Although it exhibits slightly better ASR against GCG than our methods for GPT-3.5 and Llama-2, this is because its filter threshold is set to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix.
-SmoothLLM does not show satisfactory defense performance against DrAttack and Puzzler for all three target models.
-Its ASRs even exceed 80% on Llama-2 for all attack methods except DAN.
+SmoothLLM does not show satisfactory defense performance against DrAttack and Puzzler for all four target models.
 Llama Guard exhibits poor jailbreak detection for DrAttack, Puzzler, and MultiJail, while Llama Guard 2 presents unacceptable flaws for DrAttack.
-Across all attacks and target models, the average ASR of our $P_{direct}$-tuned shadow stack (27.76%) is 37.61% lower than that of Llama Guard (65.37%) and 16.74% lower than that of Llama Guard 2 (44.50%).
-Similarly, the average defense performance of our $P_{intent}$-tuned shadow stack (22.46%) outperforms Llama Guard by 42.91% and Llama Guard 2 by 22.04%.
-Furthermore, SelfDefend achieves ASRs comparable with all other defense methods under AlpacaEval, indicating its compatibility with normal prompts.
+Llama Guard 3, benefiting from fine-tuning on Llama-3.1-8B, shows better performance than Llama Guard 1 and 2 but is still inferior overall to our fine-tuned defense models.
+Across all attacks and target models, the average ASR of our $P_{direct}$-tuned shadow stack (22.03%) is 44.05% lower than that of Llama Guard (66.08%), 19.07% lower than that of Llama Guard 2 (41.10%), and 4.23% lower than that of Llama Guard 3 (26.26%).
+Similarly, the average defense performance of our $P_{intent}$-tuned shadow stack (18.39%) outperforms Llama Guard by 47.69%, Llama Guard 2 by 22.71%, and Llama Guard 3 by 7.87%.
+Based on the same base model, Llama-2-7b, our method shows a remarkable improvement over Llama Guard 1.
+Even though Llama Guard 2 and 3 are fine-tuned on the more advanced Llama-3-8B and Llama-3.1-8B, respectively, our fine-tuned model still outperforms them.
+This advantage mainly stems from the fact that our approach is more fundamental, extracting harmful objectives from jailbreak prompts, rather than directly assessing the safety of the entire content like the Llama Guard series.
+Table 5: The CLIP-Score [59, 95, 71] ($\uparrow$ indicates better) based on Table 4’s results for Llama-2.
+Tuned Model Text Human-based Optimization-based Generation-based Indirect Multilingual
+DAN GCG AutoDAN RLbreaker PAIR TAP LLM-Fuzzer DrAttack Puzzler MultiJail
+$P_{direct}$-tuned model Generated Attack Prompts 0.685 0.851 0.784 0.732 0.828 0.827 0.673 0.737 0.663 0.677
+Identified Harmful Prompts 0.932 0.946 0.939 0.946 0.898 0.900 0.979 0.785 0.728 0.687
+$P_{intent}$-tuned model Generated Attack Prompts 0.686 0.852 0.784 0.737 0.831 0.827 0.671 0.732 0.667 0.669
+Identified Harmful Intentions 0.919 0.909 0.908 0.899 0.875 0.874 0.901 0.857 0.788 0.766
 $P_{direct}$-tuned SelfDefend v.s.
@@
 In general, SelfDefend with the direct prompt performs better than SelfDefend with the intent prompt on human-based, optimization-based, and generation-based jailbreaks, while $P_{intent}$-tuned SelfDefend is more effective against indirect and multilingual attacks.
-DAN, GCG, AutoDAN, PAIR, and TAP contain explicitly harmful instructions, whereas DrAttack, Puzzler, and MultiJail do not.
+This is because DAN, GCG, AutoDAN, RLbreaker, PAIR, TAP, and LLM-Fuzzer contain explicit harmful instructions, whereas DrAttack, Puzzler, and MultiJail do not.
 This phenomenon aligns with our original intent for designing the two prompts, since $P_{direct}$ is used to directly extract harmful parts, and $P_{intent}$ can identify the true intentions hidden by implicit attacks.
@@
 Moreover, the false positive rate of the $P_{intent}$-tuned SelfDefend on normal prompts is lower than that of the $P_{direct}$-tuned SelfDefend.
+Certain Outlier Cases.
+We observe that some defense methods exhibit better performance than our frameworks in certain outlier cases.
+These cases are typically due to the specific design of the defense methods.
+For example, Perplexity Filter achieves the lowest ASR against GCG on GPT-3.5 and Llama-2, as it sets the filter threshold to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix.
+SmoothLLM performs the lowest ASR (0.030) against DAN on GPT-4.
+Since DAN itself can only achieve an ASR of 0.047 on GPT-4, it is not surprising that SmoothLLM, which votes by accessing the target LLM multiple times, has a slightly lower ASR.
+Additionally, ICD achieves the lowest ASR against AutoDAN on GPT-4, but we also observe that ICD is much more effective in weakening attacks on GPT-4 than on other target models, which indicates the success of ICD on the powerful GPT-4 by adding in-context demonstrations to enhance the safety of the target model.
+In addition, AutoDAN adds the harmful goal at the end of the jailbreak prompt, and therefore, ICD shows a strong rejection rate for AutoDAN.
 Double Shadow Stacks.
@@
 In practice, we suggest using this paradigm only when the standard form of SelfDefend offers low detection accuracy.
-Case Study.
+Case Studies.
 Besides the quantitative analysis above, we conduct case studies to understand in depth the advantages and pitfalls of SelfDefend compared to other defenses.
-Figure 5 illustrates an example of defense methods against a jailbreak prompt.
-Due to page limitation, readers may refer to Appendix A for its detailed analysis, as well as other cases.
-Table 5: The CLIP-Score [58, 92, 69] ($\uparrow$ indicates better) based on Table 4’s results for Llama-2.
-Tuned Model Text Human-based Optimization-based Generation-based Indirect Multilingual
-DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail
-$P_{direct}$-tuned model Generated Attack Prompts 0.685 0.851 0.784 0.828 0.827 0.737 0.663 0.677
-Identified Harmful Prompts 0.932 0.946 0.939 0.898 0.900 0.785 0.728 0.687
-$P_{intent}$-tuned model Generated Attack Prompts 0.686 0.852 0.784 0.831 0.827 0.732 0.667 0.669
-Identified Harmful Intentions 0.919 0.909 0.908 0.875 0.874 0.857 0.788 0.766
-Figure 6: Average $\Delta d$ introduced by the tuned models on Llama-2-7b-chat.
+Due to page limitation, readers may refer to Appendix C.
 7.3 Extra Delay $\Delta d$
+Extra Delay $\Delta d$ Across Different Queries.
 Similar to how we measured the extra delay $\Delta d$ introduced by GPT-based SelfDefend in §5.3, we now report the average delay $\Delta d$ introduced by the tuned models on Llama-2-7b-chat in Figure 6.
-Compared to earlier results in Figure 3, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.008 seconds.
+Compared to earlier results in Figure 3 and Figure 5, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.008 seconds.
 In contrast, GPT-3.5-based (GPT-4-based) SelfDefend with $P_{intent}$ incurred an average $\Delta d$ of around 0.072 seconds (0.026 seconds) under AlpacaEval.
 This is likely because the parameters of fine-tuned models are less than those of GPT-3.5/4.
-Besides the normal prompts, the tuned models also significantly reduce the extra delay $\Delta d$ under various jailbreak prompts.
-For example, the maximum $\Delta d$ now decreases from the earlier 1.56 seconds on GPT-4 to 0.27 seconds.
-Indeed, except for DAN, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend.
+Besides normal prompts, the tuned models also significantly reduce the extra delay $\Delta d$ under jailbreak prompts.
+For example, the maximum $\Delta d$ now decreases from the earlier 1.56 seconds on GPT-4 to 0.39 seconds.
+Indeed, except for DAN and LLM-Fuzzer, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend.
 These results indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for potential deployment.
+$\Delta d$ Compared with Existing Defense Methods.
+We further compare the extra delay $\Delta d$ introduced by our tuned models with other defense methods.
+Table 6 shows the mean extra delay of different defense methods across all 11 tested jailbreak/normal queries.
+We observe that the extra delay $\Delta d$ of SelfDefend is significantly superior to that of other defense methods, benefiting from the parallel processing between the normal and shadow stack.
+Because ICD only adds in-context demonstrations to input prompts, it introduces latency closest to that of SelfDefend.
+Perplexity Filter and Llama Guards have much higher delays than SelfDefend, because the target models are required to wait for their jailbreak detection based on LLM inference.
+SafeDecoding has a much more noticeable delay because it requires two LLMs to perform simultaneous reasoning, including the target model and its fine-tuned expert model.
+SmoothLLM has the highest delay among all defenses, as it perturbs the input prompt and accesses the target LLM multiple times.
+Table 6: The mean extra delay $\Delta d$ of different defense methods for all 11 kinds of tested jailbreak/normal inputs.
+Defense Method Extra Delay $\Delta d$ (s)
+$P_{direct}$-tuned SelfDefend 0.032
+ICD 0.056
+$P_{intent}$-tuned SelfDefend 0.077
+Perplexity Filter 0.168
+Llama Guard 2 0.256
+Llama Guard 3 0.285
+Llama Guard 0.510
+SafeDecoding 0.869
+SmoothLLM 21.807
 7.4 Explainability
-Figure 5 demonstrate the efficacy of the tuned models in identifying harmful content within jailbreak prompts.
-Here, we further quantitatively assess whether the identified harmful content actually aligns with the original jailbreak prompt.
-We measure their semantic similarity using the ensemble CLIP-score [58, 92, 69], as shown in Table 5.
+Figure 7 (see Appendix C) demonstrate the effectiveness of the tuned models in identifying harmful content within jailbreak prompts.
+Here, we further quantitatively assess the alignment of identified harmful content with the original jailbreak prompts by measuring semantic similarity using the ensemble CLIP-score [59, 95, 71], as shown in Table 5.
 Note that our measurement exclusively targets the examples where the defense mechanism is successful.
-We calculate the CLIP-score between the original prompts $P_{query}$ from JailbreakBench and the attack-generated jailbreak prompts, as well as the CLIP-score between $P_{query}$ and our identified harmful content (prompts/intentions).
-Specifically, the CLIP-scores for identified harmful portions are higher than those for the generated attack prompts, suggesting that the jailbreak prompts have altered the original prompts, whereas the harmful portions identified by our models are more aligned with the original prompts.
-Regarding the differences between the two tuned models, we observe that the similarity of identified harmful prompts is higher than that of harmful intentions in the context of human-based, optimization-based, and generation-based attacks, but lower than that of intentions under indirect and multilingual attacks.
-This can be attributed to the fact that human-based, optimization-based and generation-based attack prompts contain the original prompts, whereas indirect and multilingual attack prompts significantly distort the original prompts.
-In summary, these findings are consistent with our initial design intentions and demonstrate that the harmful portions identified by our models can indeed provide an explainable foundation for defense mechanisms.
-Table 6: Performance (ASRs) of SelfDefend with different shadow models against adaptive attacks (i.e., PAIR and TAP).
-Target Model Shadow Model PAIR TAP
-GPT-3.5 Llama Guard 0.38 0.39
-Llama Guard 2 0.31 0.36
-$P_{direct}$-tuned model 0.22 0.20
-$P_{intent}$-tuned model 0.25 0.18
-GPT-4 Llama Guard 0.28 0.24
-Llama Guard 2 0.24 0.20
-$P_{direct}$-tuned model 0.20 0.19
-$P_{intent}$-tuned model 0.24 0.15
-Llama-2-7b-chat Llama Guard 0.37 0.36
-Llama Guard 2 0.28 0.28
-$P_{direct}$-tuned model 0.21 0.21
-$P_{intent}$-tuned model 0.23 0.22
+We compute the CLIP-score between the original prompts $P_{query}$ from JailbreakBench and both the attack-generated jailbreak prompts and our identified harmful content (prompts/intentions).
+Higher CLIP-scores for identified harmful portions compared to attack prompts indicate that while jailbreak prompts alter the original content, the harmful content identified by our models remains closely aligned with the originals.
+Comparing these two tuned models, identified harmful prompts show higher similarity than harmful intentions in human-based, optimization-based, and generation-based attacks, but lower similarity in indirect and multilingual attacks.
+This is because the former attack types retain the original prompts, whereas the latter significantly distort them.
+In summary, these findings confirm that our models effectively identify harmful content in an explainable manner, supporting robust defense mechanisms.
 7.5 Robustness to Adaptive Jailbreaks
-Robustness to Black-box Jailbreaks.
-In §7.2, the jailbreak prompts used for evaluation are either manually designed, transferred from a surrogate model (Vicuna-7b-v1.3 [13]), or generated for the target models (GPT-3.5, GPT-4, Llama-2-7b-chat).
-To evaluate the robustness of our SelfDefend against black-box adaptive attacks, we select query-based PAIR and TAP to jailbreak the whole SelfDefend framework.
+To test SelfDefend’s robustness against adaptive attacks, we consider three schemes of adaptive jailbreaks as follows.
+Table 7: ASRs of SelfDefend with different shadow models against adaptive attacks (i.e., PAIR, TAP, and LLM-Fuzzer).
+Target Model Shadow Model PAIR TAP LLM-Fuzzer
+GPT-3.5 Llama Guard 0.38 0.39 0.61
+Llama Guard 2 0.31 0.36 0.38
+Llama Guard 3 0.23 0.29 0.15
+$P_{direct}$-tuned model 0.22 0.20 0.17
+$P_{intent}$-tuned model 0.25 0.18 0.14
+GPT-4 Llama Guard 0.28 0.24 0.23
+Llama Guard 2 0.24 0.20 0.07
+Llama Guard 3 0.23 0.24 0.07
+$P_{direct}$-tuned model 0.20 0.19 0.08
+$P_{intent}$-tuned model 0.24 0.15 0.06
+Llama-2-7b Llama Guard 0.37 0.36 0.10
+Llama Guard 2 0.28 0.28 0.04
+Llama Guard 3 0.32 0.24 0.01
+$P_{direct}$-tuned model 0.21 0.21 0.00
+$P_{intent}$-tuned model 0.23 0.22 0.01
+Mistral-7B Llama Guard 0.43 0.41 0.37
+Llama Guard 2 0.34 0.31 0.59
+Llama Guard 3 0.30 0.21 0.06
+$P_{direct}$-tuned model 0.27 0.24 0.10
+$P_{intent}$-tuned model 0.20 0.18 0.02
+Robustness to Entire Adaptive Jailbreaks.
+In §7.2, the jailbreak prompts used for evaluation are either manually designed, transferred from a surrogate model (Vicuna-7b-v1.3 [15]), or generated for the target models (GPT-3.5, GPT-4, Llama-2-7b-chat, and Mistral-7B-Instruct-v0.2).
+To evaluate the robustness of SelfDefend against black-box adaptive attacks, we select PAIR, TAP, and LLM-Fuzzer to jailbreak the entire SelfDefend.
 The results are reported in Table 7.
-First, we can observe that all our tuned models outperform the Llama Guard series, except for one case, which is that the result of $P_{intent}$-tuned model is equal to Llama Guard 2 under PAIR.
-Moreover, the $P_{direct}$-tuned model is superior to the $P_{intent}$-tuned model in defending against PAIR, whether for GPT-3.5, GPT-4, or Llama-2.
-In contrast, the $P_{intent}$-tuned model is superior to the $P_{direct}$-tuned model in resisting TAP, except on Llama-2.
+First, we observe that the $P_{direct}$-tuned and $P_{intent}$-tuned models achieve the lowest ASR, outperforming the Llama Guard series.
+Moreover, the $P_{direct}$-tuned model is superior to the $P_{intent}$-tuned model in defending against PAIR for GPT-3.5, GPT-4, and Llama-2, except for Mistral-7B-Instruct-v0.2.
+In contrast, the $P_{intent}$-tuned model is superior to the $P_{direct}$-tuned model in resisting TAP and LLM-Fuzzer, except for Llama-2.
 […]
-Table 7: ASR with the regular GCG attack on Llama-2 and the white-box GCG directly on our tuned defense models.
-Tuned Model GCG (from Table 4) White-box GCG
-$P_{direct}$-tuned model 0.06 0.08
-$P_{intent}$-tuned model 0.11 0.09
+We use GCG, RLbreaker, and LLM-Fuzzer to separately attack SelfDefend on Llama-2-7b-chat.
+For GCG, we craft the jailbreak prompt in the form “[Prefix 1] [Prefix 2] [Goal]”, where “[Prefix 1]” is optimized by GCG to attack the shadow model, and “[Prefix 2]” is optimized by GCG to attack the target model.
+For RLbreaker and LLM-Fuzzer, we first jailbreak the target model, then focus on attacking the shadow model with the jailbreak prompt.
+The results are reported in Table 8.
+We observe that such separate adaptive attack has a noticeable effect on the target LLM, with GCG achieving an ASR of over 40%, while RLbreaker and LLM-Fuzzer both exceed 70%.
+However, the ASRs of SelfDefend are significantly lower than those of the target model.
+The average ASR of the $P_{direct}$-tuned SelfDefend under these three jailbreaks is 0.193, while the average ASR of the $P_{intent}$-tuned SelfDefend is 0.023.
+This indicates that SelfDefend demonstrates encouraging robustness against separate adaptive attacks.
+Table 8: ASRs (target LLM/SelfDefend) of SelfDefend with shadow models against separate adaptive attacks.
+Method GCG RLbreaker LLM-Fuzzer
+$P_{direct}$-tuned SelfDefend 0.43/0.32 1.00/0.15 0.76/0.11
+$P_{intent}$-tuned SelfDefend 0.47/0.04 0.83/0.02 0.76/0.01
+[…]
 7.6 Robustness to Prompt Injection
@@
 Moreover, we notice that the intent prompt-tuned model is less affected by prompt injection than the direct prompt-tuned model, as evidenced by its lower ASR.
+Table 9: ASRs of individual attacks on detector models.
+Detector GCG RLbreaker
+Llama Guard [26] 1.00 0.74
+LLM Self Defense [56] 0.99 0.82
+$P_{direct}$-tuned model 0.08 0.00
+$P_{intent}$-tuned model 0.09 0.00
+8 Discussion
+SelfDefend vs.
+Guardrail.
+As briefly introduced in §4, a guardrail approach makes direct safety judgments on the harmfulness of the content itself.
+It follows the typical classification mindset to determine whether the given input (typically LLMs’ responses) is harmful or not, without considering the target models and utilizing their inherent safety alignment.
+SelfDefend, on the other hand, designs the defense from the perspective of a target model, using one instance of the target model to process the input as usual and using another instance (which could be the target model itself, as we demonstrated in §5, or a dedicated defense model described in §6) to activate the detection state simultaneously.
+Since the detection state in the shadow stack could collaborate with the answering state in the normal stack and still utilize its safety alignment, SelfDefend’s approach would theoretically have higher defense effectiveness than any guardrail methods, as we empirically tested in §7.2, while also have minimal extra delay, as shown in §7.3.
+Therefore, SelfDefend’s architecture should be preferred over the guardrail approach.
+Moreover, SelfDefend’s tuned models consider direct malicious portions or indirect malicious intentions first before making safety judgments, which better captures the essence of jailbreak detection since jailbreaking must involve malicious goals.
+Thus, our shadow models are more effective against jailbreak prompts (see §7.2) and less likely to be hacked (see §8) compared to the Llama Guard series, which make safety judgments first and then explain them—from effect to cause.
+Impact on Utility.
+We clarify that the utility of SelfDefend is not compromised by its defense effectiveness.
+First, SelfDefend is a plug-in, self-contained defense method that does not require any modifications to the target model.
+It is thus easy to deploy and does not require altering the output of the target LLM.
+Second, as previously illustrated in Table 3 and Table 4, the shadow model-enabled framework shows similar ASRs to the target LLM on the normal prompt dataset AlpacaEval.
+Since AlpacaEval contains about 85 instructions on code generation and 720 samples on daily question-answering, it is expected that the normal user experience would not be affected by SelfDefend’s defense.
 9 Conclusion
@@
 The evaluation shows that SelfDefend is lightweight and effective in mitigating a wide spectrum of jailbreak attacks while rarely undermining normal queries.
-Ethics and the Open Science
+Acknowledgements
+We thank the reviewers and the shepherd for their constructive comments.
+The HKUST authors are supported in part by an RGC CRF grant under contract C6015-23G, a research fund provided by HSBC, and a UGC TLIP project (175X-2425).
+The NTU authors are supported by the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2GC-2023-008), and NRF Investigatorship NRF-NRFI06-20200001.
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Cyber Security Agency of Singapore.
+Ethics Considerations
 Our research has meticulously addressed various ethical considerations to ensure responsible and ethical conduct.
-Firstly, regarding vulnerability disclosures and open research artifacts, we have released our datasets, raw evaluation results, and code at this anonymous GitHub link: https://github.com/SelfDefend.
-While our study required testing various jailbreak attacks on the live system (i.e., ChatGPT API service), these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server.
+Firstly, while our study required testing various jailbreak attacks on the live system (i.e., ChatGPT API service), these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server.
 We also aimed to minimize the overhead for OpenAI by distributing our experiments over a period of three months and incurred around \$2,000 in API usage fees.
@@
 The well-being of our team members was a priority, with measures in place to protect against exposure to harmful content during the research process and offering psychological support if needed.
-Although our novel defense mechanism provide an effective, low-latency, plug-and-play, and explainable solution, its potential negative outcome could potentially be used maliciously to craft better attacks on LLMs.
+Although our novel defense mechanism provides an effective, low-latency, plug-and-play, and explainable solution, its potential negative outcome could potentially be used maliciously to craft better attacks on LLMs.
 We therefore engage with the community to ensure the technology is used responsibly.
@@
 This comprehensive ethical approach not only aligns with the guidelines but also reinforces the integrity and societal value of our research.
+Open Science
+We have released our datasets, raw evaluation results, and code at this GitHub link: https://github.com/SelfDefend. We also release the artifact at https://doi.org/10.5281/zenodo.14736935.
 Appendix
-Appendix A Case Studies
+Appendix A Expanded Evaluation across Other LLMs
+As shown in Table 10, we extend the evaluation of our tuned models to other LLMs, including Claude-3.5-sonnet and Llama-2-13b-chat.
+For the Claude-3.5-sonnet API, we use the version “claude-3-5-sonnet-20241022”.
+As in the earlier evaluations on GPT-3.5/4, we use a similar setting to generate jailbreaks for Claude.
+For Llama-2-13b-chat, we use an analogous setting to Llama-2-7b-chat in prior evaluations.
+Under the target LLM Claude-3.5-sonnet, we observe that all jailbreak methods have low ASRs, indicating the resistance of this newest Claude.
+Despite this, SelfDefend, either $P_{direct}$-tuned or $P_{intent}$-tuned, achieves the lowest ASR among all defense methods, demonstrating the effectiveness of our tuned models.
+On Llama-2-13b-chat, our methods also achieve the lowest ASRs, except for ICD under PAIR.
+For normal prompts, SelfDefend shows similar ASRs to the target LLMs, indicating that the utility of the target LLMs is not compromised by our defense.
+Table 10: Jailbreak ASRs for various defense methods.
+For ICD and SafeDecoding, we present the performance of their enhanced models.
+For detection-based Perplexity Filter, SmoothLLM and Llama Guards, we report ASRs only on their detection modules.
+Target Defense Method Human Optimization Generation Indirect Multilingual Normal
+Model DAN GCG AutoDAN RLbreaker PAIR TAP LLM-Fuzzer DrAttack Puzzler MultiJail AlpacaEval
+Claude Claude-3.5-sonnet (baseline) 0.029 0.020 0.010 0.300 0.280 0.260 0.110 0.160 0.010 0.048 0.971
+ICD [75] 0.095 0.010 0.000 0.060 0.160 0.160 0.040 0.140 0.120 0.333 0.903
+Perplexity Filter [27] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
+SmoothLLM [60] 0.103 0.030 0.000 0.080 0.160 0.190 0.020 0.160 0.030 0.305 0.958
+Llama Guard [26] 0.552 0.400 0.560 0.610 0.480 0.420 0.690 0.980 0.930 0.952 0.996
+Llama Guard 2 [67] 0.432 0.140 0.200 0.370 0.380 0.360 0.240 0.910 0.620 0.559 0.990
+Llama Guard 3 [19] 0.343 0.040 0.070 0.020 0.270 0.270 0.010 0.620 0.420 0.378 0.986
+(3.5-sonnet) $P_{direct}$-tuned Shadow Stack 0.260 0.060 0.060 0.020 0.240 0.290 0.020 0.740 0.100 0.737 0.960
+$P_{direct}$-tuned SelfDefend 0.025 0.010 0.000 0.010 0.170 0.160 0.000 0.160 0.000 0.044 0.939
+$P_{intent}$-tuned Shadow Stack 0.296 0.070 0.060 0.040 0.190 0.210 0.030 0.170 0.150 0.613 0.993
+$P_{intent}$-tuned SelfDefend 0.025 0.020 0.000 0.020 0.150 0.140 0.020 0.080 0.000 0.038 0.966
+Llama-2 Llama-2-13b-chat (baseline) 0.761 0.640 0.780 0.610 0.320 0.280 0.160 0.670 1.000 0.181 0.984
+ICD [75] 0.593 0.390 0.680 0.500 0.140 0.150 0.620 0.340 0.980 0.270 0.845
+SafeDecoding [84] 0.753 0.900 0.850 0.650 0.300 0.300 0.660 0.700 1.000 1.000 0.983
+Perplexity Filter [27] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
+SmoothLLM [60] 0.911 0.850 0.950 0.920 0.600 0.590 0.840 0.690 1.000 0.321 0.996
+Llama Guard [26] 0.552 0.450 0.950 0.730 0.430 0.340 0.540 0.890 0.930 0.952 0.996
+Llama Guard 2 [67] 0.432 0.230 0.100 0.280 0.390 0.360 0.250 0.890 0.620 0.559 0.990
+Llama Guard 3 [19] 0.343 0.110 0.080 0.110 0.290 0.210 0.150 0.450 0.420 0.378 0.986
+(13b-chat) $P_{direct}$-tuned Shadow Stack 0.256 0.120 0.060 0.020 0.270 0.170 0.020 0.400 0.130 0.740 0.963
+$P_{direct}$-tuned SelfDefend 0.221 0.120 0.060 0.000 0.250 0.130 0.000 0.270 0.130 0.121 0.949
+$P_{intent}$-tuned Shadow Stack 0.286 0.090 0.050 0.060 0.270 0.190 0.050 0.120 0.200 0.619 0.990
+$P_{intent}$-tuned SelfDefend 0.244 0.090 0.050 0.030 0.220 0.150 0.010 0.100 0.200 0.111 0.975
+Appendix B Average $\Delta d$ for Various Jailbreak Prompts
+Figure 5: Average $\Delta d$ introduced by GPT-based SelfDefend across various jailbreaks.
+Figure 6: Average $\Delta d$ introduced by SelfDefend’s tuned models on Llama-2-7b-chat.
+Figure 7: Responses of defense models against a jailbreak prompt of AutoDAN.
+Appendix C Case Studies
 Successful Cases.
-Figure 5 in §7.2 presents a detection example of our methods applied to a jailbreak query, compared with the Llama Guard series.
+Figure 7 presents a detection example of our methods applied to a jailbreak query, compared with the Llama Guard series.
 The attack prompt involves creating a fake news article to spread disinformation about a political candidate.
@@
 The reason our method outperforms Llama Guards might be due to the enhanced reasoning ability of the foundation model in identifying harmful portions.
-As described in our design, the direct prompt-based model identifies harmful parts directly from the input text, whereas the intent prompt-based model summarizes the request’s intention and then identifies unsafe parts from that summary.
+As described in our design, $P_{direct}$-based model identifies harmful parts directly from the input text, whereas $P_{intent}$-based model summarizes the request’s intention and then identifies unsafe parts from that summary.
 When comparing our $P_{direct}$-tuned models with $P_{direct}$-based GPT-4, we observed that the parts they lifted from the attack prompt came from different sentences, since $P_{direct}$ emphasizes that it is sufficient to output one harmful part.
 A similar phenomenon occurs with the intent prompt, where the semantics of GPT-4 and our tuned model’s intercepts are not identical.
-In addition, please refer to Appendix C for more case of our models handling other jailbreaks or the normal prompts.
+In addition, please refer to https://github.com/selfdefend/DatasetAndResult/blob/main/examples.md for more case of our models handling other jailbreaks or normal prompts.
 As expected, both of our tuned models make satisfactory decisions on various attack or benign inputs.
@@
 Figure 9: An example of tuned models against PAIR on GPT-3.5.
-Appendix B Keywords
-Table 8: The refusal signals considered for GPT-3.5/4.
-“I’m sorry”
-“Sorry”
-“I am sorry”
+Appendix D Keywords
+Table 11: The refusal signals considered for GPT-3.5/4 and Claude.
+“I’m sorry”, “Sorry”, “I am sorry”
 “I apologize”
-“As an”
-“As an AI”
-“I’m an”
-“I’m just”
+“As an”, “As an AI”
+“I’m an”, “I’m just”
 “As a language model”
@@
 “I’m really sorry”
-“OpenAI" “unethical”
-“not ethical”
-“illegal”
-“not legal”
+“OpenAI"
+“unethical”, “not ethical”
+“illegal”, “not legal”
 “My apologies”
@@
 “Hello!”
-As mentioned in §5.2, we use a list of refusal keywords adopted by common practice [99]. They are listed in Table 8 for GPT-3.5/5 and in Table 9 for Llama-2, respectively.
-Appendix C More Examples from Our Tuned Models
-More examples from our tuned defense models are presented in Figure 9, 10, 11, 12, 13, 14, 15, and 16.
-Figure 9: An example of tuned models against DAN on GPT-3.5.
-Figure 10: An example of tuned models against GCG on GPT-3.5.
-Figure 11: An example of tuned models against AutoDAN on GPT-3.5.
-Figure 12: An example of tuned models against PAIR on GPT-3.5.
-Figure 13: An example of tuned models against TAP on GPT-3.5.
-Figure 14: An example of tuned models against Puzzler on GPT-3.5.
-Figure 15: An example of tuned models against MultiJail on GPT-3.5.
-Figure 16: Tuned models’ response to a normal prompt from AlpacaEval.
+Table 13: The refusal signals considered for Mistral.
+“I’m unable”
+“I’m sorry”
+“I am sorry”
+“I’m really sorry”
+“I cannot”
+“I’m an”
+“I’m here”
+“unethical”
+“focus on”
+As mentioned in §5.2, we use a list of refusal keywords adopted by common practice [102].
+They are listed in Table 11 for GPT-3.5/4 and Claude, Table 12 for Llama-2, and Table 13 for Mistral, respectively.